### v1

# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

Abstract
Many real-world problems, such as those with fairness constraints, involve complex expectation constraints and large datasets, necessitating the design of efficient stochastic methods to solve them. Most existing research focuses on cases with no constraint or easy-to-project constraints or deterministic constraints. In this paper, we consider nonconvex nonsmooth stochastic optimization problems with expectation constraints, for which we build a novel exact penalty model. We first show the relationship between the penalty model and the original problem. Then on solving the penalty problem, we present a single-loop SPIDER-type stochastic subgradient method, which utilizes the subgradients of both the objective and constraint functions, as well as the constraint function value at each iteration. Under certain regularity conditions (weaker than Slater-type constraint qualification or strong feasibility assumed in existing works), we establish an iteration complexity result of $O(\epsilon^{-4})$ to reach a near-$\epsilon$ stationary point of the penalized problem in expectation, matching the lower bound for such tasks. Building on the exact penalization, an $(\epsilon,\epsilon)$-KKT point of the original problem is obtained. For a few scenarios, our complexity of either the objective sample subgradient or the constraint sample function values can be lower than the state-of-the-art results in [9, 18, 30] by a factor of $\epsilon^{-2}$. Moreover, on solving two fairness-constrained problems, our method is significantly (up to 466 times) faster than the state-of-the-art algorithms, including switching subgradient method in [18] and inexact proximal point methods in [9, 30].
keywords:
stochastic, subgradient, expectation constraints, weakly convex, fairness constrained classification.
{AMS}
90C06, 90C15, 49M37, 65K05
1 Introduction
Stochastic optimization is fundamental to the statistical sciences and serves as the backbone of data-driven learning processes. In this paper, we study this topic by designing first-order methods (FOMs) for solving the expectation-constrained nonconvex programming:
(P) $\min_{{\mathbf{x}}}\,\,{f({\mathbf{x}})},\text{ s.t. }\underbrace{\mathbb{E}_{\xi_{g}\sim\mathbb{P}_{g}}[g({\mathbf{x}},\xi_{g})]}_{=:g({\mathbf{x}})}\leq 0.$
Here, $\mathbb{P}_{g}$ is an unknown probability distribution, and we will consider a general distribution and a special uniform distribution on a finite set, while $f$ is always assumed to be stochastic. Notice that a problem with multiple constraints $g_{i}({\mathbf{x}})\leq 0,i=1,\ldots,m$ can be formulated into Eq. P by setting $g({\mathbf{x}})\equiv\max_{i=1,\ldots,m}g_{i}({\mathbf{x}})$.
Substantial research has been devoted to FOMs for solving function-constrained optimization problems. Among these FOMs, most are developed for deterministic function-constrained problems or for expectation-constrained convex optimization. This is well-documented in the literature for convex scenarios [6, 7, 21, 22, 23, 26, 29, 40, 41, 44] and nonconvex scenarios [3, 4, 10, 14, 28, 27, 31]. Only a few works (e.g., [2, 9, 18, 24]) have studied FOMs for expectation-constrained nonconvex optimization. We focus on nonconvex nonsmooth expectation-constrained problems in the form of Eq. P. Throughout this paper, we assume the follows.
Assumption 1.
The following statements hold.
$\mathrm{(a)}$ The set ${\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)\cap\Omega$ is nonempty, where $\Omega:=\{{\mathbf{x}}:g({\mathbf{x}})\leq 0\}$, and ${\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$ is bounded. Specifically, there exists $D>0$ such that $\|{\mathbf{x}}_{1}-{\mathbf{x}}_{2}\|\leq D$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
$\mathrm{(b)}$ The function $f$ is $\rho_{f}$-weakly convex, and $g$ is $\rho_{g}$-weakly convex for some $\rho_{f}>0$ and $\rho_{g}\geq 0$. That is, $f+\frac{\rho_{f}}{2}\|\cdot\|^{2}$ and $g+\frac{\rho_{g}}{2}\|\cdot\|^{2}$ are convex.
$\mathrm{(c)}$ The subdifferential sets of $f$ and $g$ are nonempty in their respective domains. In addition, it holds that $\|{\boldsymbol{\zeta}}_{f}\|\leq l_{f}$ and $\|{\boldsymbol{\zeta}}_{g}\|\leq l_{g}$ for all ${\boldsymbol{\zeta}}_{f}\in\partial f({\mathbf{x}})$ and ${\boldsymbol{\zeta}}_{g}\in\partial g({\mathbf{x}})$ for any ${\mathbf{x}}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
From items (a) and (c) in the above assumption, we have $|f({\mathbf{x}}_{1})-f({\mathbf{x}}_{2})|\leq l_{f}D=:B_{f}$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$. In item (b), we allow $\rho_{g}=0$, which indicates the convexity of $g$. However, we always assume $\rho_{f}>0$ and thus consider a nonconvex objective function $f$.
Complexity measures.  We will focus on three complexity measures: objective subgradient complexity (OGC), constraint subgradient complexity (CGC) and constraint function value complexity (CFC), detailed as follows:
OGC and CGC denote the number of calls to the sample subgradients of $f$ and $g$, respectively, while CFC represents the number of evaluations of the sample constraint function $g(\cdot,\xi_{g})$.
The nonsmooth expectation constraints in Eq. P pose significant challenges for designing efficient and reliable stochastic methods, as the projection onto an expectation-constrained set is prohibitively expensive. To circumvent the costly projection step, several methods such as online convex method [43], augmented Lagrangian method (ALM) [24], inexact proximal point method (IPP) [9, 30] and switching subgradient method (SSG) [18, 23] have been explored. Among these, the online convex method, ALM, and IPP require at least two loops. The online convex method [43], developed under a convex setting, relies on a restrictive assumption that the noise associated with stochastic constraints is bounded. Under a regularity condition, [24] and [39] demonstrate the effectiveness of the ALM applied to problem Eq. P, yielding competitive numerical results. [39] establishes convergence to a KKT point. To derive complexity results to an $(\epsilon,\epsilon)$-KKT point (see Definition 1.2), [24] assumes additional smoothness on $g$ and exploits this structure to efficiently solve each nonconvex ALM subproblem by an accelerated method. Under a certain constraint qualification (CQ), [9] and [30] respectively provide convergence results and complexity analysis for stochastic IPP methods to find an $(\epsilon,\epsilon)$-KKT point. These methods iteratively solve strongly convex subproblems by incorporating a proximal term into the objective and constraint functions. Notably, the complexity results in [9] represent the state-of-the-art results for problem Eq. P under nonsmooth weakly convex settings.
Double (or more)-loop algorithms can utilize optimal subroutines for solving inner subproblems to achieve low-order complexity results. However, implementing them needs extra efforts to design appropriate checkable stopping criteria for inner subsolvers or requires careful tuning of the inner iteration counts to reach a satisfactory empirical performance. This disadvantage is especially pronounced in stochastic settings, where the randomness complicates the verification of a stopping condition and efficient tuning. Hence, single-loop methods for solving expectation-constrained problems are preferred in practice. However, such methods are still limited. Both of [2, 36] design FOMs based on penalty or ALM for constrained stochastic programming but they require smoothness of the objective (possibly with a proximable nonsmooth convex term) and constraint functions. The authors in [18] apply a single-loop SSG [34] to problem Eq. P. The SSG is easier to implement while delivering competitive numerical performance. Under a (uniform) Slater-type CQ (see 5), the SSG adapts its stepsize depending on whether the current iterate is near feasible or not, switching between fixed and adaptive steps based on Polyak’s rules [34]. This adaptation is crucial in maintaining near feasibility to achieve a near-KKT point. One critical drawback of the SSG is the requirement of frequent checks for near feasibility at generated iterates, leading to a higher CFC and overall complexity compared to IPP in [9] for solving the nonconvex problem Eq. P. Moreover, in the stochastic setting, [18] requires a light-tail assumption.
Without the need for feasibility checking, we propose a novel single-loop stochastic FOM for solving Eq. P, named 3S-Econ. At the $k$-th iteration, 3S-Econ requires unbiased stochastic subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and $\boldsymbol{\zeta}_{g}^{(k)}$ of $g$, evaluated at the iterate ${\mathbf{x}}^{(k)}$, along with an approximation $u^{(k)}$ for $g({\mathbf{x}}^{(k)})$. Using this information, it performs the following update step:
(1) ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\right),$
where $\beta>0$, $\nu>0$, and $\alpha_{k}>0$ denotes the step size at iteration $k$. We will show that 3S-Econ not only is easy to implement but also enjoys lower-order complexity than the best-known results to produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P under milder conditions; see Table 1.
1.1 Applications
Many applications can be formulated into the nonconvex nonsmooth problem Eq. P. We highlight two representative examples: Neyman-Pearson classification [35, 42] and fairness-constrained classification [25].
Example 1.1.
Neyman-Pearson classification. Let $\phi(\mathbf{x};\mathbf{a},b)$ be a loss function for a binary classification problem. Here, $\mathbf{x}$ is the variable, and ${\mathbf{a}}$ is a data point with a label $b\in\{+1,-1\}$. Denote $c>0$ as the threshold of the false-positive error, $\left\{\mathbf{a}_{i}^{+}\right\}_{i=1}^{n^{+}}$ and $\left\{\mathbf{a}_{i}^{-}\right\}_{i=1}^{n^{-}}$ as the positive-class samples (i.e., label $=1$) and negative-class samples (i.e., label $=-1$) in the training dataset, respectively. The Neyman-Pearson classification aims to minimize the more concerned false-negative error while controlling false-positive error at an acceptable level by solving
(2) $\min_{\mathbf{x}\in\mathbb{R}^{d}}\frac{1}{n^{+}}\sum_{i=1}^{n^{+}}\phi\left(\mathbf{x};\mathbf{a}_{i}^{+},1\right),\text{s.t. }\frac{1}{n^{-}}\sum_{i=1}^{n^{-}}\phi\left(\mathbf{x};\mathbf{a}_{i}^{-},-1\right)-{c}\leq 0.$
Clearly, Eq. 2 is a special instance of Eq. P with $f$ having a finite-sum structure, and $\mathbb{P}_{g}$ being a uniform distribution on the negative-class samples.
Example 1.2.
Fairness-constrained classification. This application aims to ensure that classification models treat different groups equitably, avoiding biased outcomes. Traditional models focus on maximizing accuracy, which can lead to unfair performance across different groups of data, such as males and females. Fairness constraints are applied to balance outcomes for all groups, using principles like demographic parity [1] or equalized odds [17]. Achieving fairness requires trade-offs, as improving fairness can reduce overall accuracy. This research area is particularly important in fields like healthcare and criminal justice [11]. Let us consider a classification problem with receiver operating characteristic (ROC)-based fairness [37]. Suppose there are three groups of data: an unlabeled set $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, a protected set $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$, and an unprotected set $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$. Here, the protected set contains the features associated with the minority group, while the unprotected set includes the features from the majority group. The fairness-constrained problem can be formulated as
(3) $\vspace{-1mm}\min_{{\mathbf{x}}:\|{\mathbf{x}}\|\leq D}\Phi_{\Theta}({\mathbf{x}}),\,\,\text{{s.t.}}\,\,\Phi({\mathbf{x}})\leq\Phi^{*}+\kappa_{1},$
where $\Phi_{\Theta}({\mathbf{x}}):=\max_{\theta\in\Theta}\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\geq\theta\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\geq\theta\right)\right|$ is a ROC-based fairness, $\Phi^{*}:=\min_{{\mathbf{x}}}\Phi({\mathbf{x}})$ with $\Phi({\mathbf{x}}):=\frac{1}{n}\sum_{i=1}^{n}\left[1-b_{i}\cdot{\mathbf{x}}^{\top}{\mathbf{a}}_{i}\right]_{+}$, $\Theta$ is a finite threshold set, and $\kappa_{1}>0$ is a slackness parameter. Using the sigmoid function $\sigma(\cdot)$ as a continuous relaxation of the indicator function, we reformulate the problem as
(4) $\min_{{\mathbf{x}}:\|{\mathbf{x}}\|\leq D}\Psi_{\Theta}({\mathbf{x}}),\,\,\text{{s.t.}}\,\,\Phi({\mathbf{x}})\leq\Phi^{*}+\kappa_{1},$
where $\Psi_{\Theta}({\mathbf{x}}):=\max_{\theta\in\Theta}\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\sigma\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}-\theta\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\sigma\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}-\theta\right)\right|$. Since the max operator respects the weak convexity, Eq. 4 is also a specific instance of Eq. P.
1.2 Tools: exact penalization, smoothing, and variance reduction
A key tool in our approach is the exact penalization. This tool transforms problem Eq. P into an unconstrained one. Notably, when the penalty parameter is sufficiently large, the stationarity point of the penalized problem can align with that of the original constrained problem. Building on the exact penalization, we then design a stochastic FOM to solve the resulting unconstrained problem.
Using the penalty function $h=[\cdot]_{+}:a\to\max\{a,0\}$ for an inequality constraint, an exact penalty reformulation [8, 33, 38, 45] of problem Eq. P is given by
(5) $\min_{{\mathbf{x}}}\,\,F({\mathbf{x}})=f({\mathbf{x}})+\beta g_{+}({\mathbf{x}}),\text{ with }g_{+}({\mathbf{x}})=h(g({\mathbf{x}})),$
where $\beta>0$ is the penalty parameter. The exact penalization results regarding global minimizers are established in [33]. Note that $\mathrm{Sign}([g(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g}$ belongs to $\partial g_{+}(\mathbf{x})$, where $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$. Here, $\mathrm{Sign}:\mathbb{R}\mapsto\{-1,0,1\}$ refers to the sign operator, and $\partial$ denotes the (Clarke) subdifferential [12, Section 1.2] of a continuous function. Due to the stochastic nature of $g$, it is impractical to compute the exact value of $\mathrm{Sign}([g(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g}$. Approximating this value is also challenging, because $\mathrm{Sign}$ is discontinuous, making it impossible for the discrepancies between $\mathrm{Sign}([g(\mathbf{x})]_{+})$ and $\mathrm{Sign}([g(\mathbf{x},\mathcal{B})]_{+})$ to be consistently small enough. Here, $\mathcal{B}$ is a sample set of $\xi_{g}$ over the distribution $\mathbb{P}_{g}$, and
(6) $g(\mathbf{x},\mathcal{B}):=\frac{1}{|\mathcal{B}|}\sum_{\xi_{g}\in\mathcal{B}}g({\mathbf{x}},\xi_{g}),$
with $|\mathcal{B}|$ denoting the cardinality of $\mathcal{B}$. Consequently, solving problem Eq. 5 directly by a stochastic method is challenging. These difficulties extend to algorithm design and convergence analysis, complicating the establishment of theoretical guarantees.
To resolve this issue, we introduce a smoothed function of $h$ with a smoothing parameter $\nu>0$ as follows,
$\vspace{-3mm}h^{\nu}(z)$ $:=\max_{0\leq y\leq 1}\langle y,z\rangle-\frac{\nu}{2}y^{2}=\left\{\begin{aligned} \frac{z^{2}}{2\nu},\quad&\text{ if }0\leq z\leq\nu,\\ z-\frac{\nu}{2},&\text{ if }z>\nu,\\ 0,\quad\,\,\,&\text{ otherwise.}\end{aligned}\right.$
We then present a new penalty model of problem Eq. P:
(PP) $\min_{{\mathbf{x}}}\,\,F^{\nu}({\mathbf{x}})=f({\mathbf{x}})+\beta h^{\nu}(g({\mathbf{x}})).$
Since $h^{\nu}$ is smooth and $\nabla h^{\nu}(z)=\mathrm{Proj}_{[0,1]}\left(\frac{z}{\nu}\right)$, we can access a subgradient of $h^{\nu}(g({\mathbf{x}}))$ via $\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g}$, where $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$. In addition, as $\mathrm{Proj}_{[0,1]}$ is Lipschitz continuous, the difference between $\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g}$ and $\mathrm{Proj}_{[0,1]}\left(\frac{u}{\nu}\right)\boldsymbol{\zeta}_{g}$ can be well controlled if $u$ is a good approximation of $g({\mathbf{x}})$.
We apply a SPIDER-type (or SARAH) [16, 32] variance reduction technique to estimate $g({\mathbf{x}}^{(k)})$ to obtain $u^{(k)}$ and perform the update in Eq. 1 for each $k$. The variance reduction tool, together with the smoothing tool, enables us to achieve better complexity results for our method than state-of-the-art results to produce an $(\epsilon,\epsilon)$-KKT point of (P).
1.3 Contributions
We address key challenges in solving nonconvex nonsmooth expectation-constrained optimization problems. Our contributions are threefold.
First, to address the expectation constraint involved in Eq. P, we introduce a new smooth penalty function $h^{\nu}\circ g$. Leveraging this function, we propose solving problem Eq. PP. For any given $\epsilon>0$, we demonstrate that, under easily verifiable conditions, any near-$\epsilon$ stationary point of problem Eq. PP (in expectation) is an $(O(\epsilon),O(\epsilon))$-KKT point of problem Eq. P (in expectation). This property enables us to pursue an $(\epsilon,\epsilon)$-KKT point of problem Eq. P by finding a near-$O(\epsilon)$ stationary point of problem Eq. PP. Our framework is well-suited for many applications such as problem Eq. 4, which involves a nonconvex objective and convex expectation constraints satisfying the required assumptions.
Second, we exploit the structure of problem Eq. PP and propose a single-loop SPIDER-type stochastic FOM to solve it, with updates given in Eq. 1. At each iteration, the SPIDER technique [16] is employed to generate an approximation $u^{(k)}$ of the constraint function value. Using the tool of Moreau envelope, we demonstrate that the proposed method can, in expectation, find a near-$\epsilon$ stationary point of problem Eq. PP within $O(\epsilon^{-4})$ iterations. This result matches the known lower bound for such tasks [19], highlighting the efficiency of our approach. Combining this with the theory of exact penalization, we further analyze the complexity of obtaining an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation.
Based on the structure of $g$, we categorize the complexity results of our method as follows:
  1. 1.
    A generic stochastic $g$: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\epsilon^{-6})$
  2. 2.
    A finite-sum structured [^1] Here, we assume that $\mathbb{P}_{g}$ is a uniform distribution over a finite set of size $N$, where $N<\epsilon^{-4}$. $g$: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\sqrt{N}\epsilon^{-4})$.
When $g$ is generic stochastic, our complexity result is lower than state-of-the-art results by a factor of $\epsilon^{-2}$ in either OGC or CFC; when $g$ enjoys a finite-sum structure, the highest order among our results of OGC, CGC, and CFC is $O(\sqrt{N}\epsilon^{-4})$, representing an improvement over the previously best-known $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$ bound. Detailed comparisons are provided in Table 1. In this table, the sharpness assumption and the gradient-away assumption that we make are weaker than the (uniform) Slater-type CQ in [18, 30], which is weaker than a strong feasibility condition (see Eqn. Eq. 41 below) in [9]. These assertions will be verified in Appendix A.
Third, the ability to handle nonsmooth problems using a single-loop approach eliminates the need for tuning inner iterations, distinguishing the proposed method from double-loop methods, such as those in [9, 24, 30]. This simplification reduces computational overhead and makes the method more accessible for real-world applications. In numerical experiments, we test both the deterministic and stochastic versions of our method, for two problems across two datasets. Notably, the deterministic version is significantly faster than state-of-the-art methods, and the stochastic version achieves even more impressive speedups.
Methods SL structure of $g$ CQ on $g$ OGC CGC CFC
IPP [9] no stochastic convex $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
finite-sum convex MFCQ, and $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
stochastic WC strong feasibility $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
finite-sum WC $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
SSG [18] yes stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-8})$
finite-sum convex $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(N\epsilon^{-4})$
WC none
Ours yes stochastic convex sharpness, or $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
finite-sum convex gradient away $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
stochastic WC gradient away, and $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
finite-sum WC $B_{g}>\rho_{g}D^{2}$ $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
finite-sum WC Slater-type CQ $O(\epsilon^{-4})$ $O(N\epsilon^{-4})$ $O(N\epsilon^{-4})$
[…]
1.4 Notations and definitions
[…]
(7) $\phi^{\prime}({\mathbf{x}};{\mathbf{v}})\geq{\mathbf{v}}^{\top}{\boldsymbol{\gamma}},\text{ for any }{\boldsymbol{\gamma}}\in\partial\phi({\mathbf{x}}).$
Definition 1.1 (Moreau Envelope [13]).
For a $\rho_{\phi}$-weakly convex function $\phi$, the Moreau envelope and the proximal mapping for any $t\in(0,\frac{1}{\rho_{\phi}})$ are defined by $\phi_{t}({{\mathbf{x}}}):=\min_{{\mathbf{z}}}\left\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\right\}$ and $\operatorname{Prox}_{t\phi}({{\mathbf{x}}}):=\arg\min_{\mathbf{z}}\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\}.$
We give the definition of stationarity as follows. In deterministic cases—where randomness is not involved—the expectation operator $\mathbb{E}$ should be removed from the definitions below.
Definition 1.2 (Stationarity).
Let $\epsilon>0$ and $\delta>0$, and suppose $\psi$ and $\phi$ are $\rho_{\psi}$- and $\rho_{\phi}$-weakly convex functions.
$\mathrm{(a)}$ We call ${\mathbf{x}}^{*}$ a near-$\epsilon$ stationary point of $\min_{{\mathbf{x}}}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)$, and there exists $0<t<\frac{1}{\rho_{\phi}}$, such that $\mathbb{E}[\|\nabla\phi_{t}({\mathbf{x}}^{*})\|^{2}]=\mathbb{E}[\|\frac{1}{t}(\operatorname{Prox}_{t\phi}({{\mathbf{x}}^{*}})-{\mathbf{x}}^{*})\|^{2}]\leq~{}\epsilon^{2}.$
$\mathrm{(b)}$ We call ${\mathbf{x}}^{*}$ an $\epsilon$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if it holds that ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)\cap{\mathrm{dom}}(\psi)$, and there exists $\lambda_{{\mathbf{x}}^{*}}\geq 0$ such that
$\mathbb{E}\left[\textnormal{dist}(\mathbf{0},\partial\phi({\mathbf{x}}^{*})+\lambda_{{\mathbf{x}}^{*}}\partial\psi({\mathbf{x}}^{*}))\right]\leq\epsilon,\,\,\mathbb{E}\left[\psi({\mathbf{x}}^{*})\right]\leq\epsilon,\text{ and }|\lambda_{{\mathbf{x}}^{*}}\psi({\mathbf{x}}^{*})|\leq\epsilon.$
$\mathrm{(c)}$ We call ${\mathbf{x}}^{*}$ a $(\delta,\epsilon)$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if it holds that ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)\cap{\mathrm{dom}}(\psi)$, and $\mathbb{E}[\|{\mathbf{x}}^{*}-\widetilde{{\mathbf{x}}}^{*}\|]\leq\delta$ with $\widetilde{{\mathbf{x}}}^{*}$ being an $\epsilon$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation.
1.5 Organizations
The rest of this paper is organized as follows. Section 2 introduces the exact penalty models. In Section 3, we propose a single-loop SPIDER-type stochastic subgradient method and establish its iteration complexity results. Numerical experiments are presented in Section 4, showcasing the performance of our method. Finally, Section 5 concludes the paper.
2 Model analysis
In this section, we establish the relationships among models Eq. PEq. 5 and Eq. PP, which are presented in the following diagram for the convenience of the reader. Details are given in the mentioned lemmas and theorem.
[…]
2.1 Exact penalization with penalty function $[\cdot]_{+}$
We study the relationship between models Eq. P and Eq. 5 in this subsection. Since $g$ is $\rho_{g}$-weakly convex, it follows that $[g({\mathbf{x}})]_{+}=\max\{g({\mathbf{x}})+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2},\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}\}-\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is $\rho_{g}$-weakly convex. Recalling that $f$ is $\rho_{f}$-weakly convex, $F$ in Eq. 5 is $\rho$-weakly convex with $\rho:=\rho_{f}+\beta\rho_{g}$.
Under either of the two assumptions below, we establish the equivalence between the $\epsilon$-KKT point of problem Eq. 5 and that of Eq. P.
Assumption 2 (sharpness).
There exist $\theta>0$ and $B_{g}>0$ such that for any ${\mathbf{x}}$ satisfying $0<g({\mathbf{x}})\leq B_{g}$, it holds $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$.
Assumption 3 ((uniformly) gradient-away).
There exist $\theta>0$ and $B_{g}>0$ such that for any ${\mathbf{x}}$ satisfying $0<g({\mathbf{x}})\leq B_{g}$, it holds $\min_{\boldsymbol{{\boldsymbol{\zeta}}}_{g}\in\partial g(\mathbf{x})}\left\|\boldsymbol{{\boldsymbol{\zeta}}}_{g}\right\|\geq\theta.$
Lemma 2.1.
Suppose 1 and 2 hold, and $g$ is convex. Given $\epsilon>0$, let $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$. If ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$ is an $\epsilon$-KKT point of problem Eq. 5, then ${\mathbf{x}}^{*}\in\Omega$, i.e., ${\mathbf{x}}^{*}$ is a feasible point of problem Eq. P, and ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P.
Proof 2.2.
Since ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. 5, there exist ${\boldsymbol{\gamma}}\in\partial f({\mathbf{x}}^{*})$ and ${\boldsymbol{\zeta}}\in\partial g_{+}({\mathbf{x}}^{*})$ such that
(8) $\|{\boldsymbol{\gamma}}+\beta{\boldsymbol{\zeta}}\|\leq\epsilon.$
We then derive that for any ${\mathbf{v}}\in\mathbb{R}^{d}$,
(9) $F^{\prime}({\mathbf{x}}^{*};{\mathbf{v}})=f^{\prime}({\mathbf{x}}^{*};{\mathbf{v}})+\beta g_{+}^{\prime}({\mathbf{x}}^{*};{\mathbf{v}})\geq{\mathbf{v}}^{\top}{\boldsymbol{\gamma}}+\beta{\mathbf{v}}^{\top}{\boldsymbol{\zeta}}\geq-\epsilon\|{\mathbf{v}}\|,$
where the first inequality follows from Eq. 7, and the second inequality is by Eq. 8.
Second, we let ${\mathbf{x}}^{+}:=\mathrm{Proj}_{\Omega}({\mathbf{x}}^{*})$, and $\overline{{\mathbf{v}}}:={\mathbf{x}}^{+}-{\mathbf{x}}^{*}.$ For any $10<s<$ 1, we derive
(10) $g_{+}({\mathbf{x}}^{*}+s\overline{{\mathbf{v}}})-g_{+}({\mathbf{x}}^{*})$ $\leq(1-s)g_{+}({\mathbf{x}}^{*})+sg_{+}({\mathbf{x}}^{*}+\overline{{\mathbf{v}}})-g_{+}({\mathbf{x}}^{*})$
$=-sg_{+}({\mathbf{x}}^{*})\leq-{s\theta}{}\textnormal{dist}({\mathbf{x}}^{*},\Omega),$
where the first inequality follows from the convexity of $g$, and the second inequality holds by 2 and $\textnormal{dist}({\mathbf{x}}^{*},\Omega)=0$ if $g({\mathbf{x}}^{*})\leq 0$. In addition, we have $f({\mathbf{x}}^{*}+s\overline{{\mathbf{v}}})-f({\mathbf{x}}^{*})\leq l_{f}s\|\overline{{\mathbf{v}}}\|=l_{f}s\textnormal{dist}({\mathbf{x}}^{*},\Omega)$. Combining this with Eq. 10, we arrive at $\frac{1}{s}\left(F({\mathbf{x}}^{*}+s\overline{{\mathbf{v}}})-F({\mathbf{x}}^{*})\right)\leq-\beta\theta\textnormal{dist}({\mathbf{x}}^{*},\Omega)+l_{f}\textnormal{dist}({\mathbf{x}}^{*},\Omega)$. Letting $s\downarrow 0$, we obtain
[…]
Since $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$, the above inequality holds only if ${\mathbf{x}}^{*}\in\Omega$.
Now, if $g({\mathbf{x}}^{*})=0$, we obtain from the convexity of $g$ that $\partial g_{+}({\mathbf{x}}^{*})={\mathrm{co}}\{\partial g({\mathbf{x}}^{*}),\mathbf{0}\}$ by [12, Proposition 2.3.12]. Thus, there exists $0\leq t\leq 1$ such that ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$. Then ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P with multiplier $\beta t$ by Eq. 8 and ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$. Otherwise, i.e., if $g({\mathbf{x}}^{*})<0$, we have ${\boldsymbol{\zeta}}=\mathbf{0}$. It then yields that ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P with multiplier 0 by Eq. 8. The proof is then completed.
Lemma 2.3.
Suppose 1 and 3 hold. Given $\epsilon>0$, let $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$. If ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$ is an $\epsilon$-KKT point of problem Eq. 5, it then holds ${\mathbf{x}}^{*}\in\Omega$, and ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P.
Proof 2.4.
Since ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. 5, there exits ${\boldsymbol{\gamma}}\in\partial f({\mathbf{x}}^{*})$ and ${\boldsymbol{\zeta}}\in\partial g_{+}({\mathbf{x}}^{*})$ such that Eq. 8 holds.
We claim $g_{+}({\mathbf{x}}^{*})=0$, i.e., $g({\mathbf{x}}^{*})\leq 0$. If the claim is not true, we have $g({\mathbf{x}}^{*})>0$ and then $\partial g_{+}({\mathbf{x}}^{*})=\partial g({\mathbf{x}}^{*})$. We then derive that
(12) $\epsilon$ $\geq{}\|\beta{\boldsymbol{\zeta}}+{\boldsymbol{\gamma}}\|\geq{\beta}\|{\boldsymbol{\zeta}}\|-\|{\boldsymbol{\gamma}}\|\geq{\beta}\|{\boldsymbol{\zeta}}\|-l_{f}\geq\beta\theta-l_{f},$
where the third inequality holds by 1(c), and the fourth one follows from 3 and ${\boldsymbol{\zeta}}\in\partial g_{+}({\mathbf{x}}^{*})=\partial g({\mathbf{x}}^{*})$. This contradicts to $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$. Hence $g_{+}({\mathbf{x}}^{*})=0$.
Now, if $g({\mathbf{x}}^{*})=0$, we obtain $\partial g_{+}({\mathbf{x}}^{*})={\mathrm{co}}\{\partial g({\mathbf{x}}^{*}),\mathbf{0}\}$ from the weak convexity of $g$ and [12, Proposition 2.3.12]. Thus, there exists $0\leq t\leq 1$ such that ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$. Then ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P with multiplier $\beta t$ by Eq. 8 and ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$. Otherwise, i.e., if $g({\mathbf{x}}^{*})<0$, we have ${\boldsymbol{\zeta}}=\mathbf{0}$. It then yields that ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P with multiplier 0 by Eq. 8. The proof is then completed.
The following corollary holds directly from Lemma 2.1 and Lemma 2.3.
Corollary 2.5.
Under 1, the following statements hold with $\beta>~{}\frac{l_{f}}{\theta}$.
$\mathrm{(i)}$ Suppose 2 hold, and $g$ is convex. For all ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$, ${\mathbf{x}}^{*}\not\in\Omega$, and ${\mathbf{x}}^{*}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, it holds $\textnormal{dist}(\mathbf{0},\partial f({\mathbf{x}}^{*})+\beta\partial g_{+}({\mathbf{x}}^{*}))\geq\beta\theta-l_{f}$.
$\mathrm{(ii)}$ Suppose 3 hold. For all ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$, ${\mathbf{x}}^{*}\not\in\Omega$, and ${\mathbf{x}}^{*}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, it holds $\textnormal{dist}(\mathbf{0},\partial f({\mathbf{x}}^{*})+\beta\partial g_{+}({\mathbf{x}}^{*}))\geq\beta\theta-l_{f}$.
2.2 Exact penalization with a “smoothed” penalty approach
In this subsection, we establish the relationship between models Eq. P and Eq. PP. We begin by showing that the function $F^{\nu}$ in (PP) is weakly convex.
Lemma 2.6.
Suppose 1 holds. For any given $\nu>0$, the function $F^{\nu}$ defined in Eq. PP is $\rho$-weakly convex, where $\rho=\rho_{f}+\beta\rho_{g}$.
Proof 2.7.
To obtain the desired result, we only need to prove $h^{\nu}(g({\mathbf{x}}))$ is $\rho_{g}$-weakly convex. Notice that
$h^{\nu}(g({\mathbf{x}}))+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}=\max_{0\leq y\leq 1}y\cdot g({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}.$
Since $g$ is $\rho_{g}$-weakly convex, $y\cdot g({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex with respect to ${\mathbf{x}}$ for all $0\leq y\leq 1$. Thus by Danskin’s theorem [8], $h^{\nu}(g({\mathbf{x}}))+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex.
Next, we introduce some notations and key properties of the Moreau envelope. Denote $F_{\frac{1}{2\rho}}$ and $F^{\nu}_{\frac{1}{2\rho}}$ as the Moreau envelope of $F$ and $F^{\nu}$ with parameter $\frac{1}{2\rho}$, respectively. We let
(13) $\overline{\mathbf{x}}^{*}=\arg\min_{{\mathbf{x}}}F({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2},\quad\widehat{{\mathbf{x}}}^{*}=\arg\min_{{\mathbf{x}}}F^{\nu}({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2}.$
We then have
(14) $\|\nabla F_{\frac{1}{2\rho}}({\mathbf{x}}^{*})\|=2\rho\|\overline{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|,\ \|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{*})\|=2\rho\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|,\text{and}$
(15) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\mathrm{Proj}_{[0,1]}\left(\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\right)\partial g\left(\widehat{\mathbf{x}}^{*}\right)+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right).$
In addition, by [13, Eqn. (4.3)], it holds that
(16) $F(\overline{{\mathbf{x}}}^{*})\leq F({\mathbf{x}}^{*}),\text{ and }\textnormal{dist}(\mathbf{0},\partial F(\overline{{\mathbf{x}}}^{*}))\leq\big{\|}\nabla F_{\frac{1}{2\rho}}({\mathbf{x}}^{*})\big{\|}.$
We now present our main results to demonstrate the effectiveness of solving a smoothed exact penalty problem for a deterministic or stochastic case.
Theorem 2.8.
Suppose that 1 holds. Given $\epsilon>0$, let
(17) $\small\beta>\frac{1}{\theta}(l_{f}+\epsilon),\ 0<\overline{\epsilon}\leq\epsilon\min\left\{\frac{\beta\theta-l_{f}}{2B_{g}},\frac{\beta\theta-l_{f}}{\beta\theta},2\rho\right\},\text{ and }0<\nu\leq\min\left\{B_{g},\frac{\epsilon}{\beta},\frac{\epsilon}{2}\right\}.$
If ${\mathbf{x}}^{*}$ is a near-$\overline{\epsilon}$ stationary point of problem Eq. PP with $t=\frac{1}{2\rho}$ (see Definition 1.2$\mathrm{(a)}$) deterministically or in expectation, then $\|{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}^{*}\|\leq\epsilon$ holds deterministically or in expectation, where $\widehat{{\mathbf{x}}}^{*}$ is defined in Eq. 13. In addition, for each of the following three cases, $\widehat{{\mathbf{x}}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P deterministically or in expectation.
  • $\mathrm{(i)}$
[…]
  • $\mathrm{(ii)}$
[…]
  • $\mathrm{(iii)}$
[…]
Proof 2.9.
We only need to the prove the theorem in the expectation sense. For the deterministic case, one can simply drop the expectation notation $\mathbb{E}$.
First, according to the definition of ${\mathbf{x}}^{*}$efinition 1.2 gives $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|^{2}\right]\leq~{}\overline{\epsilon}^{2}.$ Because $\overline{\epsilon}\leq 2\rho\epsilon$, if follows that
(18) $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|\right]\leq\overline{\epsilon}\text{ and }\mathbb{E}\left[\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|\right]\leq\epsilon.$
Second, we discuss two events: ${\mathcal{A}}_{1}$ as the case of $\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\leq 1$ and ${\mathcal{A}}_{2}$ as the case of $\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}>1$. Let $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$ denote their probabilities. Also, we let $\lambda_{\widehat{\mathbf{x}}^{*}}$ be defined as $\lambda_{\widehat{\mathbf{x}}^{*}}=\beta\mathrm{Proj}_{[0,1]}(\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu})\geq 0$ when ${\mathcal{A}}_{1}$ happens and $\lambda_{\widehat{\mathbf{x}}^{*}}=0$ when ${\mathcal{A}}_{2}$ happens.
For event ${\mathcal{A}}_{1}$, we have $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\lambda_{\widehat{\mathbf{x}}^{*}}\partial g\left(\widehat{\mathbf{x}}^{*}\right)\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|$ by Eq. 15 and triangle inequality. Also, because $\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\leq 1$, we have $g(\widehat{\mathbf{x}}^{*})\leq\nu\leq\frac{\epsilon}{2}.$ Moreover, if it holds $g(\widehat{\mathbf{x}}^{*})\leq 0$, then $\lambda_{\widehat{\mathbf{x}}^{*}}=0$ by its definition, and $|\lambda_{\widehat{\mathbf{x}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)|=0\leq\epsilon$; if $g(\widehat{\mathbf{x}}^{*})>0$, then $|\lambda_{\widehat{\mathbf{x}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)|=\lambda_{\widehat{\mathbf{x}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)\leq\beta g\left(\widehat{\mathbf{x}}^{*}\right)\leq\beta\nu\leq\epsilon$. Hence, we obtain
(19) $\textstyle\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\lambda_{\widehat{{\mathbf{x}}}^{*}}\partial g\left(\widehat{\mathbf{x}}^{*}\right)\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|,\,\,g(\widehat{\mathbf{x}}^{*})\leq\nu\leq\frac{\epsilon}{2},\text{ and }|\lambda_{\widehat{{\mathbf{x}}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)|\leq\epsilon.$
For event ${\mathcal{A}}_{2}$, we have from $g\left(\widehat{\mathbf{x}}^{*}\right)>\nu$ and Eq. 15 that
(20) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\partial g_{+}\left(\widehat{\mathbf{x}}^{*}\right)+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right).$
Thus $\widehat{\mathbf{x}}^{*}$ satisfies the optimality condition of the left minimization problem in Eq. 13, and we obtain $\widehat{\mathbf{x}}^{*}=\overline{\mathbf{x}}^{*}$. Below we show $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$ under event $\mathcal{A}_{2}$ for all the three cases (i)–(iii) in the statement of Theorem 2.8.
For case (i), there exists ${\mathbf{x}}_{\mathrm{feas}}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)\cap\Omega$ by 1(a). From Eq. 13, we have
$f(\overline{{\mathbf{x}}}^{*})+\beta g_{+}(\overline{{\mathbf{x}}}^{*})+\rho\|\overline{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|^{2}$ $\leq f({\mathbf{x}}_{\mathrm{feas}})+\beta g_{+}({\mathbf{x}}_{\mathrm{feas}})+\rho\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2}$
(21) $=f({\mathbf{x}}_{\mathrm{feas}})+\rho\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2}.$
Now because $\beta\geq\frac{B_{f}+\rho_{f}D^{2}}{B_{g}}$, $\rho=\rho_{f}+\beta\rho_{g}=\rho_{f}$, and $\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2}\leq D^{2}$ from 1(a), we arrive at
$g_{+}(\widehat{{\mathbf{x}}}^{*})=g_{+}(\overline{{\mathbf{x}}}^{*})\leq\frac{1}{\beta}(f({\mathbf{x}}_{\mathrm{feas}})-f(\overline{{\mathbf{x}}}^{*})+\rho_{f}\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2})\leq\frac{B_{f}+\rho_{f}D^{2}}{\beta}\leq B_{g}.$
For case (ii), we still have Eq. 21. Thus, $f(\overline{{\mathbf{x}}}^{*})+\beta g_{+}(\overline{{\mathbf{x}}}^{*})+\rho\|\overline{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|^{2}\leq f({\mathbf{x}}_{\mathrm{feas}})+\rho D^{2}$ by 1(a). Now by $\rho_{g}D^{2}<B_{g}$, $\beta\geq\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-\rho_{g}D^{2}}$, and $\rho=\rho_{f}+\beta\rho_{g}$, it follows
$g_{+}(\widehat{{\mathbf{x}}}^{*})=g_{+}(\overline{{\mathbf{x}}}^{*})\leq\frac{1}{\beta}\Bigl{(}f({\mathbf{x}}_{\mathrm{feas}})-f(\overline{{\mathbf{x}}}^{*})+\rho D^{2}\Bigr{)}\leq\frac{B_{f}+\rho D^{2}}{\beta}\leq B_{g}.$
For case (iii), we have $f(\overline{{\mathbf{x}}}^{*})+\beta g_{+}(\overline{{\mathbf{x}}}^{*})+\rho\|\overline{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|^{2}\leq f({{\mathbf{x}}}^{*})+\beta g_{+}({{\mathbf{x}}}^{*})$ by the definition of $\overline{\mathbf{x}}^{*}$. Using $|f(\overline{{\mathbf{x}}}^{*})-f({{\mathbf{x}}}^{*})|\leq B_{f}$, $\beta\geq\frac{2B_{f}}{B_{g}}$, and $g_{+}({{\mathbf{x}}}^{*})\leq\frac{B_{g}}{2}$, we derive that $g_{+}(\widehat{{\mathbf{x}}}^{*})=g_{+}(\overline{{\mathbf{x}}}^{*})\leq\frac{B_{f}}{\beta}+g_{+}({{\mathbf{x}}}^{*})\leq B_{g}$.
Hence, for all three cases under event $\mathcal{A}_{2}$, we have $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$. Thus, by 1(c) and $\lambda_{\widehat{\mathbf{x}}^{*}}=0$, we have
(22) $\textnormal{dist}\left(\mathbf{0},\partial f(\widehat{\mathbf{x}}^{*})+\lambda_{\widehat{\mathbf{x}}^{*}}\partial g(\widehat{\mathbf{x}}^{*})\right)\leq l_{f},\,\,g(\widehat{\mathbf{x}}^{*})\leq B_{g},\text{ and }\lambda_{\widehat{\mathbf{x}}^{*}}g(\widehat{\mathbf{x}}^{*})=0.$
Furthermore, $g(\widehat{\mathbf{x}}^{*})>\nu$ implies $\widehat{\mathbf{x}}^{*}\notin\Omega$. Hence by Corollary 2.5 and $g(\widehat{\mathbf{x}}^{*})\leq B_{g}$, it holds $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\partial g_{+}(\widehat{\mathbf{x}}^{*})\right)\geq\beta\theta-l_{f}$. This yields
[…]
Now, we combine events $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. From Eq. 19 and Eq. 23, and by the definition of $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$, we have
[…]
Hence from $\overline{\epsilon}\leq\frac{\beta\theta-l_{f}}{\beta\theta}\epsilon$, it follows that $\mathrm{Prob}_{2}\leq\frac{\overline{\epsilon}}{\beta\theta-l_{f}}\leq\frac{\epsilon}{\beta\theta}.$ We finally compute expectations based on these probabilities. It holds
[…]
$\leq\overline{\epsilon}+\frac{\overline{\epsilon}l_{f}}{\beta\theta-l_{f}}\leq\epsilon.$
In addition, we have $\mathbb{E}[g(\widehat{\mathbf{x}}^{*})]{\leq}\nu\mathrm{Prob}_{1}+B_{g}\mathrm{Prob}_{2}\leq\nu+\frac{B_{g}\overline{\epsilon}}{\beta\theta-l_{f}}\leq\epsilon,\text{ and }|\lambda_{\widehat{\mathbf{x}}^{*}}g(\widehat{\mathbf{x}}^{*})|\leq\epsilon,$ where we have used Eq. 19Eq. 22, $\nu\leq\frac{\epsilon}{2}$ and $\overline{\epsilon}\leq\frac{\beta\theta-l_{f}}{2B_{g}}$. Therefore, $\widehat{\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P in expectation. This completes the proof.
3 A single-loop SPIDER-type stochastic subgradient algorithm
In this section, we present our algorithm for solving problem Eq. P. It is developed based on a subgradient method for problem Eq. PP by using the SPIDER technique to estimate constraint function values. The algorithm framework is outlined in Section 3.1. In Section 3.2, we prove that in expectation, the algorithm can find a near-$\epsilon$ stationary point of problem Eq. PP. Building on the results presented in the previous section, we further establish in Section 3.3 that an $(\epsilon,\epsilon)$-KKT point in expectation of problem Eq. P can be obtained.
3.1 Algorithm framework
We employ the SPIDER technique [16] to obtain an estimator $u^{(k)}$ of $g$ at the $k$-th iteration. Specifically, every $q$ iterations, we draw a big batch of $S_{1}$ samples from $\mathbb{P}_{g}$, while for other iterations, a smaller batch of $S_{2}$ samples is drawn. The estimator is set as
(25) $u^{(k)}=\left\{\begin{array}[]{lll}g({\mathbf{x}}^{(k)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{1},&\text{ if }\mathrm{mod}(k,q)=0,\\[2.84526pt] u^{(k-1)}+g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k-1)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{2},&\text{ otherwise},\end{array}\right.$
where $\mathcal{B}_{k}$ contains identically independent samples from $\mathbb{P}_{g}$, and $g(\mathbf{x},\mathcal{B})$ is defined in Eq. 6. With $u^{(k)}$, we then take a sample subgradient of $f$ and $g$ at ${\mathbf{x}}^{(k)}$ and perform the update in Eq. 1. The pseudocode is given in Algorithm 1.
Algorithm 1 A SPIDER-type Stochastic Subgradient algorithm
1:  Input: initial point ${\mathbf{x}}^{(0)}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, $\beta>0$, $q>0$, $S_{1},S_{2}>0$, and $K>0$.
2:  for $k=0,1,\cdots,Kq-1$ do
3:     Set $u^{(k)}$ by Eq. 25.
4:     Choose $\alpha_{k}>0$, generate a (stochastic) subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and a (stochastic) subgradient $\boldsymbol{\zeta}_{g}^{(k)}$ of $g$.
5:     Update ${\mathbf{x}}$ by Eq. 1.
6:  end for
For the convenience of analysis, we denote
(26) $\boldsymbol{\zeta}_{F^{\nu}}^{(k)}=\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\text{, and }$
${\mathbf{w}}^{(k)}=\left(\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g}^{(k)}.$
Then the update in Eq. 1 can be rewritten as
(27) ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right).$
Throughout this section, we make the following assumption, where for brevity of notations, we let $\mathbb{E}_{k}[\cdot]:=\mathbb{E}\left[\cdot\mid{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\right]$, representing the conditional expectation given $\{{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\}$ for any $k\geq 0$. In addition, $\mathrm{Var}(\boldsymbol{\zeta}):=\mathbb{E}[\|\boldsymbol{\zeta}-\mathbb{E}[\boldsymbol{\zeta}]\|^{2}]$ for a random vector $\boldsymbol{\zeta}$.
Assumption 4.
The following conditions hold.
  • (a)
    Unbiased function estimate and Lipschitz continuity of $g$: For all $k\geq 0$, the samples in $\mathcal{B}_{k}$ are mutually independent. It holds $\mathbb{E}_{k}[g({\mathbf{x}}^{(k)},\xi_{g})]=g({\mathbf{x}}^{(k)})$ and $\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k)})\right|^{2}\right]\leq\sigma^{2}$ for any $\xi_{g}\in\mathcal{B}_{k}$ and some $\sigma>0$. Moreover, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathrm{dom}}(g)$, it holds $\mathbb{E}\left[\left|g({\mathbf{x}},\xi_{g})-g({\mathbf{y}},\xi_{g})\right|^{2}\right]\leq L_{g}^{2}\|{\mathbf{x}}-{\mathbf{y}}\|^{2}.$
  • (b)
    Unbiased Stochastic Subgradients: For each $k$, $\mathbb{E}_{k}[\boldsymbol{\zeta}_{f}^{(k)}]\in\partial f({\mathbf{x}}^{(k)})$ and $\mathbb{E}_{k}[\boldsymbol{\zeta}_{g}^{(k)}]\in\partial g({\mathbf{x}}^{(k)})$. In addition, there are $\sigma_{1}\geq 0$ and $\sigma_{2}\geq 0$ such that $\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})\leq\sigma_{1}^{2}$ and $\mathrm{Var}(\boldsymbol{\zeta}_{g}^{(k)})\leq\sigma_{2}^{2}$.
By 4(b), it follows that $\mathbb{E}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$.
3.2 Convergence analysis and complexity results
In this subsection, we present the convergence analysis of Algorithm 1. First, we derive a bound on the difference between $u^{(k)}$ and $g({\mathbf{x}}^{(k)})$.
Lemma 3.1.
Under 1 and 4, let $\{{\mathbf{x}}^{(k)},u^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1. Define $r_{g}^{(k)}:=\mathbb{E}\left[|u^{(k)}-g({\mathbf{x}}^{(k)})|^{2}\right]$. Then, for all $k\geq 0$, it holds $r_{g}^{(k)}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
Proof 3.2.
Case (i): When $\mathrm{mod}(k,q)=0$,  Eq. 25 and 4(a) give
(28) $r_{g}^{(k)}=\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k)})\right|^{2}\right]\leq\frac{\sigma^{2}}{S_{1}}.$
Case (ii): When $\mathrm{mod}(k,q)>0$, it follows that
[…]
$=r_{g}^{(k-1)}+\mathbb{E}\left[\left|g({\mathbf{x}}^{(k-1)})+g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k-1)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k)})\right|^{2}\right]$
$=r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{g}\in\mathcal{B}_{k}}\mathbb{E}\left[\left|g({\mathbf{x}}^{(k-1)})+g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k-1)},\xi_{g})-g({\mathbf{x}}^{(k)})\right|^{2}\right]$
$\leq r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{g}\in\mathcal{B}_{k}}\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k-1)},\xi_{g})\right|^{2}\right]$
$\leq r_{g}^{(k-1)}+\frac{L_{g}^{2}}{S_{2}}\mathbb{E}\left[\|{\mathbf{x}}^{(k)}-{\mathbf{x}}^{(k-1)}\|^{2}\right]\leq r_{g}^{(k-1)}+\frac{\alpha_{k-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}},$
where the second equality holds due to the unbiasedness of $g({\mathbf{x}},\xi_{g})$, the third equality follows from the mutual independence of samples in $\mathcal{B}_{k}$, the second inequality follows by 4(a), and the last one uses
(30) $\mathbb{E}\left[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|^{2}\right]=\textstyle\alpha_{k}^{2}\mathbb{E}\left[\left\|\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\right)\right\|^{2}\right]$
$\leq 2\alpha_{k}^{2}\left(\mathbb{E}\big{[}\|\boldsymbol{\zeta}_{f}^{(k)}\|^{2}+\beta^{2}\|\boldsymbol{\zeta}_{g}^{(k)}\|^{2}\big{]}\right)$
$\leq 2\alpha_{k}^{2}\left(\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})+\|\mathbb{E}[\boldsymbol{\zeta}_{f}^{(k)}]\|^{2}+\beta^{2}\mathrm{Var}(\boldsymbol{\zeta}_{g}^{(k)})+\beta^{2}\|\mathbb{E}[\boldsymbol{\zeta}_{g}^{(k)}]\|^{2}\right)$
$\leq 2\alpha_{k}^{2}(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})=\alpha_{k}^{2}L_{F}^{2}.$
Since $\mathrm{mod}(k,q)>0$, there exist nonnegative integers $t$ and $s<q$ such that $k=tq+s$. Summing Eq. 29 for all $k=tq+1,\ldots,tq+s-1,tq+s$, and applying Eq. 28, we have
(31) $r_{g}^{(k)}\leq r_{g}^{(tq)}+\sum_{i=1}^{s}\frac{\alpha_{tq+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}.$
The proof is then completed by combining the above two cases.
The following corollary is obtained immediately from Eq. 31 by noticing $r_{g}^{(tq)}=0$ if $\mathbb{P}_{g}$ is a uniform distribution over $N$ data points and ${\mathcal{B}}_{tq}$ contains all the data points.
Corollary 3.3.
Under 1 and 4, suppose that $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$. Let ${\mathcal{B}}_{k}=\Xi$ if $\mathrm{mod}(k,q)=0$. Then $r_{g}^{(k)}=0$ if $\mathrm{mod}(k,q)=0$ and $r_{g}^{(k)}\leq\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$ otherwise, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
Next, we bound the difference between the Moreau envelope at two consecutive iterates.
Lemma 3.4.
Under 1 and 4, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1. With $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$, it holds that
(32) $\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k+1)})\right]$
$\leq$ $\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right]-\frac{\alpha_{k}}{{4}}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]+\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)}+{\alpha_{k}^{2}{\rho}L_{F}^{2}}{}.$
Proof 3.5.
For any $k\geq 0$, let $\widehat{{\mathbf{x}}}^{(k)}=\operatorname{Prox}_{\frac{F^{\nu}}{2\rho}}({{\mathbf{x}}}^{(k)})$. We deduce that
$\mathbb{E}_{k}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k+1)})\right]\leq\mathbb{E}_{k}\left[F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})+\rho\left\|\widehat{{\mathbf{x}}}^{(k)}-{\mathbf{x}}^{(k+1)}\right\|^{2}\right]$
[…]
$\quad\quad\quad\quad+\rho\mathbb{E}_{k}\left[\left\|{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k+1)}\right\|^{2}\right]$
$=$ $F^{\nu}_{\frac{1}{2\rho}}(\widehat{{\mathbf{x}}}^{(k)})+2\rho\alpha_{k}\mathbb{E}_{k}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right\rangle\right]+\rho\mathbb{E}_{k}\left[\left\|{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k+1)}\right\|^{2}\right]$
$\leq$ $F^{\nu}_{\frac{1}{2\rho}}(\widehat{{\mathbf{x}}}^{(k)})+2\rho\alpha_{k}\left(F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})-F^{\nu}({{\mathbf{x}}}^{(k)})+\frac{\rho}{2}\left\|{{\mathbf{x}}}^{(k)}-\widehat{{\mathbf{x}}}^{(k)}\right\|^{2}\right)$
(33) $\quad\quad\quad\quad+2\rho\alpha_{k}\mathbb{E}_{k}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]+\rho\mathbb{E}_{k}\left[\left\|{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k+1)}\right\|^{2}\right],$
where the first inequality follows directly from the definition of $F^{\nu}_{\frac{1}{2\rho}}$, and the last one uses $\mathbb{E}_{k}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$ by 4(b) and the $\rho$-weak convexity of $F^{\nu}$ by Lemma 2.6.
Since the function ${\mathbf{x}}\mapsto F^{\nu}({\mathbf{x}})+\rho\left\|{\mathbf{x}}-{{\mathbf{x}}}^{(k)}\right\|^{2}$ is $\rho$-strongly convex, we have
(34) $F^{\nu}({{\mathbf{x}}}^{(k)})-F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})-\frac{\rho}{2}\left\|{{\mathbf{x}}}^{(k)}-\widehat{{\mathbf{x}}}^{(k)}\right\|^{2}$
$=$ $\left(F^{\nu}({{\mathbf{x}}}^{(k)})+\rho\|{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|^{2}\right)-\left(F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})+\rho\|{{\mathbf{x}}}^{(k)}-\widehat{{\mathbf{x}}}^{(k)}\|^{2}\right)+\frac{\rho}{2}\left\|{{\mathbf{x}}}^{(k)}-\widehat{{\mathbf{x}}}^{(k)}\right\|^{2}$
[…]
Substituting Eq. 34 into Eq. 33, and using the law of expectation and Eq. 30, we obtain
(35) $\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k+1)})\right]$ $\leq\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right]-\mathbb{E}\left[\frac{\alpha_{k}}{2}\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]$
$\quad+\mathbb{E}\left[2\rho\alpha_{k}\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]+{\alpha_{k}^{2}\rho L_{F}^{2}}{}.$
For the third term in the right hand side (RHS) of Eq. 35, it holds that
$\textstyle 2\rho\alpha_{k}\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]\leq\mathbb{E}\left[{\alpha_{k}\rho^{2}}\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|^{2}\right]+\mathbb{E}\left[{\alpha_{k}\beta^{2}}{}\|{\mathbf{w}}^{(k)}\|^{2}\right]$
$=$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+{\alpha_{k}\beta^{2}}{}\mathbb{E}\left[\left\|\left(\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}\right]$
$\leq$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\left|\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right|^{2}\right]$
$\leq$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\left|\frac{u^{(k)}}{\nu}-\frac{g({\mathbf{x}}^{(k)})}{\nu}\right|^{2}\right]$
$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)},$
where the first equality follows by Eq. 14 and Eq. 26, the second inequality comes from 1(c) and 4(b), and the third inequality uses 1-Lipschitz of Proj operator. Substituting the above inequality back into Eq. 35 yields Eq. 32.
We now give the convergence result in the following theorem.
Theorem 3.6.
Under 1 and 4, given $\overline{\epsilon}>0$, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1 with
(36) $S_{1}=\lceil 16\overline{\epsilon}^{-2}\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\beta^{2}\sigma^{2}\rceil,\,q=S_{2}=\lceil\sqrt{S_{1}}\rceil,\,$
(37) $\alpha_{k}=\alpha=\min\left\{\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta\sqrt{l_{g}^{2}+\sigma_{2}^{2}}L_{g}L_{F}}\right\},\text{ and }K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil.$
Here, $\Delta:=\mathbb{E}[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})]-\min_{{\mathbf{x}}}\mathbb{E}[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}})]$ and $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$. It holds
(38) $\frac{1}{Kq}\sum_{k=0}^{Kq-1}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq\overline{\epsilon}^{2}.$
Proof 3.7.
Since $\alpha_{k}=\alpha$Lemma 3.1 implies
$\sum_{k=0}^{Kq-1}\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)}\leq\frac{Kq\alpha\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right).$
Summing Eq. 32 from $k=0$ to $Kq-1$ and applying the above bound, we derive
$\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(Kq)})\right]\leq$ $\,\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})\right]-\frac{\alpha}{{4}}\sum_{k=0}^{Kq-1}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]$
(39) $+\frac{Kq\alpha\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+Kq{\alpha^{2}{\rho}L_{F}^{2}}{}.$
Rearranging the terms and multiplying both sides of Eq. 39 by $\frac{4}{Kq\alpha}$, we obtain
$\frac{1}{Kq}\sum_{k=0}^{Kq-1}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq$ $\frac{4}{Kq\alpha}\left(\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})\right]-\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(Kq)})\right]\right)$
$\quad\quad\quad+\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}$
(40) $\leq$ $\frac{4\Delta}{Kq\alpha}+\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}.$
Now we bound the RHS of Eq. 40. From $\alpha\leq\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}}$, it follows that $4{\alpha{\rho}L_{F}^{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$. The choice $S_{1}=\lceil 16\overline{\epsilon}^{-2}\beta^{2}\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\sigma^{2}\rceil$ yields $\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\sigma^{2}}{S_{1}}\leq\frac{\overline{\epsilon}^{2}}{4}$. The conditions $q=S_{2}=\lceil\sqrt{S_{1}}\rceil$ and $\alpha\leq\frac{\overline{\epsilon}\nu}{4\beta\sqrt{l_{g}^{2}+\sigma_{2}^{2}}L_{g}L_{F}}$ imply $\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$. Also, we obtain $\frac{4\Delta}{Kq\alpha}\leq\frac{\overline{\epsilon}^{2}}{4}$ from $K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil$. Thus we have the desired result.
Under the conditions of Corollary 3.3, we have Eq. 40 with the vanishment of the term $\frac{\sigma^{2}}{S_{1}}$. Thus the following theorem can be established immediately.
Theorem 3.8.
Under the conditions of Corollary 3.3, given $\overline{\epsilon}>0$, let $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1 with $S_{1}=N$, (i.e., taking the whole data), and other parameters set to those in Eq. 36Eq. 37. Then Eq. 38 holds.
From Theorem 3.6 and Theorem 3.8, we can directly derive the complexity of Algorithm 1 to produce a near-$\overline{\epsilon}$ stationary solution of Eq. PP in expectation by choosing an iterate from $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ uniformly at random.
Corollary 3.9.
Under 1 and 4Algorithm 1 can produce a near-$\overline{\epsilon}$ stationary solution of Eq. PP in expectation with $Kq=O\left(\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic subgradients and $K(S_{1}+(q-1)S_{2})=O\left(\overline{\epsilon}^{-5}\nu^{-1}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic function evaluations on $g$. If $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, the number of stochastic function evaluations becomes $O\left(N+\sqrt{N}\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$.
3.3 Identifying an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation
In this subsection, we demonstrate that Algorithm 1 can find an $(\epsilon,\epsilon)$-KKT point for problem Eq. P under certain specific scenarios. The following proposition addresses the case with a convex constraint or a special weakly-convex case.
Proposition 3.10.
[…]
  • $\mathrm{(i)}$
    $O(\epsilon^{-4})$ OGC/CGC and $O(\epsilon^{-6})$ CFC if $\mathbb{P}_{g}$ is a general distribution;
  • $\mathrm{(ii)}$
    $O(\epsilon^{-4})$ OGC/CGC and $O(N+\sqrt{N}\epsilon^{-4})$ CFC if $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
Proof 3.11.
Choose $\overline{\epsilon}$, $\beta$, and $\nu$ so that the conditions in Eq. 17 hold and $\beta>~{}\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-\rho_{g}D^{2}}$. In addition, set $S_{1}$ as in Eq. 36 if $\mathbb{P}_{g}$ is a general distribution and $S_{1}=N$ , (i.e., taking the whole data), if $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$. Then set $S_{2},$ $q$, $\alpha_{k}$ and $K$ to those in Eq. 36 and Eq. 37. The claims now directly follow from Theorem 2.8(i)-(ii) and Corollary 3.9.
Remark 3.12.
If the strong feasibility condition assumed by [9] holds, i.e.,
(41) $\exists\,{\mathbf{y}}\in{{\mathrm{dom}}(g)}\text{ such that }g({\mathbf{y}})\leq-2\rho_{g}D^{2},$
we then have 2 with any (big) $B_{g}>0$ when $g$ is convex and 3 with any (big) $B_{g}>0$ when $g$ is weakly convex. Hence, the conditions in the two cases (i)-(ii) of Proposition 3.10 are satisfied.
Next, we consider a finite-sum weakly convex constraint $g$. For this case, we can obtain an ($\epsilon,\epsilon$)-KKT point of problem Eq. P with a weaker assumption than Eq. 41. Specifically, we require the (uniform) Slater-type CQ assumption that is assumed in [18, 30]. The purpose of making this assumption is to ensure $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all $k\geq 0$ so that Theorem 2.8(iii) can be applied.
Assumption 5 ((uniform) Slater-type CQ).
There exist $B_{g}>~{}0$, $B>0$, and $\overline{\rho}>\rho_{g}\geq 0$, such that for any ${\mathbf{x}}$ satisfying $g({\mathbf{x}})\leq B_{g}$, it holds $g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for some $\mathbf{y}\in{\mathrm{dom}}(g)$.
Remark 3.13.
When $g$ is convex, [18] assumes 1(a) and the existence of ${\mathbf{y}}_{\mathrm{feas}}$ such that $g({\mathbf{y}}_{\mathrm{feas}})<0$. These assumptions naturally lead to 5, with the parameters given by ${\mathbf{y}}={\mathbf{y}}_{\mathrm{feas}}$, $\overline{\rho}=\frac{-g({\mathbf{y}}_{\mathrm{feas}})}{D^{2}}$, $B=\frac{-g({\mathbf{y}}_{\mathrm{feas}})}{2}$, and any (big) $B_{g}>0$. When $g$ is convex, 5 implies 2 with the same value of $B_{g}$, and when $g$ is $\rho_{g}$-weakly convex with $\rho_{g}\geq 0$, it also implies 3 with the same value of $B_{g}$. We will verify these two assertions in Appendix A. Moreover, 5 is clearly implied by the strong feasibility condition Eq. 41.
Before giving our complexity result in this case, we establish a lower bound relating $g_{+}({\mathbf{x}})$ to the distance from ${\mathbf{x}}$ to the feasible set $\Omega$ in the following lemma. Its proof is given in Lemma A.5.
Lemma 3.14.
Suppose that $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, and 1 and 5 hold. If $\textnormal{dist}({\mathbf{x}},\Omega)\leq\min\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\}$, then $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\gamma=\sqrt{B(\overline{\rho}-\rho_{g})/2}$.
In the next lemma, we establish that Algorithm 1 guarantees $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all iterations.
Lemma 3.15.
Suppose that $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, and 14, and 5 hold. In Algorithm 1, let $u^{(k)}=g({\mathbf{x}}^{(k)})$, i.e., ${\mathcal{B}}_{k}=\Xi$ for all $k\geq 0$ in Eq. 25, and choose $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$. Suppose $\|\boldsymbol{\zeta}_{f}^{(k)}\|\leq l_{f}$ for all $k\geq 0$. Set
[…]
(43) $0<\alpha_{k}\leq\min\left\{\frac{C_{0}}{\beta l_{g}^{2}},{\frac{1}{\beta\rho_{g}},}\frac{\gamma}{2l_{f}\rho_{g}},\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}},\,\frac{B_{g}}{4l_{f}l_{g}}\right\},$
where $\gamma=\sqrt{B(\overline{\rho}-\rho_{g})/2}$. If the initial point ${\mathbf{x}}^{(0)}$ satisfies $\textnormal{dist}({\mathbf{x}}^{(0)},\Omega)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$, then it holds $\textnormal{dist}({\mathbf{x}}^{(k)},\Omega)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ and $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all $k\geq 0$.
Proof 3.16.
We prove the desired results by induction on $k$. The claim holds for $k=0$, by the assumption on ${\mathbf{x}}^{(0)}$. Suppose it holds up to iteration $k$. We show that the same bounds hold for $k+1$. We consider the two cases: (i) $g({\mathbf{x}}^{(k)})\geq~{}C_{0}$; (ii) $g({\mathbf{x}}^{(k)})<~{}C_{0}$.
(i) In this case, since $\nu\leq C_{0}\leq g({\mathbf{x}}^{(k)})$, the updating rule from Eq. 1 becomes
${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\boldsymbol{\zeta}_{g}^{(k)}\right)=\left({\mathbf{x}}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}\right)-\alpha_{k}\boldsymbol{\zeta}_{f}^{(k)}.$
Since $0<g({\mathbf{x}}^{(k)})\leq l_{g}\textnormal{dist}\left(\mathbf{x}^{(k)},\Omega\right)<B_{g}$, it follows from Lemma A.1 that $\boldsymbol{\zeta}_{g}^{(k)}\neq\mathbf{0}$. For simplicity of notation, let ${\mathbf{y}}^{(k)}={\mathbf{x}}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}$ and $\mathbf{x}^{\dagger(k)}=\operatorname{Proj}_{\Omega}(\mathbf{x}^{(k)})$. It holds
$\small{\textnormal{dist}^{2}({{\mathbf{y}}}^{(k)},\Omega)\leq\left\|{{\mathbf{y}}}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}=\left\|\mathbf{x}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}}$
$=\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)-2\alpha_{k}\beta\left\langle\boldsymbol{\zeta}_{g}^{(k)},\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\rangle+(\alpha_{k}\beta)^{2}\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}}$
$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+2\alpha_{k}\beta\left(g(\mathbf{x}^{\dagger(k)})-g(\mathbf{x}^{(k)})+\frac{\rho_{g}}{2}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)+(\alpha_{k}\beta)^{2}\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}}$
$=\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-g_{+}(\mathbf{x}^{(k)})+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)+\alpha_{k}\beta\left(\alpha_{k}\beta\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}-g(\mathbf{x}^{(k)})\right)}$
$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-g_{+}(\mathbf{x}^{(k)})+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)}$
$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-{\gamma}{}+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|\right)\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|}$
(44) $\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)-\alpha_{k}\beta\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2},}$
where the third equality comes from $g(\mathbf{x}^{\dagger(k)})=0$ and $g(\mathbf{x}^{(k)})=g_{+}(\mathbf{x}^{(k)})$, the second inequality holds by the $\rho_{g}$-weak convexity of $g$, the third one uses $g(\mathbf{x}^{(k)})\geq C_{0}$, $\|\boldsymbol{\zeta}_{g}^{(k)}\|\leq l_{g}$, and $0<\alpha_{k}\beta\leq\frac{C_{0}}{l_{g}^{2}}$, the fourth one follows by Lemma 3.14, and the last one holds because $-{\gamma}{}\leq-2\rho_{g}\textnormal{dist}({\mathbf{x}}^{(k)},\Omega)=-2\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|$ by induction. Then it follows that
$\vspace{-1mm}\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)$ $\leq\textnormal{dist}({{\mathbf{y}}}^{(k)},\Omega)+\alpha_{k}l_{f}\leq\sqrt{1-\alpha_{k}\beta\rho_{g}}\textnormal{dist}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}l_{f}$
$\leq\textnormal{dist}(\mathbf{x}^{(k)},\Omega),$
where the second inequality uses Eq. 44, and the last one follows from $\beta\geq\frac{2l_{f}l_{g}}{\rho_{g}C_{0}}$ and $\textnormal{dist}\left(\mathbf{x}^{(k)},\Omega\right)\geq\frac{g({\mathbf{x}}^{(k)})}{l_{g}}\geq~{}\frac{C_{0}}{l_{g}}$. Thus $\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ by induction, and $g({{\mathbf{x}}}^{(k+1)})\leq l_{g}\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\frac{B_{g}}{2}$.
(ii) In the second case, we have $g({\mathbf{x}}^{(k)})\leq C_{0}.$ From Eq. 1, we have
$\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)\leq$ $\textnormal{dist}({{\mathbf{x}}}^{(k)},\Omega)+\alpha_{k}(l_{f}+\beta l_{g})$
(45) $\leq$ $\min\left\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\right\}+\alpha_{k}l_{f}+\frac{C_{0}}{l_{g}}\leq\min\left\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\right\},$
where the second inequality follows from $\alpha_{k}\leq\frac{C_{0}}{\beta l_{g}^{2}}$ and $\textnormal{dist}\left({{\mathbf{x}}}^{(k)},\Omega\right)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ by induction hypothesis, and the last one is derived using $\alpha_{k}\leq\min\{\frac{\gamma}{2l_{f}\rho_{g}},\frac{B_{g}}{4l_{f}l_{g}}\}$, and $C_{0}\leq\min\{\frac{\gamma l_{g}}{\rho_{g}},\frac{B_{g}}{4}\}$. We then obtain
$\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)\leq\frac{g({\mathbf{x}}^{(k+1)})}{\gamma}\leq\frac{1}{\gamma}\left(g({\mathbf{x}}^{(k)})+l_{g}\|{{\mathbf{x}}}^{(k+1)}-{{\mathbf{x}}}^{(k)}\|\right)$
(46) $\leq$ $\frac{1}{\gamma}\left(g({\mathbf{x}}^{(k)})+\alpha_{k}l_{g}(l_{f}+\beta l_{g})\right)\leq\frac{C_{0}}{\gamma}+\frac{1}{\gamma}\left(\alpha_{k}l_{g}l_{f}+{C_{0}}{}\right)\leq\min\left\{\frac{\gamma}{2\rho_{g}},\,\frac{B_{g}}{2l_{g}}\right\},$
where the first inequality uses Lemma 3.14 by Eq. 45, the second/third one comes from the updating rule Eq. 1 and 1(c), the fourth one holds because $g({\mathbf{x}}^{(k)})\leq C_{0}$ and $\alpha_{k}\leq\frac{C_{0}}{\beta l_{g}^{2}}$, and the last one follows by $C_{0}\,\leq\,\min\{\frac{\gamma^{2}}{8\rho_{g}},\,\frac{\gamma B_{g}}{8l_{g}}\},$ and $\alpha_{k}\,\leq\,\min\{\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}}\}.$ Hence $g({{\mathbf{x}}}^{(k+1)})\leq l_{g}\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\frac{B_{g}}{2}$.
Finally, combining the two above cases and applying the induction hypothesis, we complete the proof.
Now we are ready to show the complexity result for the finite-sum weakly convex case under 5.
Proposition 3.17.
Suppose $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, and 14, and 5 hold. In Algorithm 1, let $u^{(k)}=g({\mathbf{x}}^{(k)})$, i.e., ${\mathcal{B}}_{k}=\Xi$ for all $k\geq 0$ in Eq. 25, and choose $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$. Suppose $\|\boldsymbol{\zeta}_{f}^{(k)}\|\leq l_{f}$ for all $k\geq 0$. Then given $\epsilon>0$Algorithm 1 can produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation with $O(\epsilon^{-4})$ stochastic subgradients of $f$, and $O(\epsilon^{-4})$ deterministic subgradients and function evaluations of $g$.
Proof 3.18.
Choose $\overline{\epsilon}$ and $\beta$ to satisfy the conditions in Eq. 17 and Eq. 42 and $\beta>\frac{2B_{f}}{B_{g}}$, and let $\nu$ satisfy $0<\nu\leq\min\{C_{0},{B}_{g},\frac{\epsilon}{\beta},\frac{\epsilon}{2}\},$ where $C_{0}$ is given in Lemma 3.15. Also, choose the stepsize as
$\alpha_{k}=\alpha=\min\left\{\frac{C_{0}}{\beta l_{g}^{2}},{\frac{1}{\beta\rho_{g}},}\frac{\gamma}{2l_{f}\rho_{g}},\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}},\,\frac{B_{g}}{4l_{f}l_{g}},\,\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta l_{g}L_{g}L_{F}}\right\}.\vspace{-2mm}$
Since $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$ and $u^{(k)}=g({\mathbf{x}}^{(k)})$, it holds $\sigma=\sigma_{2}=0$. Hence, we can still obtain Eq. 38 from Eq. 40 by setting $q=1$, $S_{2}=N$, (i.e., taking the whole data), and $K=\lceil 16\overline{\epsilon}^{-2}\alpha^{-1}\Delta\rceil$, thus a near-$\overline{\epsilon}$ stationary point ${\mathbf{x}}^{*}$ of Eq. PP in expectation can be produced within $K$ iterations. Now by Theorem 2.8(iii) and Lemma 3.15, ${\mathbf{x}}^{*}$ is also an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation. It takes $K=O(\epsilon^{-4})$ stochastic subgradient of $f$ and $K=O(\epsilon^{-4})$ deterministic subgradients and function evaluations of $g$. The proof is then completed.
4 Numerical experiments
In this section, we evaluate the empirical performance of the proposed Algorithm 1 on two fairness-constrained problems, which are instances of Eq. P. We compare its performance with three state-of-the-art approaches, including two different double-loop IPP [9, 30] and the SSG in [18]. All the experiments are performed on iMac with Apple M1, a 3.2 GHz 8-core processor, and 16GB of RAM running MATLAB R2024a.
4.1 Datasets and tested problems
The experiments are conducted on two datasets: a9a [20], and COMPAS [5]. The a9a dataset contains 48,842 data points with 123 features, and is used to predict income levels; gender is the protected attribute, and fairness is assessed between males and females. The COMPAS dataset includes 6,172 data points and 16 features, focused on predicting recidivism risk, with fairness evaluated between Caucasian and non-Caucasian groups. We split each dataset into two subsets with a ratio of 2:1. The larger set is used as $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, and the smaller set is further split into a protected set $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and an unprotected set $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$.
Problem 1. The first tested problem is a classification problem with ROC-based fairness [37]. It is formulated in Eq. 4. For this problem, we have $\rho_{g}=0$ and $\rho_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}$. That is, $\Phi(\cdot)-\Phi^{*}-\kappa_{1}$ is convex. We obtain the minimum $\Phi^{*}$ and a solution ${\mathbf{x}}^{*}$ by applying a stochastic subgradient method to a large number of iterations on $\min_{{\mathbf{x}}}\Phi({\mathbf{x}})$. Afterward, we set $\kappa_{1}=0.001\cdot\Phi^{*}$, $D=5\|{\mathbf{x}}^{*}\|$, and let $\Theta$ consist of 400 points equally spaced between $\uline{{\mathbf{a}}}-0.5\left(\overline{{\mathbf{a}}}-\uline{{\mathbf{a}}}\right)$ and $\overline{{\mathbf{a}}}+0.5\left(\overline{{\mathbf{a}}}-\uline{{\mathbf{a}}}\right)$, where $\overline{{\mathbf{a}}}=\max_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$ and $\uline{{\mathbf{a}}}=\min_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$. For all the compared methods, we initialize ${\mathbf{x}}^{(0)}={\mathbf{x}}^{*}$.
Problem 2. The second problem is a classification problem with demographic parity [1] and the smoothly clipped absolute deviation (SCAD) regularization term [15]. It is formulated as
(47) $\min_{{\mathbf{x}}}\Phi({\mathbf{x}})+\lambda\|\varphi({\mathbf{x}})\|_{1},\,\,\text{{s.t.}}\,\,\Psi_{0}({\mathbf{x}})\leq\kappa_{2},$
where $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, $\Psi_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}$, $\Phi$ is the same as that in Eq. 4, $\lambda>0$, $\kappa_{2}>0$ is a slackness parameter. The functions $\Psi_{0}$ and $\varphi$ are defined as follows:
$\small{\Psi_{0}({\mathbf{x}}):=\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\sigma\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\sigma\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\right)\right|,\text{ and }}$
$\small{[\varphi({\mathbf{x}})]_{i}=\begin{cases}2\left|x_{i}\right|,&0\leq\left|x_{i}\right|\leq 1,\\ -x_{i}^{2}+4\left|x_{i}\right|+1,&1<\left|x_{i}\right|\leq 2,\\ 3,&\left|x_{i}\right|>2,\end{cases}\quad\text{ for }i=1,2,\ldots,d}$
where $\varphi$ promotes sparsity in ${\mathbf{x}}$. Here, $\Psi_{0}$ is a continuous relaxation of the demographic parity $\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\geq 0\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\geq 0\right)\right|$. For this problem, we have $\rho_{f}=\rho_{g}=\max\left\{2\lambda,\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}\right\}.$ In the experiments, we set $\lambda=0.02$, $\kappa_{2}=0.02$, and ${\mathbf{x}}^{(0)}=\mathbf{0}$ for all the compared methods.
4.2 Implementation details
For our method, we use $\beta=10$ and $\nu=10^{-5}$ as the default setting. We implement both deterministic and stochastic versions of our method (denoted as 3S-Econ-D and 3S-Econ-S, respectively). For 3S-Econ-D, we set $q=1$ and $S_{1}=S_{2}=n$ (i.e., taking the whole data) and select $\alpha_{k}=10^{-2}$ as our default setting. For 3S-Econ-S, we set $S_{1}=n$ and $S_{2}=q=\lceil\sqrt{n}\rceil$ and select $\alpha_{k}=\frac{1}{100\sqrt{k/q}}$ as our default setting [^2] To determine the default settings, we test 3S-Econ-D and 3S-Econ-S on Problem 1 using the a9a dataset, with varying choices of $\beta$ and $\alpha_{k}$. Specifically, for all $k\geq 0$, we choose $\beta\in\{1,10,100,1000\}$ and $\alpha_{k}\in\{0.001,0.01,\frac{1}{100\sqrt{k/q}},\frac{1}{1000\sqrt{k/q}}\}$. Through our experiments, we observe that 3S-Econ-D achieves the best performance with $\beta=10$ and $\alpha_{k}=0.01$. Similarly, for 3S-Econ-S, $\beta=10$ and $\alpha_{k}=\frac{1}{100\sqrt{k/q}}$ yield the best performance. Based on these observations, we fix these parameters as the default settings and apply them across all datasets and tested problems. .
We adopt the tuning techniques in [9, 30] for the two distinct double-loop IPP methods and the techniques in [18] for the SSG. A strongly convex constrained subproblem is solved approximately in each outer iteration of the IPPs. For this subproblem, we employ the SSG from [18] and the ConEx in [9] as inner solvers. We refer to the two implemented IPPs as IPP-SSG and IPP-ConEx, respectively. The stepsize $\alpha_{k}$ for IPP-SSG is selected from $\{2\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3}\}$. Following the notation in [30], we set $\theta_{t}=\frac{t}{t+1},\eta_{t}=c_{1}(t+1)$, and $\tau_{t}=\frac{c_{2}}{t+1}$ for IPP-ConEx, where $c_{1}$ is chosen from $\{20,50,100,200\}$ and $c_{2}$ from $\{0.002,0.005,0.01,0.02\}$. The optimal parameter set for these methods is selected based on the smallest objective value after 5000 iterations.
We report the objective function value (FV), i.e., $f({\mathbf{x}})$, the constraint violation (CVio), i.e., $[g({\mathbf{x}})]_{+}$, and the stationarity violation (SVio). At each iteration, we solve
$\min_{{\mathbf{x}}}\big{\{}f({\mathbf{x}})+\rho_{f}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2},\,\text{{s.t.}}\,g({\mathbf{x}})+\rho_{g}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2}\leq 0\big{\}}$
to a desired accuracy to obtain $\widehat{{\mathbf{x}}}^{(k)}$ and then use $\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|$ to measure SVio [^3] Specically, we solve the strongly convex problem by the SSG [18] to a tolerance of $10^{-2}$..
4.3 Comparisons among different methods
In Fig. 1, we compare the 3S-Econ-D method with three state-of-the-art deterministic approaches on solving Problems 1–2. It is clear that 3S-Econ-D outperforms the others in FVs and SVio values, not only converging faster but also achieving lower errors. In terms of CVio, 3S-Econ-D shows more fluctuations. Nevertheless, we observe that it remains feasible for most of the iterations. Since for the deterministic version, we can explicitly check the feasibility. Thus by keeping feasible iterates, 3S-Econ-D will significantly outperform the other three deterministic methods.
(a) Problem 1, a9a
(b) Problem 1, a9a
(c) Problem 1, a9a
(d) Problem 1, COMPAS
(e) Problem 1, COMPAS
(f) Problem 1, COMPAS
(g) Problem 2, a9a
(h) Problem 2, a9a
(i) Problem 2, a9a
(j) Problem 2, COMPAS
(k) Problem 2, COMPAS
(l) Problem 2, COMPAS
Figure 1: Comparisons between 3S-Econ-D (deterministic version of our method) and other approaches: SSG [18], IPP-SSG [30], and IPP-ConEx [9].
To evaluate the numerical behavior of the stochastic version of our method (denoted as 3S-Econ-S), we first compare it with 3S-Econ-D. The FV and CVio values on Problems 1-2 with two datasets are shown in Fig. 2, “data passes” for the $x$-axis label is a shorthand for “data passes ($g$)” (DP ($g$)), referring to the number of times to access all the data involved in the constraint function. Despite more oscillations due to the stochasticity, 3S-Econ-S can produce comparable results as 3S-Econ-D with significantly fewer data passes.
(a) Prob. 1, a9a
(b) Prob. 1, COMPAS
(c) Prob. 2, a9a
(d) Prob. 2, COMPAS
Figure 2: Comparisons between the deterministic and stochastic versions of our method
In Table 2, we show more comparison results of 3S-Econ-D and 3S-Econ-S with IPP-SSG, IPP-ConEx, SSG, and the stochastic version of SSG (denoted as SSG-S) [^4] It is worth noting that the stochastic versions of IPP-SSG and IPP-ConEx produce numerical results inferior to their deterministic counterparts when solving Problems 1–2. Thus, we only present results for their deterministic versions. This observation can be explained by the fact that, although their stochastic variants utilize the stochastic subgradient of $f$ rather than the full subgradient, both methods still require accessing all data in the constraints during each iteration to access a constraint stochastic subgradient and the constraint deterministic function value. . We terminate all the deterministic methods if SVio is less than $10^{-3}$, while for the stochastic algorithms, the threshold is set to $5\times 10^{-3}$. In addition, we impose a maximum “DP ($g$)” to 200,000 for Problem 1 and 720,000 for Problem 2. Since computing $\widehat{{\mathbf{x}}}^{(k)}$ with high precision at each iteration $k$ is time-consuming, we calculate SVio every few steps. Upon termination, in terms of “DP ($g$)” that dominate “DP ($f$)” for all methods, we observe that for Problem 1, 3S-Econ-D is 5 times faster than other deterministic methods on the a9a dataset and 2 times faster on the COMPAS dataset; for Problem 2, 3S-Econ-D is 2-15 times faster than other deterministic methods on the a9a dataset, and 1.5-2 times faster on the COMPAS dataset. Furthermore, 3S-Econ-S delivers even more impressive speedups: for Problem 1, it is 219 times faster than SSG-S on the a9a dataset and 108 times faster on the COMPAS dataset; for Problem 2, it is 466 times faster than SSG-S on the a9a dataset and 26 times faster on the COMPAS dataset. Additionally, in most cases, our methods demonstrate superior performance in terms of both FV and SVio when compared to other algorithms.
measures SSG IPP-SSG IPP-ConEx 3S-Econ-D SSG-S 3S-Econ-S
Problem 1 a9a iteration 1.00e+05 1.00e+05 1.00e+05 1.50e+04 1.99e+05 8.24e+04
DP ($f$) 1.00e+05 1.00e+05 1.00e+05 1.50e+04 1.10e+03 455
DP ($g$) 2.00e+05 2.00e+05 2.00e+05 3.00e+04 2.00e+05 910
FV 9.26e-02 1.03e-01 9.70e-02 8.47e-02 8.08e-02 8.48e-02
CVio 3.76e-05 6.57e-06 0 1.41e-04 6.34e-04 0
SVio 2.41e-03 5.25e-03 3.60e-03 9.56e-04 1.36e-02 4.69e-03
COMPAS iteration 1.00e+05 1.00e+05 1.00e+05 3.70e+04 1.96e+05 6.00e+04
DP ($f$) 1.00e+05 1.00e+05 1.00e+05 3.70e+04 3.06e+03 938
DP ($g$) 2.00e+05 2.00e+05 2.00e+05 7.40e+04 2.00e+05 1.85e+03
FV 6.91e-03 6.79e-03 6.85e-03 6.31e-03 6.49e-03 6.41e-03
CVio 9.27e-06 5.65e-06 2.17e-06 1.49e-05 5.66e-04 0
SVio 1.65e-03 1.52e-03 1.60e-03 9.99e-04 1.43e-02 4.77e-03
Problem 2 a9a iteration 4.02e+04 3.60e+05 3.60e+05 2.10e+04 5.10e+04 9.90e+03
DP ($f$) 4.02e+04 3.60e+05 3.60e+05 2.10e+04 282 55
DP ($g$) 8.04e+04 7.20e+05 7.20e+05 4.20e+04 5.13e+04 110
FV 5.11e-01 5.07e-01 5.14e-01 5.05e-01 5.07e-01 5.10e-01
CVio 5.11e-05 0 0 0 0 0
SVio 9.58e-04 2.72e-03 1.15e-02 8.58e-04 1.99e-03 2.19e-03
COMPAS iteration 2.89e+05 3.60e+05 3.10e+05 1.54e+05 1.15e+05 1.39e+05
DP ($f$) 2.89e+05 3.60e+05 3.10e+05 1.54e+05 1.80e+03 2.17e+03
DP ($g$) 5.78e+05 7.20e+05 6.20e+05 3.07e+05 1.17e+05 4.35e+03
FV 2.04e-01 2.04e-01 2.04e-01 2.04e-01 2.06e-01 2.06e-01
CVio 0 0 0 0 1.24e-05 0
SVio 8.54e-04 8.30e-04 7.99e-04 9.43e-04 4.06e-03 4.70e-03
Table 2: Comparisons among all the methods.
5 Conclusions
We have presented a novel single-loop first-order method for solving nonconvex nonsmooth stochastic optimization with expectation constraints. Built on an exact penalty formulation that applies a smoothed penalty function on the inequality expectation constraint, our method incorporates the SPIDER-type estimation of the constraint function into the stochastic subgradient method. Through establishing the equivalence between the near-stationary solution of the penalty problem and the near-KKT solution of the original problem, we achieve an optimal $O(\epsilon^{-4})$ iteration complexity result to produce an $(\epsilon,\epsilon)$-KKT point. In terms of sample complexity, our result is lower than a few state-of-the-art results by a factor of $\epsilon^{-2}$ in either constraint sample subgradient or constraint function values, or by a factor of $\sqrt{N}$ when the constraint is the average of $N$ sample functions. Moreover, our method delivers significantly superior numerical performance over state-of-the-art methods on solving fairness-constrained problems.
Acknowledgement
We would like to express our sincere gratitude to Dr. Yankun Huang and Prof. Qihang Lin for providing their code in [18].
Appendix A Comparisons among our CQs and others
As mentioned in Remark 3.13, the strong feasibility condition is stronger than the (uniform) Slater-type CQ stated in 5. In the following lemmas, we demonstrate that 2 and 3 can be implied by the uniform Slate-type CQ under mild conditions.
Lemma A.1.
Under 1 and 5, if $0\leq g({\mathbf{x}})\leq B_{g}$, then $\min_{\boldsymbol{{\boldsymbol{\zeta}}_{g}}\in\partial g(\mathbf{x})}\left\|\boldsymbol{{\boldsymbol{\zeta}}_{g}}\right\|\geq\theta,$ where $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$.
Proof A.2.
Becasue $g({\mathbf{x}})\leq B_{g}$ and 5, There exists a point $\mathbf{y}$ such that $g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$. Since $g(\mathbf{z})+\frac{\overline{\rho}}{2}\|\mathbf{z}-\mathbf{x}\|^{2}$ is $(\overline{\rho}-\rho_{g})$-strongly convex, we obtain that for any $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$, it holds
(48) $g(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}+\left\langle\boldsymbol{\zeta}_{g},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B.$
Applying Young’s inequality to Eq. 48 and noticing $g({\mathbf{x}})\geq 0$, we derive
$-\frac{\left\|\boldsymbol{\zeta}_{g}\right\|^{2}}{2(\overline{\rho}-\rho_{g})}\leq\left\langle\boldsymbol{\zeta}_{g},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B-g({\mathbf{x}})\leq-B.$
This indicates $\|\boldsymbol{\zeta}_{g}\|\geq\sqrt{2B(\overline{\rho}-\rho_{g})}$ and completes the proof.
Lemma A.3.
Suppose that 1 and 5 hold and $g$ is convex. Then for any ${\mathbf{x}}$ satisfying $g({\mathbf{x}})\leq B_{g}$, it holds that $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\theta$ is given in Lemma A.1.
Proof A.4.
Define $\mathbf{x}^{+}=\operatorname{Proj}_{\Omega}(\mathbf{x})\in\underset{{\mathbf{y}}:g(\mathbf{y})\leq 0}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}.$ If $\mathbf{x}\in\Omega$, the result holds trivially. Therefore, we assume $\mathbf{x}\not\in\Omega$, which implies $g\left(\mathbf{x}^{+}\right)=0$. Under the Slater-type CQ, there exist $\lambda^{+}\geq 0$ and $\boldsymbol{\zeta}_{g}\in\partial g\left(\mathbf{x}^{+}\right)$ such that $\mathbf{x}^{+}-\mathbf{x}+\lambda^{+}\boldsymbol{\zeta}_{g}=\mathbf{0}.$ We assert that $\lambda^{+}>0$, because otherwise, the above inclusion would imply $\mathbf{x}=\mathbf{x}^{+}$, which contradicts the assumption that $\mathbf{x}\not\in\Omega$. By the convexity of $g$, we have
(49) $\lambda^{+}\left(g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\right)\geq\lambda^{+}\left\langle\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle=\left\langle\lambda^{+}\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle=\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}.$
Dividing both sides of the inequalities by $\lambda^{+}$ in Eq. 49 yields
$g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\geq\frac{\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}}{\lambda^{+}}=\|\boldsymbol{\zeta}_{g}\|\|\mathbf{x}-\mathbf{x}^{+}\|\geq\sqrt{2B(\overline{\rho}-\rho_{g})}\left\|\mathbf{x}-\mathbf{x}^{+}\right\|,$
where the last inequality is by Lemma A.1. Noting that $g(\mathbf{x})=g_{+}(\mathbf{x})$ for $\mathbf{x}\in\Omega$ and $g\left(\mathbf{x}^{+}\right)=0$, we obtain $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, and complete the proof.
Finally, we remove the convexity assumption from the last lemma and present the following result.
Lemma A.5.
Suppose that 1 and 5 hold, and $\rho_{g}>0$. Then for any ${\mathbf{x}}$ satisfying $\textnormal{dist}({\mathbf{x}},\Omega)\leq\min\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\}$, it holds that $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\gamma=\sqrt{\frac{B(\overline{\rho}-\rho_{g})}{2}}$.
Proof A.6.
Define $\mathbf{x}^{+}=\operatorname{Proj}_{\Omega}(\mathbf{x})\in\underset{{\mathbf{y}}:g(\mathbf{y})\leq 0}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}.$ If $\mathbf{x}\in\Omega$, the result holds trivially. Therefore, we assume $\mathbf{x}\not\in\Omega$, which implies $g\left(\mathbf{x}^{+}\right)=0$. It is Straightforward that $\mathbf{x}^{+}$ is also the optimal solution of the convex problem $\underset{{\mathbf{y}}:g(\mathbf{y})+\frac{\overline{\rho}}{2}\|{\mathbf{y}}-\mathbf{x}^{+}\|^{2}\leq 0}{\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$. By 5, the Slater’s condition holds for the above problem. Hence, there exist $\lambda^{+}\geq 0$ and $\boldsymbol{\zeta}_{g}\in\partial g\left(\mathbf{x}^{+}\right)$ such that $\mathbf{x}^{+}-\mathbf{x}+\lambda^{+}\boldsymbol{\zeta}_{g}=\mathbf{0}.$ We assert that $\lambda^{+}>0$, as otherwise, the above inclusion would imply $\mathbf{x}=\mathbf{x}^{+}$, which contradicts to $\mathbf{x}\not\in\Omega$. By the $\rho_{g}$-weak convexity of $g$, we have
(50) $\lambda^{+}\left(g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\right)\geq\lambda^{+}\left(\left\langle\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}\right)$
$=$ $\left\langle\lambda^{+}\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle-\frac{\lambda^{+}\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}=\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}-\frac{\lambda^{+}\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}.$
Since $0<g({\mathbf{x}})\leq l_{g}\textnormal{dist}\left(\mathbf{x},\Omega\right)\leq B_{g}$, dividing both sides of Eq. 50 by $\lambda^{+}$ yields
$g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)$ $\geq\frac{\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}}{\lambda^{+}}-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}=\|\boldsymbol{\zeta}_{g}\|\|\mathbf{x}-\mathbf{x}^{+}\|-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}$
$\geq\sqrt{2B(\overline{\rho}-\rho_{g})}\left\|\mathbf{x}-\mathbf{x}^{+}\right\|-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2},$
where the last inequality follows from Lemma A.1. Noting $g(\mathbf{x})=g_{+}(\mathbf{x})$ for $\mathbf{x}\not\in\Omega$, $g\left(\mathbf{x}^{+}\right)=0$ and $\|\mathbf{x}-\mathbf{x}^{+}\|=\textnormal{dist}({\mathbf{x}},\Omega)\leq\frac{\sqrt{2B(\overline{\rho}-\rho_{g})}}{\rho_{g}}$, we obtain $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$ and complete the proof.

### v2

@@
-A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization
+A SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization
 Abstract
 Many real-world problems, such as those with fairness constraints, involve complex expectation constraints and large datasets, necessitating the design of efficient stochastic methods to solve them.
-Most existing research focuses on cases with no constraint or easy-to-project constraints or deterministic constraints.
-In this paper, we consider nonconvex nonsmooth stochastic optimization problems with expectation constraints, for which we build a novel exact penalty model.
+Most existing research focuses on cases, which have no constraint, or have constraints that admit an easy projection, or only have deterministic constraints.
+In this paper, we consider weakly convex stochastic optimization problems with expectation constraints, for which we build an exact penalty model.
 We first show the relationship between the penalty model and the original problem.
-Then on solving the penalty problem, we present a single-loop SPIDER-type stochastic subgradient method, which utilizes the subgradients of both the objective and constraint functions, as well as the constraint function value at each iteration.
-Under certain regularity conditions (weaker than Slater-type constraint qualification or strong feasibility assumed in existing works), we establish an iteration complexity result of $O(\epsilon^{-4})$ to reach a near-$\epsilon$ stationary point of the penalized problem in expectation, matching the lower bound for such tasks.
+Then on solving the penalty problem, we present a SPIDER-type stochastic subgradient method, which utilizes the subgradients of both the objective and constraint functions, as well as the constraint function value at each iteration.
+Under the Slater-type constraint qualification (CQ), we establish an iteration complexity result of $O(\epsilon^{-4})$ to reach a near-$\epsilon$ stationary point of the penalized problem in expectation, matching the lower bound for such tasks.
 Building on the exact penalization, an $(\epsilon,\epsilon)$-KKT point of the original problem is obtained.
 For a few scenarios, our complexity of either the objective sample subgradient or the constraint sample function values can be lower than the state-of-the-art results in [10, 27, 46] by a factor of $\epsilon^{-2}$.
-Moreover, on solving two fairness-constrained problems, our method is significantly (up to 466 times) faster than the state-of-the-art algorithms, including switching subgradient method in [18] and inexact proximal point methods in [9, 30].
+Moreover, on solving two fairness-constrained problems and a multi-class Neyman-Pearson classification problem, our method is significantly (up to 466 times in terms of data pass) faster than the state-of-the-art algorithms, including the switching subgradient method in [27] and the inexact proximal point methods in [10, 46].
 keywords:
-stochastic, subgradient, expectation constraints, weakly convex, fairness constrained classification.
+stochastic, subgradient, expectation constraints, weakly convex, fairness-constrained classification.
 {AMS}
@@
 1 Introduction
-Stochastic optimization is fundamental to the statistical sciences and serves as the backbone of data-driven learning processes.
+Stochastic optimization is fundamental to the statistical learning and serves as the backbone of data-driven learning processes.
 In this paper, we study this topic by designing first-order methods (FOMs) for solving the expectation-constrained nonconvex programming:
-(P) $\min_{{\mathbf{x}}}\,\,{f({\mathbf{x}})},\text{ s.t. }\underbrace{\mathbb{E}_{\xi_{g}\sim\mathbb{P}_{g}}[g({\mathbf{x}},\xi_{g})]}_{=:g({\mathbf{x}})}\leq 0.$
-Here, $\mathbb{P}_{g}$ is an unknown probability distribution, and we will consider a general distribution and a special uniform distribution on a finite set, while $f$ is always assumed to be stochastic.
-Notice that a problem with multiple constraints $g_{i}({\mathbf{x}})\leq 0,i=1,\ldots,m$ can be formulated into Eq.
-P by setting $g({\mathbf{x}})\equiv\max_{i=1,\ldots,m}g_{i}({\mathbf{x}})$.
+(P) $\min_{{\mathbf{x}}\in\mathcal{X}}\,\,{f({\mathbf{x}})},\text{ s.t.
+}\underbrace{\mathbb{E}_{\xi_{{\mathbf{g}}}\sim\mathbb{P}_{{\mathbf{g}}}}[{\mathbf{g}}({\mathbf{x}},\xi_{{\mathbf{g}}})]}_{=:{\mathbf{g}}({\mathbf{x}})}\leq 0.$
+Here, $\mathcal{X}\subset\mathbb{R}^{d}$ is a closed convex set that allows a computationally easy projection operator, ${\mathbf{g}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is a vector-valued function consisting of $m$ components $g_{1},\ldots,g_{m}$, and $\mathbb{P}_{{\mathbf{g}}}$ is an unknown distribution from which we can draw independent samples.
+We will consider both general distributions and the special case where $\mathbb{P}_{{\mathbf{g}}}$ is uniform over a finite sample set; for the latter we can obtain deterministic subgradient of each $g_{i}$ if needed.
+For $f$, we assume only its stochastic subgradient is available.
 Substantial research has been devoted to FOMs for solving function-constrained optimization problems.
 Among these FOMs, most are developed for deterministic function-constrained problems or for expectation-constrained convex optimization.
-This is well-documented in the literature for convex scenarios [6, 7, 21, 22, 23, 26, 29, 40, 41, 44] and nonconvex scenarios [3, 4, 10, 14, 28, 27, 31].
+This is well-documented in the literature for convex scenarios [7, 8, 31, 32, 33, 40, 45, 59, 60, 64] and nonconvex scenarios [3, 4, 13, 19, 22, 25, 28, 42, 43, 47, 54].
 Only a few works (e.g., [2, 10, 27, 38]) have studied FOMs for expectation-constrained nonconvex optimization.
@@
 P.
-Throughout this paper, we assume the follows.
+Throughout this paper, we make the following assumption.
 Assumption 1.
 The following statements hold.
-$\mathrm{(a)}$ The set ${\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)\cap\Omega$ is nonempty, where $\Omega:=\{{\mathbf{x}}:g({\mathbf{x}})\leq 0\}$, and ${\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$ is bounded.
-Specifically, there exists $D>0$ such that $\|{\mathbf{x}}_{1}-{\mathbf{x}}_{2}\|\leq D$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
-$\mathrm{(b)}$ The function $f$ is $\rho_{f}$-weakly convex, and $g$ is $\rho_{g}$-weakly convex for some $\rho_{f}>0$ and $\rho_{g}\geq 0$.
-That is, $f+\frac{\rho_{f}}{2}\|\cdot\|^{2}$ and $g+\frac{\rho_{g}}{2}\|\cdot\|^{2}$ are convex.
-$\mathrm{(c)}$ The subdifferential sets of $f$ and $g$ are nonempty in their respective domains.
-In addition, it holds that $\|{\boldsymbol{\zeta}}_{f}\|\leq l_{f}$ and $\|{\boldsymbol{\zeta}}_{g}\|\leq l_{g}$ for all ${\boldsymbol{\zeta}}_{f}\in\partial f({\mathbf{x}})$ and ${\boldsymbol{\zeta}}_{g}\in\partial g({\mathbf{x}})$ for any ${\mathbf{x}}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
-From items (a) and (c) in the above assumption, we have $|f({\mathbf{x}}_{1})-f({\mathbf{x}}_{2})|\leq l_{f}D=:B_{f}$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
-In item (b), we allow $\rho_{g}=0$, which indicates the convexity of $g$.
-However, we always assume $\rho_{f}>0$ and thus consider a nonconvex objective function $f$.
+$\mathrm{(a)}$ The set $\mathcal{X}\cap\{{\mathbf{x}}:{\mathbf{g}}({\mathbf{x}})\leq 0\}$ is nonempty, and $\mathcal{X}$ is bounded.
+Specifically, there exists $D>0$ such that $\|{\mathbf{x}}_{1}-{\mathbf{x}}_{2}\|\leq D$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in\mathcal{X}$.
+$\mathrm{(b)}$ The function $f$ is $\rho_{f}$-weakly convex on $\mathcal{X}$}, and $g_{i}$ is $\rho_{g}$-weakly convex on $\mathcal{X}$} for some $\rho_{f}\geq 0$, $\rho_{g}\geq 0$ and each $i\in[m]:=\{1,\ldots,m\}$.
+That is, $f+\frac{\rho_{f}}{2}\|\cdot\|^{2}$ and $g_{i}+\frac{\rho_{g}}{2}\|\cdot\|^{2}$ are convex on $\mathcal{X}$} for each $i\in[m]$.
+$\mathrm{(c)}$ The subdifferential sets of $f$ and $g_{i}$ are nonempty in their respective domains for each $i\in[m]$.
+In addition, it holds that $\|{\boldsymbol{\zeta}}_{f}\|\leq l_{f}$ and $\|\boldsymbol{{\boldsymbol{\zeta}}}_{g_{i}}\|\leq l_{g}$ for all ${\boldsymbol{\zeta}}_{f}\in\partial f({\mathbf{x}})$ and $\boldsymbol{{\boldsymbol{\zeta}}}_{g_{i}}\in\partial g_{i}({\mathbf{x}})$ for any ${\mathbf{x}}\in\mathcal{X}$ and any $i\in[m]$.
+From items (a) and (c) in the above assumption, we have
+[…]
+In item (c), $\partial$ refers to the subdifferential, see Definition 1.1. Item (c) holds naturally when $f$ and $\{g_{i}\}$ are all Lipschitz continuous.
 Complexity measures.
 We will focus on three complexity measures: objective subgradient complexity (OGC), constraint subgradient complexity (CGC) and constraint function value complexity (CFC), detailed as follows:
-OGC and CGC denote the number of calls to the sample subgradients of $f$ and $g$, respectively, while CFC represents the number of evaluations of the sample constraint function $g(\cdot,\xi_{g})$.
+OGC and CGC denote the number of calls to the sample subgradients of $f$ and $\mathbf{g}$}, respectively, while CFC represents the number of evaluations of the sample constraint function ${\mathbf{g}}(\cdot,\xi_{{\mathbf{g}}})$.
 The nonsmooth expectation constraints in Eq.
-P pose significant challenges for designing efficient and reliable stochastic methods, as the projection onto an expectation-constrained set is prohibitively expensive.
-To circumvent the costly projection step, several methods such as online convex method [43], augmented Lagrangian method (ALM) [24], inexact proximal point method (IPP) [9, 30] and switching subgradient method (SSG) [18, 23] have been explored.
-Among these, the online convex method, ALM, and IPP require at least two loops.
-The online convex method [43], developed under a convex setting, relies on a restrictive assumption that the noise associated with stochastic constraints is bounded.
-Under a regularity condition, [24] and [39] demonstrate the effectiveness of the ALM applied to problem Eq.
+P pose significant challenges for designing efficient and reliable stochastic methods, as the projection onto an expectation-constraint set is prohibitively expensive.
+To circumvent the costly projection step, several methods such as online convex method [63], augmented Lagrangian method (ALM) [38], inexact proximal point method (IPP) [10, 28, 46] and switching subgradient method (SSG) [27, 33] have been explored.
+Among these methods, the online convex method, ALM, and IPP all require at least an inner subsolver to address their respective subproblems.
+The online convex method [63] is developed specifically for problems with convex constraints.
+With nonconvex constraints and under suitable regularity conditions, Li et al.
+[38] and Xiao et al.
+[58] demonstrate the effectiveness of the ALM applied to problem Eq.
 P, yielding competitive numerical results.
+Xiao et al.
 [58] establishes convergence to a KKT point.
-To derive complexity results to an $(\epsilon,\epsilon)$-KKT point (see Definition 1.2), [24] assumes additional smoothness on $g$ and exploits this structure to efficiently solve each nonconvex ALM subproblem by an accelerated method.
-Under a certain constraint qualification (CQ), [9] and [30] respectively provide convergence results and complexity analysis for stochastic IPP methods to find an $(\epsilon,\epsilon)$-KKT point.
+To derive complexity results to an $(\epsilon,\epsilon)$-KKT point (see Definition 1.3), Li et al.
+[38] assume additional smoothness on $\mathbf{g}$} and exploit this structure to efficiently solve each nonconvex ALM subproblem by an accelerated method.
+Under a certain constraint qualification (CQ), Boob et al.
+[10] and Ma et al.
+[46] respectively provide convergence results and complexity analysis for stochastic IPP methods to find an $(\epsilon,\epsilon)$-KKT point.
 These methods iteratively solve strongly convex subproblems by incorporating a proximal term into the objective and constraint functions.
@@
 P under nonsmooth weakly convex settings.
-Double (or more)-loop algorithms can utilize optimal subroutines for solving inner subproblems to achieve low-order complexity results.
-However, implementing them needs extra efforts to design appropriate checkable stopping criteria for inner subsolvers or requires careful tuning of the inner iteration counts to reach a satisfactory empirical performance.
+The aforementioned methods utilize optimal subroutines for solving inner subproblems to achieve low-order complexity results.
+Implementing them needs extra efforts to design appropriate checkable stopping criteria for inner subsolvers or requires careful tuning of the inner iteration counts to reach a satisfactory empirical performance.
 This disadvantage is especially pronounced in stochastic settings, where the randomness complicates the verification of a stopping condition and efficient tuning.
-Hence, single-loop methods for solving expectation-constrained problems are preferred in practice.
+Therefore, algorithms that avoid solving subproblems by an iterative subroutine and instead update the solution directly at each iteration are often more desirable for solving expectation-constrained problems in practice.
 However, such methods are still limited.
-Both of [2, 36] design FOMs based on penalty or ALM for constrained stochastic programming but they require smoothness of the objective (possibly with a proximable nonsmooth convex term) and constraint functions.
-The authors in [18] apply a single-loop SSG [34] to problem Eq.
+Both [2, 54] design FOMs based on penalty or ALM for constrained stochastic programming but they require smoothness of the objective (possibly with a proximable nonsmooth convex term) and constraint functions.
+Huang and Lin [27] apply SSG [52] to problem Eq.
 P.
 The SSG is easier to implement while delivering competitive numerical performance.
-Under a (uniform) Slater-type CQ (see 5), the SSG adapts its stepsize depending on whether the current iterate is near feasible or not, switching between fixed and adaptive steps based on Polyak’s rules [34].
-This adaptation is crucial in maintaining near feasibility to achieve a near-KKT point.
-One critical drawback of the SSG is the requirement of frequent checks for near feasibility at generated iterates, leading to a higher CFC and overall complexity compared to IPP in [9] for solving the nonconvex problem Eq.
+Under a (uniform) Slater-type CQ (see 2), the SSG adapts its stepsize depending on whether the current iterate is near feasible or not, switching between fixed and adaptive steps based on Polyak’s rules [52].
+This adaptation is crucial in maintaining near feasibility to achieve a near KKT point.
+However, the SSG requires frequent checks for near-feasibility at generated iterates, resulting in a higher CFC and overall computational complexity compared to the IPP method in [10] when solving the stochastic nonconvex problem Eq.
 P.
-Moreover, in the stochastic setting, [18] requires a light-tail assumption.
-Without the need for feasibility checking, we propose a novel single-loop stochastic FOM for solving Eq.
+Furthermore, in the stochastic setting, Huang et al.
+[27] establish convergence guarantees only under the assumptions that $\mathbf{g}$} is convex and satisfies a light-tail condition.
+Without the need for feasibility checking, we propose a simple SPIDER-type Stochastic Subgradient FOM for solving the Expectation constrained problem Eq.
 P, named 3S-Econ.
-At the $k$-th iteration, 3S-Econ requires unbiased stochastic subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and $\boldsymbol{\zeta}_{g}^{(k)}$ of $g$, evaluated at the iterate ${\mathbf{x}}^{(k)}$, along with an approximation $u^{(k)}$ for $g({\mathbf{x}}^{(k)})$.
+At the $k$-th iteration, 3S-Econ requires unbiased stochastic subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and $\boldsymbol{\zeta}_{g_{i}}^{(k)}$ of $g_{i}$, evaluated at the iterate ${\mathbf{x}}^{(k)}$, along with an approximation $u_{i}^{(k)}$ of $g_{i}({\mathbf{x}}^{(k)})$, for all $i\in[m]$.
 Using this information, it performs the following update step:
-(1) ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\right),$
-where $\beta>0$, $\nu>0$, and $\alpha_{k}>0$ denotes the step size at iteration $k$.
-We will show that 3S-Econ not only is easy to implement but also enjoys lower-order complexity than the best-known results to produce an $(\epsilon,\epsilon)$-KKT point of problem Eq.
-P under milder conditions; see Table 1.
+(2) ${\mathbf{x}}^{(k+1)}=\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\textstyle\frac{u_{i}^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\right)\right),$
+where $\beta>0$ is a penalty parameter, $\nu>0$ is a smoothing parameter, and $\alpha_{k}>0$ denotes the step size at iteration $k$.
+We will show that 3S-Econ is not only easy to implement but also enjoys lower-order complexity than the best-known results to produce an $(\epsilon,\epsilon)$-KKT point of problem Eq.
+P; see Table 1.
 1.1 Applications
@@
 Neyman-Pearson classification.
-Let $\phi(\mathbf{x};\mathbf{a},b)$ be a loss function for a binary classification problem.
-Here, $\mathbf{x}$ is the variable, and ${\mathbf{a}}$ is a data point with a label $b\in\{+1,-1\}$.
-Denote $c>0$ as the threshold of the false-positive error, $\left\{\mathbf{a}_{i}^{+}\right\}_{i=1}^{n^{+}}$ and $\left\{\mathbf{a}_{i}^{-}\right\}_{i=1}^{n^{-}}$ as the positive-class samples (i.e., label $=1$) and negative-class samples (i.e., label $=-1$) in the training dataset, respectively.
-The Neyman-Pearson classification aims to minimize the more concerned false-negative error while controlling false-positive error at an acceptable level by solving
-(2) $\min_{\mathbf{x}\in\mathbb{R}^{d}}\frac{1}{n^{+}}\sum_{i=1}^{n^{+}}\phi\left(\mathbf{x};\mathbf{a}_{i}^{+},1\right),\text{s.t.
-}\frac{1}{n^{-}}\sum_{i=1}^{n^{-}}\phi\left(\mathbf{x};\mathbf{a}_{i}^{-},-1\right)-{c}\leq 0.$
-Clearly, Eq. 2 is a special instance of Eq. P with $f$ having a finite-sum structure, and $\mathbb{P}_{g}$ being a uniform distribution on the negative-class samples.
+Consider a classification problem involving training data partitioned into $M$ classes, denoted by $\mathcal{D}_{i}\subseteq\mathbb{R}^{d}$ for $i\in[M]$.
+Let $\phi$ be a loss function, $\{{\mathbf{x}}_{i}\}_{i\in[M]}$ denote model variables, $D>0$, and $\kappa$ be a prescribed threshold.
+The multi-class Neyman-Pearson classification aims to minimize the average pairwise loss for a specified class (e.g., class ${\mathcal{D}}_{1}$), while controlling the average pairwise losses for all other classes below a threshold.
+Formally, the optimization problem is given by
+(3) $\min_{\left\|\mathbf{x}_{i}\right\|\leq D,i\in[M]}\frac{1}{\left|\mathcal{D}_{1}\right|}\sum_{l>1}\sum_{{\mathbf{a}}_{1}\in\mathcal{D}_{1}}\phi\left(\mathbf{x}_{1}^{\top}{\mathbf{a}}_{1}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{1}\right),$
+$\text{ s.t.
+}\frac{1}{\left|\mathcal{D}_{i}\right|}\sum^{M}_{l=2,l\neq i}\sum_{{\mathbf{a}}_{i}\in\mathcal{D}_{i}}\phi\left(\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}\right)\leq\kappa,\quad i=2,3,\ldots,M.$
+Clearly, Eq.
+3 is a special instance of Eq.
+P with $f$ having a finite-sum structure, and $\mathbb{P}_{{\mathbf{g}}}$ corresponds to a uniform distribution over sample sets associated with $\{{\mathcal{D}}_{i}\}_{i=2,\ldots,M}$.
 Example 1.2.
@@
 Traditional models focus on maximizing accuracy, which can lead to unfair performance across different groups of data, such as males and females.
-Fairness constraints are applied to balance outcomes for all groups, using principles like demographic parity [1] or equalized odds [17].
+Fairness constraints are applied to balance outcomes for all groups, using principles like demographic parity [1, 35] or equalized odds [26].
 Achieving fairness requires trade-offs, as improving fairness can reduce overall accuracy.
@@
 Let us consider a classification problem with receiver operating characteristic (ROC)-based fairness [56].
-Suppose there are three groups of data: an unlabeled set $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, a protected set $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$, and an unprotected set $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$.
-Here, the protected set contains the features associated with the minority group, while the unprotected set includes the features from the majority group.
+Suppose the data are partitioned into three parts: a dataset $\mathcal{D}=\left\{(\mathbf{a}_{i},b_{i})\right\}_{i=1}^{n}$, and two labeled subgroups $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$, defined according to a binary protected attribute (e.g., race or gender).
+Here, group $\mathcal{D}_{\mathrm{p}}$ corresponds to the minority subgroup with historically less favorable treatment, while group $\mathcal{D}_{\mathrm{u}}$ represents the majority subgroup.
 The fairness-constrained problem can be formulated as
-(3) $\vspace{-1mm}\min_{{\mathbf{x}}:\|{\mathbf{x}}\|\leq D}\Phi_{\Theta}({\mathbf{x}}),\,\,\text{{s.t.}}\,\,\Phi({\mathbf{x}})\leq\Phi^{*}+\kappa_{1},$
+(4) $\vskip-2.84526pt\min_{{\mathbf{x}}:\|{\mathbf{x}}\|\leq D}\Phi_{\Theta}({\mathbf{x}}),\,\,\text{{s.t.}}\,\,\Phi({\mathbf{x}})\leq\Phi^{*}+\kappa_{1},$
 where $\Phi_{\Theta}({\mathbf{x}}):=\max_{\theta\in\Theta}\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\geq\theta\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\geq\theta\right)\right|$ is a ROC-based fairness, $\Phi^{*}:=\min_{{\mathbf{x}}}\Phi({\mathbf{x}})$ with $\Phi({\mathbf{x}}):=\frac{1}{n}\sum_{i=1}^{n}\left[1-b_{i}\cdot{\mathbf{x}}^{\top}{\mathbf{a}}_{i}\right]_{+}$, $\Theta$ is a finite threshold set, and $\kappa_{1}>0$ is a slackness parameter.
@@
 This tool transforms problem Eq.
-P into an unconstrained one.
-Notably, when the penalty parameter is sufficiently large, the stationarity point of the penalized problem can align with that of the original constrained problem.
-Building on the exact penalization, we then design a stochastic FOM to solve the resulting unconstrained problem.
-Using the penalty function $h=[\cdot]_{+}:a\to\max\{a,0\}$ for an inequality constraint, an exact penalty reformulation [8, 33, 38, 45] of problem Eq. P is given by
-(5) $\min_{{\mathbf{x}}}\,\,F({\mathbf{x}})=f({\mathbf{x}})+\beta g_{+}({\mathbf{x}}),\text{ with }g_{+}({\mathbf{x}})=h(g({\mathbf{x}})),$
+P into a problem without function constraints.
+Notably, when the penalty parameter is sufficiently large, the stationary point of the penalized problem can align with the KKT point of the original problem.
+Building on the exact penalization, we then design a stochastic FOM to solve the resulting problem.
+Let the penalty function $h({\mathbf{z}})={\mathbf{e}}^{\top}[{\mathbf{z}}]_{+}$, where ${\mathbf{e}}\in\mathbb{R}^{m}$ is the all-ones vector and ${\mathbf{z}}]_{+}:=(\max\{z_{1},0\},\ldots,\max\{z_{m},0\})^{\top}$ for any ${\mathbf{z}}\in\mathbb{R}^{m}$.
+Then an exact penalty reformulation [9, 50, 51, 57, 62, 65] of problem Eq.
+P is
+(6) $\min_{{\mathbf{x}}\in\mathcal{X}}\,\,F({\mathbf{x}}):=f({\mathbf{x}})+\beta g_{+}({\mathbf{x}}),\text{ with }g_{+}({\mathbf{x}})=h({\mathbf{g}}({\mathbf{x}}))=\sum_{i=1}^{m}[g_{i}({\mathbf{x}})]_{+},$
 where $\beta>0$ is the penalty parameter.
 The exact penalization results regarding global minimizers are established in [50].
-Note that $\mathrm{Sign}([g(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g}$ belongs to $\partial g_{+}(\mathbf{x})$, where $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$.
-Here, $\mathrm{Sign}:\mathbb{R}\mapsto\{-1,0,1\}$ refers to the sign operator, and $\partial$ denotes the (Clarke) subdifferential [12, Section 1.2] of a continuous function.
-Due to the stochastic nature of $g$, it is impractical to compute the exact value of $\mathrm{Sign}([g(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g}$.
-Approximating this value is also challenging, because $\mathrm{Sign}$ is discontinuous, making it impossible for the discrepancies between $\mathrm{Sign}([g(\mathbf{x})]_{+})$ and $\mathrm{Sign}([g(\mathbf{x},\mathcal{B})]_{+})$ to be consistently small enough.
-Here, $\mathcal{B}$ is a sample set of $\xi_{g}$ over the distribution $\mathbb{P}_{g}$, and
-(6) $g(\mathbf{x},\mathcal{B}):=\frac{1}{|\mathcal{B}|}\sum_{\xi_{g}\in\mathcal{B}}g({\mathbf{x}},\xi_{g}),$
-with $|\mathcal{B}|$ denoting the cardinality of $\mathcal{B}$.
-Consequently, solving problem Eq.
-5 directly by a stochastic method is challenging.
-These difficulties extend to algorithm design and convergence analysis, complicating the establishment of theoretical guarantees.
-To resolve this issue, we introduce a smoothed function of $h$ with a smoothing parameter $\nu>0$ as follows,
-$\vspace{-3mm}h^{\nu}(z)$ $:=\max_{0\leq y\leq 1}\langle y,z\rangle-\frac{\nu}{2}y^{2}=\left\{\begin{aligned} \frac{z^{2}}{2\nu},\quad&\text{ if }0\leq z\leq\nu,\\ z-\frac{\nu}{2},&\text{ if }z>\nu,\\ 0,\quad\,\,\,&\text{ otherwise.}\end{aligned}\right.$
-We then present a new penalty model of problem Eq. P:
-(PP) $\min_{{\mathbf{x}}}\,\,F^{\nu}({\mathbf{x}})=f({\mathbf{x}})+\beta h^{\nu}(g({\mathbf{x}})).$
-Since $h^{\nu}$ is smooth and $\nabla h^{\nu}(z)=\mathrm{Proj}_{[0,1]}\left(\frac{z}{\nu}\right)$, we can access a subgradient of $h^{\nu}(g({\mathbf{x}}))$ via $\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g}$, where $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$.
-In addition, as $\mathrm{Proj}_{[0,1]}$ is Lipschitz continuous, the difference between $\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g}$ and $\mathrm{Proj}_{[0,1]}\left(\frac{u}{\nu}\right)\boldsymbol{\zeta}_{g}$ can be well controlled if $u$ is a good approximation of $g({\mathbf{x}})$.
-We apply a SPIDER-type (or SARAH) [16, 32] variance reduction technique to estimate $g({\mathbf{x}}^{(k)})$ to obtain $u^{(k)}$ and perform the update in Eq.
-1 for each $k$.
-The variance reduction tool, together with the smoothing tool, enables us to achieve better complexity results for our method than state-of-the-art results to produce an $(\epsilon,\epsilon)$-KKT point of (P).
+Note that $\sum_{i=1}^{m}\mathrm{Sign}([g_{i}(\mathbf{x})]_{+})\boldsymbol{\zeta}_{g_{i}}$ belongs to $\partial g_{+}(\mathbf{x})$, where $\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}(\mathbf{x})$, and $\mathrm{Sign}:\mathbb{R}\mapsto\{-1,0,1\}$ refers to the sign operator.
+Due to the stochastic nature of $\mathbf{g}$}, it is impractical to compute the exact value of $\mathrm{Sign}([g_{i}(\mathbf{x})]_{+})$ for each $i$.
+Approximating this value is also challenging, because $\mathrm{Sign}$ is discontinuous, making it impossible for the discrepancies between $\mathrm{Sign}([g_{i}(\mathbf{x})]_{+})$ and $\mathrm{Sign}([g_{i}(\mathbf{x},\mathcal{B})]_{+})$ to be consistently small enough.
+Here, $\mathcal{B}$ is a sample set of $\xi_{{\mathbf{g}}}$ over the distribution $\mathbb{P}_{{\mathbf{g}}}$, and we define
+(7) $g_{i}(\mathbf{x},\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{\xi_{{\mathbf{g}}}\in\mathcal{B}}g_{i}({\mathbf{x}},\xi_{{\mathbf{g}}}),\forall\,i\in[m],\quad{\mathbf{g}}({\mathbf{x}},{\mathcal{B}})=\left(g_{1}(\mathbf{x},\mathcal{B}),\ldots,g_{m}(\mathbf{x},\mathcal{B})\right)^{\top},$
+with $|\mathcal{B}|$ denoting the cardinality of $\mathcal{B}$. Consequently, solving problem Eq. 6 directly by a stochastic method is challenging.
+To resolve this issue, we employ the Huber-type [^1] The smoothing technique is a powerful tool for handling nonsmooth problems, and we refer interested readers to [11, 12, 15] for further details and illustrative examples.
+Specifically, the function $\cdot\,$$_{+}$ admits various smooth approximations, including the Softplus function [36, 37], the square-root function [14], and the log-sum-exp function [15, 41].
+Our complexity results remain valid as long as the smoothing function $h^{\nu}$ is Lipschitz continuous, its gradient has a Lipschitz constant that grows linearly with $1/\nu$, and the composition $h^{\nu}({\mathbf{g}}(\cdot))$ is weakly convex.
+The mentioned Huber-type, Softplus and log-sum-exp smoothing functions satisfy all these conditions.
+We choose the Huber-type smoothing because its gradient takes the particularly simple form $\mathrm{Proj}_{[0,1]}(\cdot/\nu)$, which facilitates a clean and efficient algorithmic implementation.
+smoothing [16, 48, 55] function $h^{\nu}$ given as follows,
+$\vskip-8.53581pth^{\nu}({\mathbf{z}})$ $:=\sum_{i=1}^{m}H^{\nu}(z_{i}),\text{ with }H^{\nu}(z_{i})=\max_{0\leq y\leq 1}\left[\langle y,z_{i}\rangle-\frac{\nu}{2}y^{2}\right]=\left\{\begin{aligned} \frac{z_{i}^{2}}{2\nu},\quad&\text{ if }0\leq z_{i}\leq\nu,\\ z_{i}-\frac{\nu}{2},&\text{ if }z_{i}>\nu,\\ 0,\quad\,\,\,&\text{ otherwise},\end{aligned}\right$.
+where $\nu>0$ is a smoothing parameter. Then instead of directly working on Eq. 6, we present a new penalty model of problem Eq. P:
+($P$) $\min_{{\mathbf{x}}\in\mathcal{X}}\,\,F^{\nu}({\mathbf{x}})=f({\mathbf{x}})+\beta h^{\nu}({\mathbf{g}}({\mathbf{x}})).$
+Since $h^{\nu}$ is smooth and $\nabla h^{\nu}({\mathbf{z}})=\mathrm{Proj}_{[0,1]^{m}}\left(\frac{{\mathbf{z}}}{\nu}\right)$, we can access a subgradient of $h^{\nu}({\mathbf{g}}({\mathbf{x}}))$ via $\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}$, where $\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}(\mathbf{x})$ for each $i\in[m]$.
+In addition, as $\mathrm{Proj}_{[0,1]}$ is Lipschitz continuous, the difference between $\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}})}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}$ and $\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}$ can be well controlled if $u_{i}$ is a good approximation of $g_{i}({\mathbf{x}})$.
+For each $i\in[m]$ and each $k$, we apply a SPIDER-type (or SARAH) [24, 49] variance reduction technique to estimate $g_{i}({\mathbf{x}}^{(k)})$ to obtain $u_{i}^{(k)}$ and perform the update in Eq.
+2.
+The variance reduction tool, together with the smoothing tool, enables us to achieve better complexity results for our method than state-of-the-art results to produce an $(\epsilon,\epsilon)$-KKT point of Eq.
+P.
 1.3 Contributions
@@
 First, to address the expectation constraint involved in Eq.
-P, we introduce a new smooth penalty function $h^{\nu}\circ g$.
+P, we introduce a Huber-type smoothing penalty function $h^{\nu}\circ{\mathbf{g}}$.
 Leveraging this function, we propose solving problem Eq.
@@
 Our framework is well-suited for many applications such as problem Eq.
-4, which involves a nonconvex objective and convex expectation constraints satisfying the required assumptions.
+5, which involves a nonconvex objective and convex expectation constraints satisfying the required assumptions.
 Second, we exploit the structure of problem Eq.
-PP and propose a single-loop SPIDER-type stochastic FOM to solve it, with updates given in Eq.
-1.
-At each iteration, the SPIDER technique [16] is employed to generate an approximation $u^{(k)}$ of the constraint function value.
+PP and propose a SPIDER-type stochastic FOM to solve it, with updates given in Eq.
+2.
+At each iteration, the SPIDER technique [24] is employed to generate an approximation $u_{i}^{(k)}$ of the constraint function value $g_{i}({\mathbf{x}}^{(k)})$ for each $i\in[m]$.
 Using the tool of Moreau envelope, we demonstrate that the proposed method can, in expectation, find a near-$\epsilon$ stationary point of problem Eq.
@@
 P in expectation.
-Based on the structure of $g$, we categorize the complexity results of our method as follows:
+Based on the structure of $\mathbf{g}$}, we categorize the complexity results of our method as follows:
   1. 1.
-    A generic stochastic $g$: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\epsilon^{-6})$
+    A generic stochastic $\mathbf{g}$}: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\epsilon^{-6})$ .
   2. 2.
-    A finite-sum structured [^1] Here, we assume that $\mathbb{P}_{g}$ is a uniform distribution over a finite set of size $N$, where $N<\epsilon^{-4}$.
-$g$: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\sqrt{N}\epsilon^{-4})$.
-When $g$ is generic stochastic, our complexity result is lower than state-of-the-art results by a factor of $\epsilon^{-2}$ in either OGC or CFC; when $g$ enjoys a finite-sum structure, the highest order among our results of OGC, CGC, and CFC is $O(\sqrt{N}\epsilon^{-4})$, representing an improvement over the previously best-known $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$ bound.
+    A finite-sum structured [^2] Here, we assume that $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite set of size $N$, where $N<\epsilon^{-4}$.
+$\mathbf{g}$}: OGC is $O(\epsilon^{-4})$, CGC is $O(\epsilon^{-4})$, and CFC is $O(\sqrt{N}\epsilon^{-4})$.
+When $\mathbf{g}$} is generic stochastic, our complexity result is lower than state-of-the-art results by a factor of $\epsilon^{-2}$ in either OGC or CFC; when $\mathbf{g}$} enjoys a finite-sum structure, the highest order among our results of OGC, CGC, and CFC is $O(\sqrt{N}\epsilon^{-4})$, representing an improvement over the previously best-known $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$ bound.
 Detailed comparisons are provided in Table 1.
-In this table, the sharpness assumption and the gradient-away assumption that we make are weaker than the (uniform) Slater-type CQ in [18, 30], which is weaker than a strong feasibility condition (see Eqn.
-Eq.
-41 below) in [9].
-These assertions will be verified in Appendix A.
-Third, the ability to handle nonsmooth problems using a single-loop approach eliminates the need for tuning inner iterations, distinguishing the proposed method from double-loop methods, such as those in [9, 24, 30].
+Third, our approach does not need to solve subproblems by an iterative subroutine and thus eliminates the need for tuning inner iterations, distinguishing our method from those in [10, 38, 46].
 This simplification reduces computational overhead and makes the method more accessible for real-world applications.
-In numerical experiments, we test both the deterministic and stochastic versions of our method, for two problems across two datasets.
+In numerical experiments, we test both the deterministic and stochastic versions of our method, for three problems across three datasets.
 Notably, the deterministic version is significantly faster than state-of-the-art methods, and the stochastic version achieves even more impressive speedups.
-Methods SL structure of $g$ CQ on $g$ OGC CGC CFC
-IPP [9] no stochastic convex $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
-finite-sum convex MFCQ, and $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
-stochastic WC strong feasibility $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
-finite-sum WC $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
-SSG [18] yes stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-8})$
+Methods RIS structure of $\mathbf{g}$} CQ on $\mathbf{g}$} OGC CGC CFC
+IPP [10] yes stochastic convex MFCQ, and $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
+finite-sum convex Slater’s condition $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
+stochastic WC MFCQ, and $O(\epsilon^{-6})$ $O(\epsilon^{-6})$
+finite-sum WC strong feasibility $O(\epsilon^{-4})$ $O(\min\{N\epsilon^{-4},\epsilon^{-6}\})$
+SSG [27] no stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-8})$
 finite-sum convex $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(N\epsilon^{-4})$
 WC none
-Ours yes stochastic convex sharpness, or $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
-finite-sum convex gradient away $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
-stochastic WC gradient away, and $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
-finite-sum WC $B_{g}>\rho_{g}D^{2}$ $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
-finite-sum WC Slater-type CQ $O(\epsilon^{-4})$ $O(N\epsilon^{-4})$ $O(N\epsilon^{-4})$
-[…]
+Ours no stochastic convex Slater-type CQ $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
+finite-sum convex $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
+stochastic WC Slater-type CQ, with $O(\epsilon^{-4})$ $O(\epsilon^{-4})$ $O(\epsilon^{-6})$
+finite-sum WC $B_{g}>m\rho_{g}D^{2}$ $O(\epsilon^{-4})$ $O(\sqrt{N}\epsilon^{-4})$
+Table 1: Comparison of complexity results among our method and existing ones for obtaining an $(\epsilon,\epsilon)$-KKT point of problem Eq.
+P with a stochastic nonsmooth weakly convex objective function $f$.
+RIS is short for “Require Inner Solver”, and WC is for “weakly convex”.
+Strong feasibility implies the Slater-type CQ with any $B_{g}>0$; see Proposition 2.1(ii).
+The Slater-type CQ is directly assumed in [27] when $\{g_{i}\}$ are weakly convex.
+Under the setting where $g_{i}$ is convex for each $i\in[m]$, Slater’s condition is equivalent to the Slater-type CQ; see Proposition 2.1(i).
+In the finite-sum setting, we assume that $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite set of size $N<\epsilon^{-4}$.
+When $\{g_{i}\}_{i\in[m]}$ are weakly convex, Huang et al.
+[27] provide complexity results only for the case where both $f$ and $\mathbf{g}$} are deterministic, excluding the setting where $f$ is stochastic.
 1.4 Notations and definitions
-[…]
-(7) $\phi^{\prime}({\mathbf{x}};{\mathbf{v}})\geq{\mathbf{v}}^{\top}{\boldsymbol{\gamma}},\text{ for any }{\boldsymbol{\gamma}}\in\partial\phi({\mathbf{x}}).$
+We use ${\mathrm{dom}}(\cdot)$ to denote the domain of a function.
+$[M]$ denotes $\{1,\ldots,M\}$ for a positive integer $M$.
+The 0-$\infty$ indicator function of a set $\mathcal{X}$} is denoted as $\iota_{\mathcal{X}}$.
+The $l_{2}$-norm is written as $\|\cdot\|$.
+$\mathrm{co}(\cdot)$} and $\mathrm{relint}(\cdot)$ represent the closed convex hull and relative interior of a set, respectively.
+$\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$ refers to the normal cone of a closed convex set $\mathcal{X}$ at $\mathbf{x}$}.
+We use $\mathrm{Proj}$, $\mathrm{Prob}$, and $\mathrm{dist}$ to denote operators of the projection, the probability, and the distance, respectively.
+Definition 1.1 (subdifferential).
+For a convex function $f$, its subdifferential at ${\mathbf{x}}\in{\mathrm{dom}}(f)$ is defined as [18]
+$\partial f({\mathbf{x}})=\big{\{}{\boldsymbol{\xi}}:f({\mathbf{x}}^{\prime})-f({\mathbf{x}})\geq\langle{\boldsymbol{\xi}},{\mathbf{x}}^{\prime}-{\mathbf{x}}\rangle\ \text{ for all }{\mathbf{x}}^{\prime}\in{\mathrm{dom}}(f)\big{\}}.$
+The subdifferential of a $\rho$-weakly convex function $f$ at ${\mathbf{x}}\in{\mathrm{dom}}(f)$ is given by
+$\partial f({\mathbf{x}}):=\textstyle\partial\big{(}f(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}\big{)}({\mathbf{x}})-\rho{\mathbf{x}}.$
 Definition 1.2 (Moreau Envelope [21]).
-For a $\rho_{\phi}$-weakly convex function $\phi$, the Moreau envelope and the proximal mapping for any $t\in(0,\frac{1}{\rho_{\phi}})$ are defined by $\phi_{t}({{\mathbf{x}}}):=\min_{{\mathbf{z}}}\left\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\right\}$ and $\operatorname{Prox}_{t\phi}({{\mathbf{x}}}):=\arg\min_{\mathbf{z}}\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\}.$
-We give the definition of stationarity as follows. In deterministic cases—where randomness is not involved—the expectation operator $\mathbb{E}$ should be removed from the definitions below.
+For a $\rho_{\phi}$-weakly convex function $\phi$, its Moreau envelope and proximal mapping for any $t\in(0,\frac{1}{\rho_{\phi}})$ on $\mathcal{X}$ are defined by $\phi_{t}({{\mathbf{x}}})=\min_{{\mathbf{z}}\in\mathcal{X}}\left\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\right\}$, $\operatorname{Prox}_{t\phi}({{\mathbf{x}}})=\arg\min_{{\mathbf{z}}\in\mathcal{X}}\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\}.$
+We give the definition of stationarity as follows. In deterministic cases, where randomness is not involved, the expectation operator $\mathbb{E}$ should be removed from the definition below.
 Definition 1.3 (stationarity).
-Let $\epsilon>0$ and $\delta>0$, and suppose $\psi$ and $\phi$ are $\rho_{\psi}$- and $\rho_{\phi}$-weakly convex functions.
-$\mathrm{(a)}$ We call ${\mathbf{x}}^{*}$ a near-$\epsilon$ stationary point of $\min_{{\mathbf{x}}}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)$, and there exists $0<t<\frac{1}{\rho_{\phi}}$, such that $\mathbb{E}[\|\nabla\phi_{t}({\mathbf{x}}^{*})\|^{2}]=\mathbb{E}[\|\frac{1}{t}(\operatorname{Prox}_{t\phi}({{\mathbf{x}}^{*}})-{\mathbf{x}}^{*})\|^{2}]\leq~{}\epsilon^{2}.$
-$\mathrm{(b)}$ We call ${\mathbf{x}}^{*}$ an $\epsilon$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if it holds that ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)\cap{\mathrm{dom}}(\psi)$, and there exists $\lambda_{{\mathbf{x}}^{*}}\geq 0$ such that
-$\mathbb{E}\left[\textnormal{dist}(\mathbf{0},\partial\phi({\mathbf{x}}^{*})+\lambda_{{\mathbf{x}}^{*}}\partial\psi({\mathbf{x}}^{*}))\right]\leq\epsilon,\,\,\mathbb{E}\left[\psi({\mathbf{x}}^{*})\right]\leq\epsilon,\text{ and }|\lambda_{{\mathbf{x}}^{*}}\psi({\mathbf{x}}^{*})|\leq\epsilon.$
-$\mathrm{(c)}$ We call ${\mathbf{x}}^{*}$ a $(\delta,\epsilon)$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if it holds that ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)\cap{\mathrm{dom}}(\psi)$, and $\mathbb{E}[\|{\mathbf{x}}^{*}-\widetilde{{\mathbf{x}}}^{*}\|]\leq\delta$ with $\widetilde{{\mathbf{x}}}^{*}$ being an $\epsilon$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation.
-1.5 Organizations
+Let $\epsilon>0$. Suppose that $\{\psi_{i}\}_{i=1}^{m}$ are all $\rho_{\psi}$-weakly convex and $\phi$ is $\rho_{\phi}$-weakly convex on $\mathcal{X}$}.
+$\mathrm{(a)}$ $\mathbf{x}$$^{*}$ is called a near-$\epsilon$ stationary point of $\min_{{\mathbf{x}}\in{\mathcal{X}}}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in{\mathcal{X}}$, and there exists $0<t<\frac{1}{\rho_{\phi}}$, such that $\mathbb{E}[\|\nabla\phi_{t}({\mathbf{x}}^{*})\|^{2}]\leq~\epsilon^{2}.$
+$\mathrm{(b)}$ $\mathbf{x}$$^{*}$ is called an $\epsilon$-KKT point of $\min_{{\mathbf{x}}\in\mathcal{X},\boldsymbol{\psi}({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in\mathcal{X}$ and there exists $\boldsymbol{\lambda}_{{\mathbf{x}}^{*}}\geq 0$ such that $\mathbb{E}\left[\psi_{i}({\mathbf{x}}^{*})\right]\leq\epsilon,$ $|({\lambda}_{{\mathbf{x}}^{*}})_{i}{\psi}_{i}({\mathbf{x}}^{*})|\leq\epsilon$ for all $i\in[m]$, and $\mathbb{E}\left[\textnormal{dist}(\mathbf{0},\partial\phi({\mathbf{x}}^{*})+\sum_{i=1}^{m}(\lambda_{{\mathbf{x}}^{*}})_{i}\partial\psi_{i}({\mathbf{x}}^{*})+\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*}))\right]\leq\epsilon.$
+$\mathrm{(c)}$ $\mathbf{x}$$^{*}$ is called an $(\epsilon,\epsilon)$-KKT point of $\min_{{\mathbf{x}}\in\mathcal{X},\boldsymbol{\psi}({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in\mathcal{X}$ and $\mathbb{E}[\|{\mathbf{x}}^{*}-\widetilde{{\mathbf{x}}}^{*}\|]\leq\epsilon$ with $\widetilde{{\mathbf{x}}}^{*}$ being an $\epsilon$-KKT point of $\min_{{\mathbf{x}}\in\mathcal{X},\boldsymbol{\psi}({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation.
+1.5 Organization
 The rest of this paper is organized as follows.
 Section 2 introduces the exact penalty models.
-In Section 3, we propose a single-loop SPIDER-type stochastic subgradient method and establish its iteration complexity results.
-Numerical experiments are presented in Section 4, showcasing the performance of our method.
+In Section 3, we propose a SPIDER-type stochastic subgradient method and establish its iteration complexity results.
+Numerical results are presented in Section 4, showcasing the performance of our method.
 Finally, Section 5 concludes the paper.
@@
 6 and Eq.
-PP, which are presented in the following diagram for the convenience of the reader.
-Details are given in the mentioned lemmas and theorem.
+PP, which are presented in the following diagram for the convenience of readers.
+Details are given in the mentioned lemma and theorem.
 […]
-2.1 Exact penalization with penalty function $[\cdot]_{+}$
-We study the relationship between models Eq.
+2.1 Regularity conditions
+To derive the exact penalization result, we assume the following (uniform) Slater-type CQ, which holds for tested instances of Examples 1.1 and 1.2 in Section 4.
+Assumption 2 ((uniform) Slater-type CQ [27, 46]).
+There exist $B_{g}>~0$, $B>0$, and $\overline{\rho}>\rho_{g}\geq 0$, such that for any ${\mathbf{x}}{\in{\mathcal{X}}}$ satisfying $0<g_{+}({\mathbf{x}})\leq B_{g}$ with $g_{+}$ defined in (6), it holds
+(8) $\max_{i\in[m]}g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B,\text{ for some }\mathbf{y}\in\mathrm{relint}(\mathcal{X}).$
+Below we give the relationships among the Slater-type CQ, Slater’s condition and the strong feasibility condition assumed in [10], where the latter two conditions are stated as follows:
+(9) $\text{Slater's condition: }\exists\,{\mathbf{y}}_{\mathrm{feas}}\in\mathrm{relint}({\mathcal{X}})\text{ such that }\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})<0;$
+(10) $\text{strong feasibility: }\exists\,{\mathbf{y}}_{\mathrm{feas}}\in\mathcal{X}\text{ such that }\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})\leq-2\rho_{g}D^{2}.$
+Proposition 2.1.
+Under 1, the following statements hold:
+  • $(\mathrm{i})$
+    If $\{g_{i}\}_{i\in[m]}$ are convex, Slater’s condition and the Slater-type CQ are equivalent.
+  • $(\mathrm{ii})$
+    If the strong feasibility condition holds, so does the Slater-type CQ.
+Proof 2.2.
+(i) Suppose Slater’s condition holds.
+Let ${\mathbf{y}}={\mathbf{y}}_{\mathrm{feas}},$ $\overline{\rho}=\frac{-\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})}{D^{2}},$ $B=\frac{-\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})}{2}$, and $B_{g}$ be any positive constant.
+Then Eq.
+8 trivially holds with the chosen $\mathbf{y}$}, $\overline{\rho}$, $B$, and any ${\mathbf{x}}\in{\mathcal{X}}$.
+Thus Slater-type CQ holds since $\rho_{g}=0<\overline{\rho}$.
+Conversely, if the Slater-type CQ holds, then by definition there exist ${\mathbf{y}}\in\mathrm{relint}({\mathcal{X}})$ and $B>0$ such that $\max_{i\in[m]}g_{i}({\mathbf{y}})\leq-B<0,$ which immediately implies Slater’s condition.
+Hence, we establish the equivalence.
+(ii) Because ${\mathcal{X}}\subset\mathbb{R}^{d}$ is nonempty convex and ${\mathbf{y}}_{\mathrm{feas}}\in\mathcal{X}$, there must exist ${\mathbf{y}}\in\mathrm{relint}({\mathcal{X}})$ such that $\|{\mathbf{y}}-{\mathbf{y}}_{\mathrm{feas}}\|\leq\frac{\rho_{g}D^{2}}{2l_{g}}$.
+By the $l_{g}$-Lipschitz continuity of each $g_{i}$ from 1(c), it holds $g_{i}(\mathbf{y})-g_{i}(\mathbf{y}_{\mathrm{feas}})\leq l_{g}\|{\mathbf{y}}-{\mathbf{y}}_{\mathrm{feas}}\|\leq\rho_{g}D^{2}/2,\,\forall\,i\in[m]$.
+Thus $\max_{i\in[m]}g_{i}(\mathbf{y})\leq\max_{i\in[m]}g_{i}(\mathbf{y}_{\mathrm{feas}})+\rho_{g}D^{2}/2$.
+Now let $\overline{\rho}=2\rho_{g}$.
+We have for any ${\mathbf{x}}\in{\mathcal{X}}$ that
+$\max_{i\in[m]}g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$ $\leq\max_{i\in[m]}g_{i}(\mathbf{y}_{\mathrm{feas}})+\frac{\rho_{g}D^{2}}{2}+\rho_{g}\|\mathbf{y}-\mathbf{x}\|^{2}$
+$\leq-2\rho_{g}D^{2}+\frac{\rho_{g}D^{2}}{2}+\rho_{g}D^{2}=-\frac{\rho_{g}D^{2}}{2}.$
+Hence, the Slater-type CQ holds with $B=\frac{\rho_{g}D^{2}}{2}$, $\overline{\rho}=2\rho_{g}$, and any $B_{g}>0$.
+We highlight that the strong feasibility condition in Eq.
+10 ensures the Slater-type CQ with any $B_{g}>0$.
+Later in the proof of Theorem 2.9, we assume Slater-type CQ with $B_{g}>m\rho_{g}D^{2}$, which is still implied by the strong feasibility condition.
+However, the converse does not hold in general.
+The following example illustrates that a Slater-type CQ with $B_{g}>m\rho_{g}D^{2}$ does not imply the strong feasibility condition.
+Example 2.3.
+Consider the nonconvex function $g(x)=-6+8x+x^{2}-x^{3}$ over the set $\mathcal{X}=[0,1].$ Since $g^{\prime\prime}(x)=2-6x$, it holds $g^{\prime\prime}(x)\geq-4,\,\forall x\in{\mathcal{X}}$.
+Then $g$ is $4$-weakly convex on $\mathcal{X}$}.
+Take $y=0.01\in\mathrm{relint}({\mathcal{X}})$.
+It is straightforward to verify that $g(y)+\frac{5}{2}|x-y|^{2}\leq-3$ for any $x\in\mathcal{X}$.
+Thus, the Slater-type CQ holds with parameters $B_{g}=1000,B=3$, and $\overline{\rho}=5>4=\rho_{g}.$ Thus, we have $B_{g}>\rho_{g}mD^{2}$, where $m=1$, and $D=1$.
+However, the strong feasibility condition fails because it holds that $\min_{x\in[0,1]}g(x)=-6\;>\;-8=-2\rho_{g}D^{2}.$
+2.2 Exact penalization with penalty function $\cdot$$_{+}$
+In this subsection, we study the relationship between Eq.
 P and Eq.
-5 in this subsection.
-Since $g$ is $\rho_{g}$-weakly convex, it follows that $[g({\mathbf{x}})]_{+}=\max\{g({\mathbf{x}})+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2},\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}\}-\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is $\rho_{g}$-weakly convex.
-Recalling that $f$ is $\rho_{f}$-weakly convex, $F$ in Eq.
-5 is $\rho$-weakly convex with $\rho:=\rho_{f}+\beta\rho_{g}$.
-Under either of the two assumptions below, we establish the equivalence between the $\epsilon$-KKT point of problem Eq. 5 and that of Eq. P.
-Assumption 2 (sharpness).
-There exist $\theta>0$ and $B_{g}>0$ such that for any ${\mathbf{x}}$ satisfying $0<g({\mathbf{x}})\leq B_{g}$, it holds $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$.
-Assumption 3 ((uniformly) gradient-away).
-There exist $\theta>0$ and $B_{g}>0$ such that for any ${\mathbf{x}}$ satisfying $0<g({\mathbf{x}})\leq B_{g}$, it holds $\min_{\boldsymbol{{\boldsymbol{\zeta}}}_{g}\in\partial g(\mathbf{x})}\left\|\boldsymbol{{\boldsymbol{\zeta}}}_{g}\right\|\geq\theta.$
+6.
+Since each $g_{i}$ is $\rho_{g}$-weakly convex for $i\in[m]$, and $f$ is $\rho_{f}$-weakly convex, $F$ in Eq.
+6 is $\rho$-weakly convex with $\rho:=\rho_{f}+\beta m\rho_{g}$.
+First, we present a key lemma as follows, whose proof is given in Appendix A.
 Lemma 2.4.
-Suppose 1 and 2 hold, and $g$ is convex.
-Given $\epsilon>0$, let $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$.
-If ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$ is an $\epsilon$-KKT point of problem Eq.
-5, then ${\mathbf{x}}^{*}\in\Omega$, i.e., ${\mathbf{x}}^{*}$ is a feasible point of problem Eq.
-P, and ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
+Under 1 and 2, let $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$ and ${\mathbf{x}}\in\mathcal{X}$ satisfy $0<g_{+}({\mathbf{x}})\leq B_{g}$.
+Given $\mathcal{J}_{1}\subset[m]$ such that $\mathcal{J}_{1}\cap\{i:g_{i}({\mathbf{x}})\geq 0\}\neq\emptyset$, let $\mathcal{J}_{2}=[m]\setminus\mathcal{J}_{1}$ and choose $\{\lambda_{i}\}_{i\in\mathcal{J}_{2}}\subset\mathbb{R}_{+}$ satisfying $\lambda_{i}g_{i}({\mathbf{x}})\geq 0$ for all $i\in\mathcal{J}_{2}$.
+Then,
+$\min_{\{\boldsymbol{\zeta}_{g_{i}}\in\partial[g_{i}({\mathbf{x}})]_{+}\}_{i\in\mathcal{J}_{1}},\{\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}({\mathbf{x}})\}_{i\in\mathcal{J}_{2}},{\mathbf{u}}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}})}\left\|\sum_{i\in\mathcal{J}_{1}}\zeta_{g_{i}}+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\zeta_{g_{i}}+{\mathbf{u}}\right\|\geq\theta.$
+Now, we establish the relationship between the $\epsilon$-stationary point of problem Eq. 6 and the $\epsilon$-KKT point of Eq. P.
+Lemma 2.5.
+Suppose 1 and 2 hold.
+Given $\epsilon>0$, let the penalty parameter $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$ in Eq.
+6 with $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$.
+If ${\mathbf{x}}^{*}\in{\mathcal{X}}$ satisfying $g_{+}({\mathbf{x}}^{*})\leq B_{g}$ is an $\epsilon$-stationary point of problem Eq.
+6, then $\mathbf{x}$$^{*}$ satisfies $g_{i}({\mathbf{x}}^{*})\leq 0$ for all $i\in[m]$, and is an $\epsilon$-KKT point of problem Eq.
 P.
-Proof 2.2.
-Since ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-5, there exist ${\boldsymbol{\gamma}}\in\partial f({\mathbf{x}}^{*})$ and ${\boldsymbol{\zeta}}\in\partial g_{+}({\mathbf{x}}^{*})$ such that
-(8) $\|{\boldsymbol{\gamma}}+\beta{\boldsymbol{\zeta}}\|\leq\epsilon.$
-We then derive that for any ${\mathbf{v}}\in\mathbb{R}^{d}$,
-(9) $F^{\prime}({\mathbf{x}}^{*};{\mathbf{v}})=f^{\prime}({\mathbf{x}}^{*};{\mathbf{v}})+\beta g_{+}^{\prime}({\mathbf{x}}^{*};{\mathbf{v}})\geq{\mathbf{v}}^{\top}{\boldsymbol{\gamma}}+\beta{\mathbf{v}}^{\top}{\boldsymbol{\zeta}}\geq-\epsilon\|{\mathbf{v}}\|,$
-where the first inequality follows from Eq. 7, and the second inequality is by Eq. 8.
-Second, we let ${\mathbf{x}}^{+}:=\mathrm{Proj}_{\Omega}({\mathbf{x}}^{*})$, and $\overline{{\mathbf{v}}}:={\mathbf{x}}^{+}-{\mathbf{x}}^{*}.$ For any $10<s<$ 1, we derive
-(10) $g_{+}({\mathbf{x}}^{*}+s\overline{{\mathbf{v}}})-g_{+}({\mathbf{x}}^{*})$ $\leq(1-s)g_{+}({\mathbf{x}}^{*})+sg_{+}({\mathbf{x}}^{*}+\overline{{\mathbf{v}}})-g_{+}({\mathbf{x}}^{*})$
-$=-sg_{+}({\mathbf{x}}^{*})\leq-{s\theta}{}\textnormal{dist}({\mathbf{x}}^{*},\Omega),$
-where the first inequality follows from the convexity of $g$, and the second inequality holds by 2 and $\textnormal{dist}({\mathbf{x}}^{*},\Omega)=0$ if $g({\mathbf{x}}^{*})\leq 0$.
-In addition, we have $f({\mathbf{x}}^{*}+s\overline{{\mathbf{v}}})-f({\mathbf{x}}^{*})\leq l_{f}s\|\overline{{\mathbf{v}}}\|=l_{f}s\textnormal{dist}({\mathbf{x}}^{*},\Omega)$.
-Combining this with Eq.
-10, we arrive at $\frac{1}{s}\left(F({\mathbf{x}}^{*}+s\overline{{\mathbf{v}}})-F({\mathbf{x}}^{*})\right)\leq-\beta\theta\textnormal{dist}({\mathbf{x}}^{*},\Omega)+l_{f}\textnormal{dist}({\mathbf{x}}^{*},\Omega)$.
-Letting $s\downarrow 0$, we obtain
+Proof 2.6.
+Since $\mathbf{x}$$^{*}$ is an $\epsilon$-stationary point of problem Eq.
+6, there exist ${\boldsymbol{\gamma}}\in\partial f({\mathbf{x}}^{*})$, $\{{\boldsymbol{\zeta}}_{g_{i}}\in\partial[g_{i}({\mathbf{x}}^{*})]_{+}\}_{i\in[m]}$, and ${\mathbf{u}}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*})$ such that
+(11) $\left\|\textstyle\beta\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}}+{\boldsymbol{\gamma}}\right\|\leq\epsilon.$
+We claim $g_{i}({\mathbf{x}}^{*})\leq 0$ for all $i\in[m]$.
+Otherwise, it holds $0<g_{+}({\mathbf{x}}^{*})\leq B_{g}$ and $m]\cap\{i:g_{i}({\mathbf{x}}^{*})\geq 0\}\neq\emptyset$.
+We then derive from Eq.
+11 and the triangle inequality that
+(12) $\epsilon$ $\geq{\beta}\left\|\textstyle\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}+\frac{{\mathbf{u}}}{\beta}\right\|-\|{\boldsymbol{\gamma}}\|\geq{\beta}\left\|\textstyle\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}+\frac{{\mathbf{u}}}{\beta}\right\|-l_{f}\geq\beta\theta-l_{f},$
+where the second inequality holds by 1(c), and the third one follows from Lemma 2.4 with ${\mathcal{J}}_{1}=[m]$.
+This contradicts the condition $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$.
+Hence $g_{i}({\mathbf{x}}^{*})\leq 0$ for all $i\in[m]$.
+Now we find a corresponding multiplier as follows.
+For each $i\in[m]$, if $g_{i}({\mathbf{x}}^{*})=0$, we obtain $\partial[g_{i}({\mathbf{x}}^{*})]_{+}={\mathrm{co}}(\partial g_{i}({\mathbf{x}}^{*})\cup\{\mathbf{0}\})$ by [18, Proposition 2.3.12] and the weak convexity of $g_{i}]_{+}$.
+Thus, there exists $\lambda_{i}\in[0,1]$ such that ${\boldsymbol{\zeta}}_{g_{i}}\in\lambda_{i}\partial g_{i}({\mathbf{x}}^{*})$.
+In this case, we set $t_{i}=\lambda_{i}$.
+If $g_{i}({\mathbf{x}}^{*})<0$, we have $\partial[g_{i}({\mathbf{x}}^{*})]_{+}=\{\mathbf{0}\}$ so ${\boldsymbol{\zeta}}_{g_{i}}=\mathbf{0}$, and we set $t_{i}=0$.
+For the constructed $\mathbf{t}$}, we have $t_{i}g_{i}({\mathbf{x}}^{*})=0,\forall\,i\in[m]$.
+In addition, $\beta\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}\in\sum_{i=1}^{m}\beta t_{i}\partial g_{i}({\mathbf{x}}^{*})$, and thus Eq.
+11 indicates $\textnormal{dist}\big{(}\mathbf{0},\partial f({\mathbf{x}}^{*})+\sum_{i=1}^{m}\beta t_{i}\partial g_{i}({\mathbf{x}}^{*})+{\mathcal{N}}_{\mathcal{X}}({\mathbf{x}}^{*})\big{)}\leq\epsilon$.
+Therefore $\mathbf{x}$$^{*}$ is an $\epsilon$-KKT point of problem Eq.
+P with the multiplier $\beta{\mathbf{t}}$.
+2.3 Exact penalization with a “smoothed” penalty approach
+In this subsection, we establish the relationship between models Eq. P and Eq. PP. We begin by showing that the function $F^{\nu}$ in Eq. PP is weakly convex.
+Lemma 2.7.
+Suppose 1 holds. For any $\nu>0$, the function $F^{\nu}$ defined in Eq. PP is $\rho$-weakly convex, where $\rho=\rho_{f}+\beta m\rho_{g}$.
+Proof 2.8.
+To obtain the desired result, we only need to prove $h^{\nu}({\mathbf{g}}({\mathbf{x}}))$ is $m\rho_{g}$-weakly convex. Notice that
+$h^{\nu}({\mathbf{g}}({\mathbf{x}}))+\frac{m\rho_{g}}{2}\|{\mathbf{x}}\|^{2}=\sum_{i=1}^{m}\max_{0\leq y\leq 1}\left[y\cdot g_{i}({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}\right].$
+Since $g_{i}$ is $\rho_{g}$-weakly convex for each $i\in[m]$, $y\cdot g_{i}({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex with respect to $\mathbf{x}$} for all $0\leq y\leq 1$.
+Thus by Danskin’s theorem [9], $h^{\nu}({\mathbf{g}}({\mathbf{x}}))+\frac{m\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex.
+Next, we introduce some notations and key properties of the Moreau envelope.
+Denote $F_{\frac{1}{2\rho}}$ and $F^{\nu}_{\frac{1}{2\rho}}$ as the Moreau envelope of $F$ and $F^{\nu}$ with parameter $\frac{1}{2\rho}$ on $\mathcal{X}$, respectively, where $\rho=\rho_{f}+\beta m\rho_{g}$.
+We let
+(13) $\overline{{\mathbf{x}}}^{*}=\arg\min_{{\mathbf{x}}\in\mathcal{X}}F({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2},\quad\widehat{{\mathbf{x}}}^{*}=\arg\min_{{\mathbf{x}}\in\mathcal{X}}F^{\nu}({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2}.$
+Then we have
+(14) $\nabla F_{\frac{1}{2\rho}}({\mathbf{x}}^{*})=2\rho({\mathbf{x}}^{*}-\overline{{\mathbf{x}}}^{*}),\ \nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{*})=2\rho({\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}^{*}),\text{ and}$
+(15) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\right)\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*}).$
+In addition, by [21, Eqn. (4.3)], it holds that
 […]
-Since $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$, the above inequality holds only if ${\mathbf{x}}^{*}\in\Omega$.
-Now, if $g({\mathbf{x}}^{*})=0$, we obtain from the convexity of $g$ that $\partial g_{+}({\mathbf{x}}^{*})={\mathrm{co}}\{\partial g({\mathbf{x}}^{*}),\mathbf{0}\}$ by [12, Proposition 2.3.12].
-Thus, there exists $0\leq t\leq 1$ such that ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$.
-Then ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-P with multiplier $\beta t$ by Eq.
-8 and ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$.
-Otherwise, i.e., if $g({\mathbf{x}}^{*})<0$, we have ${\boldsymbol{\zeta}}=\mathbf{0}$.
-It then yields that ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-P with multiplier 0 by Eq.
-8.
-The proof is then completed.
-Lemma 2.3.
-Suppose 1 and 3 hold.
-Given $\epsilon>0$, let $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$.
-If ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$ is an $\epsilon$-KKT point of problem Eq.
-5, it then holds ${\mathbf{x}}^{*}\in\Omega$, and ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-P.
-Proof 2.4.
-Since ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-5, there exits ${\boldsymbol{\gamma}}\in\partial f({\mathbf{x}}^{*})$ and ${\boldsymbol{\zeta}}\in\partial g_{+}({\mathbf{x}}^{*})$ such that Eq.
-8 holds.
-We claim $g_{+}({\mathbf{x}}^{*})=0$, i.e., $g({\mathbf{x}}^{*})\leq 0$.
-If the claim is not true, we have $g({\mathbf{x}}^{*})>0$ and then $\partial g_{+}({\mathbf{x}}^{*})=\partial g({\mathbf{x}}^{*})$.
-We then derive that
-(12) $\epsilon$ $\geq{}\|\beta{\boldsymbol{\zeta}}+{\boldsymbol{\gamma}}\|\geq{\beta}\|{\boldsymbol{\zeta}}\|-\|{\boldsymbol{\gamma}}\|\geq{\beta}\|{\boldsymbol{\zeta}}\|-l_{f}\geq\beta\theta-l_{f},$
-where the third inequality holds by 1(c), and the fourth one follows from 3 and ${\boldsymbol{\zeta}}\in\partial g_{+}({\mathbf{x}}^{*})=\partial g({\mathbf{x}}^{*})$.
-This contradicts to $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$.
-Hence $g_{+}({\mathbf{x}}^{*})=0$.
-Now, if $g({\mathbf{x}}^{*})=0$, we obtain $\partial g_{+}({\mathbf{x}}^{*})={\mathrm{co}}\{\partial g({\mathbf{x}}^{*}),\mathbf{0}\}$ from the weak convexity of $g$ and [12, Proposition 2.3.12].
-Thus, there exists $0\leq t\leq 1$ such that ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$.
-Then ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-P with multiplier $\beta t$ by Eq.
-8 and ${\boldsymbol{\zeta}}\in t\partial g({\mathbf{x}}^{*})$.
-Otherwise, i.e., if $g({\mathbf{x}}^{*})<0$, we have ${\boldsymbol{\zeta}}=\mathbf{0}$.
-It then yields that ${\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-P with multiplier 0 by Eq.
-8.
-The proof is then completed.
-The following corollary holds directly from Lemma 2.1 and Lemma 2.3.
-Corollary 2.5.
-Under 1, the following statements hold with $\beta>~{}\frac{l_{f}}{\theta}$.
-$\mathrm{(i)}$ Suppose 2 hold, and $g$ is convex.
-For all ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$, ${\mathbf{x}}^{*}\not\in\Omega$, and ${\mathbf{x}}^{*}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, it holds $\textnormal{dist}(\mathbf{0},\partial f({\mathbf{x}}^{*})+\beta\partial g_{+}({\mathbf{x}}^{*}))\geq\beta\theta-l_{f}$.
-$\mathrm{(ii)}$ Suppose 3 hold.
-For all ${\mathbf{x}}^{*}$ satisfying $g({\mathbf{x}}^{*})\leq B_{g}$, ${\mathbf{x}}^{*}\not\in\Omega$, and ${\mathbf{x}}^{*}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, it holds $\textnormal{dist}(\mathbf{0},\partial f({\mathbf{x}}^{*})+\beta\partial g_{+}({\mathbf{x}}^{*}))\geq\beta\theta-l_{f}$.
-2.2 Exact penalization with a “smoothed” penalty approach
-In this subsection, we establish the relationship between models Eq. P and Eq. PP. We begin by showing that the function $F^{\nu}$ in (PP) is weakly convex.
-Lemma 2.6.
-Suppose 1 holds. For any given $\nu>0$, the function $F^{\nu}$ defined in Eq. PP is $\rho$-weakly convex, where $\rho=\rho_{f}+\beta\rho_{g}$.
-Proof 2.7.
-To obtain the desired result, we only need to prove $h^{\nu}(g({\mathbf{x}}))$ is $\rho_{g}$-weakly convex. Notice that
-$h^{\nu}(g({\mathbf{x}}))+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}=\max_{0\leq y\leq 1}y\cdot g({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}.$
-Since $g$ is $\rho_{g}$-weakly convex, $y\cdot g({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex with respect to ${\mathbf{x}}$ for all $0\leq y\leq 1$.
-Thus by Danskin’s theorem [8], $h^{\nu}(g({\mathbf{x}}))+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex.
-Next, we introduce some notations and key properties of the Moreau envelope.
-Denote $F_{\frac{1}{2\rho}}$ and $F^{\nu}_{\frac{1}{2\rho}}$ as the Moreau envelope of $F$ and $F^{\nu}$ with parameter $\frac{1}{2\rho}$, respectively.
-We let
-(13) $\overline{\mathbf{x}}^{*}=\arg\min_{{\mathbf{x}}}F({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2},\quad\widehat{{\mathbf{x}}}^{*}=\arg\min_{{\mathbf{x}}}F^{\nu}({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2}.$
-We then have
-(14) $\|\nabla F_{\frac{1}{2\rho}}({\mathbf{x}}^{*})\|=2\rho\|\overline{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|,\ \|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{*})\|=2\rho\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|,\text{and}$
-(15) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\mathrm{Proj}_{[0,1]}\left(\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\right)\partial g\left(\widehat{\mathbf{x}}^{*}\right)+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right).$
-In addition, by [13, Eqn. (4.3)], it holds that
-(16) $F(\overline{{\mathbf{x}}}^{*})\leq F({\mathbf{x}}^{*}),\text{ and }\textnormal{dist}(\mathbf{0},\partial F(\overline{{\mathbf{x}}}^{*}))\leq\big{\|}\nabla F_{\frac{1}{2\rho}}({\mathbf{x}}^{*})\big{\|}.$
-We now present our main results to demonstrate the effectiveness of solving a smoothed exact penalty problem for a deterministic or stochastic case.
+We now present our main results to demonstrate the effectiveness of solving a smoothed exact penalty problem for both deterministic and stochastic settings.
 Theorem 2.9.
-Suppose that 1 holds. Given $\epsilon>0$, let
-(17) $\small\beta>\frac{1}{\theta}(l_{f}+\epsilon),\ 0<\overline{\epsilon}\leq\epsilon\min\left\{\frac{\beta\theta-l_{f}}{2B_{g}},\frac{\beta\theta-l_{f}}{\beta\theta},2\rho\right\},\text{ and }0<\nu\leq\min\left\{B_{g},\frac{\epsilon}{\beta},\frac{\epsilon}{2}\right\}.$
+Suppose that 1 holds and 2 holds with $B_{g}>m\rho_{g}D^{2}$. Given $\epsilon>0$, let $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$, $\rho=\rho_{f}+\beta m\rho_{g}$,
+(17) $\beta>\max\left\{\frac{1}{\theta}(l_{f}+\epsilon),\frac{2(B_{f}+\rho_{f}D^{2})}{B_{g}-m\rho_{g}D^{2}}\right\},0<\overline{\epsilon}\leq\epsilon\min\left\{\frac{\beta\theta-l_{f}}{2B_{g}},\frac{\beta\theta-l_{f}}{m\beta l_{g}+\beta\theta-l_{f}},2\rho\right\},$
+$\text{ and }0<\nu\leq\min\left\{B_{g},\frac{\epsilon}{\beta},\frac{B_{g}-m\rho_{g}D^{2}}{2m},\frac{\epsilon}{2}\right\}.$
 If $\mathbf{x}$$^{*}$ is a near-$\overline{\epsilon}$ stationary point of problem Eq.
-PP with $t=\frac{1}{2\rho}$ (see Definition 1.2$\mathrm{(a)}$) deterministically or in expectation, then $\|{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}^{*}\|\leq\epsilon$ holds deterministically or in expectation, where $\widehat{{\mathbf{x}}}^{*}$ is defined in Eq.
+PP with $t=\frac{1}{2\rho}=\frac{1}{2(\rho_{f}+\beta m\rho_{g})}$ deterministically or in expectation, then $\|{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}^{*}\|\leq\epsilon$ holds deterministically or in expectation, where $\widehat{{\mathbf{x}}}^{*}$ is defined in Eq.
 13.
-In addition, for each of the following three cases, $\widehat{{\mathbf{x}}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
+In addition, $\widehat{{\mathbf{x}}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
 P deterministically or in expectation.
+Proof 2.10.
+We only need to prove the theorem in the expectation case, when $\mathbf{x}$$^{*}$ is a random vector. For the deterministic case, one can simply drop the expectation notation $\mathbb{E}$.
+First, from Definition 1.3 and Eq.
+14, it follows $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|^{2}\right]\leq~\overline{\epsilon}^{2}$ by the near-$\overline{\epsilon}$ stationarity of $\mathbf{x}$$^{*}$ in expectation.
+Because $\overline{\epsilon}\leq 2\rho\epsilon$, it follows that
+(18) $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|\right]\leq\overline{\epsilon}\text{ and }\mathbb{E}\left[\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|\right]\leq\epsilon.$
+Second, we define two index sets: $\mathcal{J}_{2}=\{i:g_{i}(\widehat{\mathbf{x}}^{*})/\nu\leq 1\}$ and $\mathcal{J}_{1}=[m]\setminus\mathcal{J}_{2}$.
+Construct the vector $\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}}$ by setting $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}=\beta\mathrm{Proj}_{[0,1]}(g_{i}(\widehat{\mathbf{x}}^{*})/\nu)\geq 0$ for $i\in\mathcal{J}_{2}$ and $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}=0$ for $i\in\mathcal{J}_{1}$.
+We next bound $|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})|$ for each $i$.
+For $i\in\mathcal{J}_{1}$, we directly have $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})=0$.
+For $i\in\mathcal{J}_{2}$, if $g_{i}(\widehat{\mathbf{x}}^{*})\leq 0$, then $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}=0$ and $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})=0$; if $g_{i}(\widehat{\mathbf{x}}^{*})>0$, we have from $g_{i}(\widehat{\mathbf{x}}^{*})/\nu\leq 1$ that $g_{i}(\widehat{\mathbf{x}}^{*})\leq\nu\leq\epsilon/2$, and thus $|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})|=(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})\leq\beta g_{i}(\widehat{\mathbf{x}}^{*})\leq\beta\nu\leq\epsilon.$ Hence,
+(19) $|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})|\leq\epsilon,\forall\,i\in[m].$
+Third, we show $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$. By Eq. 15 and the definition of ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$, we have
+(20) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\partial\sum_{i\in\mathcal{J}_{1}}[g_{i}\left(\widehat{\mathbf{x}}^{*}\right)]_{+}+\sum_{i\in\mathcal{J}_{2}}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right).$
+By 1, there exists $\mathbf{x}_{\mathrm{feas}}\in{\mathcal{X}}\cap\{{\mathbf{x}}:{\mathbf{g}}({\mathbf{x}})\leq 0\}$. Hence, from Eq. 13, it holds
+(21) $f\left(\widehat{\mathbf{x}}^{*}\right)+\beta h^{\nu}\left({\mathbf{g}}\left(\widehat{\mathbf{x}}^{*}\right)\right)+\rho\left\|\widehat{\mathbf{x}}^{*}-\mathbf{x}^{*}\right\|^{2}$
+$\leq$ $f\left(\mathbf{x}_{\mathrm{feas}}\right)+\beta h^{\nu}\left({\mathbf{g}}\left(\mathbf{x}_{\mathrm{feas}}\right)\right)+\rho\left\|\mathbf{x}_{\mathrm{feas}}-\mathbf{x}^{*}\right\|^{2}=f\left(\mathbf{x}_{\mathrm{feas}}\right)+\rho\left\|\mathbf{x}_{\mathrm{feas}}-\mathbf{x}^{*}\right\|^{2}.$
+Thus, $f(\widehat{{\mathbf{x}}}^{*})+\beta h^{\nu}({\mathbf{g}}(\widehat{{\mathbf{x}}}^{*}))+\rho\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|^{2}\leq f({\mathbf{x}}_{\mathrm{feas}})+\rho D^{2}$ by 1(a).
+In addition, we have $\beta\geq\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-m\nu-m\rho_{g}D^{2}}$ by $\beta>\frac{2(B_{f}+\rho_{f}D^{2})}{B_{g}-m\rho_{g}D^{2}}$ and $\nu\leq\frac{B_{g}-m\rho_{g}D^{2}}{2m}$.
+Now further by $m\rho_{g}D^{2}<B_{g}$, $\rho=\rho_{f}+\beta m\rho_{g}$, and Eq.
+1, it follows that
+(22) $\textstyle h^{\nu}({\mathbf{g}}(\widehat{{\mathbf{x}}}^{*}))\leq\frac{1}{\beta}\Bigl{(}f({\mathbf{x}}_{\mathrm{feas}})-f(\widehat{{\mathbf{x}}}^{*})+\rho D^{2}\Bigr{)}\leq\frac{B_{f}+\rho D^{2}}{\beta}\leq B_{g}-m\nu.$
+This implies $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq h^{\nu}({\mathbf{g}}(\widehat{{\mathbf{x}}}^{*}))+m\nu\leq B_{g}$ by the definition of $h^{\nu}$.
+We now discuss two events: ${\mathcal{A}}_{1}$ as the event of $\mathcal{J}_{1}=\emptyset$ and ${\mathcal{A}}_{2}$ as the event of $\mathcal{J}_{1}\neq\emptyset$.
+Let $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$ denote their probabilities.
+When ${\mathcal{A}}_{1}$ happens, we have $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\sum_{i=1}^{m}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|$ by Eq.
+20.
+Hence,
+(23) $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\sum_{i=1}^{m}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|,$
+$g_{i}(\widehat{\mathbf{x}}^{*})\leq\frac{\epsilon}{2},\text{ and }|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}\left(\widehat{\mathbf{x}}^{*}\right)|\leq\epsilon,\forall\,i\in[m].$
+When ${\mathcal{A}}_{2}$ happens, we have $\mathcal{J}_{1}\cap\{i:g_{i}(\widehat{\mathbf{x}}^{*})\geq 0\}\neq\emptyset$, and $(\lambda_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})\geq 0$ for all $i\in\mathcal{J}_{2}$.
+Using Lemma 2.4 and 1(c), we obtain that
+(24) $2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|$
+[…]
+Next, we combine events $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. From Eq. 24 and by the definition of $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$, we have
+[…]
+Thus $\mathrm{Prob}_{2}\leq\frac{\overline{\epsilon}}{\beta\theta-l_{f}}.$ Splitting the integration for full expectation to ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ gives
+(25) $\mathbb{E}\left[\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\sum_{i=1}^{m}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})\right)\right]$
+[…]
+where the last inequality uses $\overline{\epsilon}\leq\frac{\beta\theta-l_{f}}{m\beta l_{g}+\beta\theta-l_{f}}\epsilon$. In addition, we have
+$\mathbb{E}[g_{i}(\widehat{\mathbf{x}}^{*})]{\leq}\nu\mathrm{Prob}_{1}+B_{g}\mathrm{Prob}_{2}\leq\nu+\frac{B_{g}\overline{\epsilon}}{\beta\theta-l_{f}}\leq\epsilon,\forall\,i\in[m],$
+where we have used $\nu\leq\frac{\epsilon}{2}$ and $\overline{\epsilon}\leq\frac{\beta\theta-l_{f}}{2B_{g}}$.
+Therefore, by recalling Eq.
+19, we have that $\widehat{\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
+P in expectation and complete the proof.
+Remark 2.11.
+Theorem 2.9 relies on several key assumptions, which we clarify below to highlight their implications and scope:
   • $\mathrm{(i)}$
-[…]
+    The assumed near-$\bar{\epsilon}$ stationary point $\mathbf{x}$$^{*}$ of the penalized problem Eq. PP can be found by 3S-Econ introduced in the next section.
   • $\mathrm{(ii)}$
-[…]
+    The required Slater-type CQ with $B_{g}>m\rho_{g}D^{2}$ will reduce to the classical Slater’s condition under convex case ($\rho_{g}=0$), as formally established in Proposition 2.1(i).
+The condition is milder than the strong feasibility condition in Eq.
+10; see Proposition 2.1(ii).
+It could potentially be further relaxed when $\mathbf{g}$} is deterministic; see also Remark 3.12.
+Additionally, the parameter $B_{g}$, which serves as an upper bound for $\sum_{i\in[m]}[g_{i}({\mathbf{x}})]_{+}$, can scale linearly with $m$.
   • $\mathrm{(iii)}$
-[…]
-Proof 2.9.
-We only need to the prove the theorem in the expectation sense. For the deterministic case, one can simply drop the expectation notation $\mathbb{E}$.
-First, according to the definition of ${\mathbf{x}}^{*}$efinition 1.2 gives $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|^{2}\right]\leq~{}\overline{\epsilon}^{2}.$ Because $\overline{\epsilon}\leq 2\rho\epsilon$, if follows that
-(18) $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|\right]\leq\overline{\epsilon}\text{ and }\mathbb{E}\left[\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|\right]\leq\epsilon.$
-Second, we discuss two events: ${\mathcal{A}}_{1}$ as the case of $\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\leq 1$ and ${\mathcal{A}}_{2}$ as the case of $\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}>1$.
-Let $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$ denote their probabilities.
-Also, we let $\lambda_{\widehat{\mathbf{x}}^{*}}$ be defined as $\lambda_{\widehat{\mathbf{x}}^{*}}=\beta\mathrm{Proj}_{[0,1]}(\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu})\geq 0$ when ${\mathcal{A}}_{1}$ happens and $\lambda_{\widehat{\mathbf{x}}^{*}}=0$ when ${\mathcal{A}}_{2}$ happens.
-For event ${\mathcal{A}}_{1}$, we have $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\lambda_{\widehat{\mathbf{x}}^{*}}\partial g\left(\widehat{\mathbf{x}}^{*}\right)\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|$ by Eq.
-15 and triangle inequality.
-Also, because $\frac{g\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\leq 1$, we have $g(\widehat{\mathbf{x}}^{*})\leq\nu\leq\frac{\epsilon}{2}.$ Moreover, if it holds $g(\widehat{\mathbf{x}}^{*})\leq 0$, then $\lambda_{\widehat{\mathbf{x}}^{*}}=0$ by its definition, and $|\lambda_{\widehat{\mathbf{x}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)|=0\leq\epsilon$; if $g(\widehat{\mathbf{x}}^{*})>0$, then $|\lambda_{\widehat{\mathbf{x}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)|=\lambda_{\widehat{\mathbf{x}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)\leq\beta g\left(\widehat{\mathbf{x}}^{*}\right)\leq\beta\nu\leq\epsilon$.
-Hence, we obtain
-(19) $\textstyle\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\lambda_{\widehat{{\mathbf{x}}}^{*}}\partial g\left(\widehat{\mathbf{x}}^{*}\right)\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|,\,\,g(\widehat{\mathbf{x}}^{*})\leq\nu\leq\frac{\epsilon}{2},\text{ and }|\lambda_{\widehat{{\mathbf{x}}}^{*}}g\left(\widehat{\mathbf{x}}^{*}\right)|\leq\epsilon.$
-For event ${\mathcal{A}}_{2}$, we have from $g\left(\widehat{\mathbf{x}}^{*}\right)>\nu$ and Eq. 15 that
-(20) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\partial g_{+}\left(\widehat{\mathbf{x}}^{*}\right)+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right).$
-Thus $\widehat{\mathbf{x}}^{*}$ satisfies the optimality condition of the left minimization problem in Eq.
-13, and we obtain $\widehat{\mathbf{x}}^{*}=\overline{\mathbf{x}}^{*}$.
-Below we show $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$ under event $\mathcal{A}_{2}$ for all the three cases (i)–(iii) in the statement of Theorem 2.8.
-For case (i), there exists ${\mathbf{x}}_{\mathrm{feas}}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)\cap\Omega$ by 1(a). From Eq. 13, we have
-$f(\overline{{\mathbf{x}}}^{*})+\beta g_{+}(\overline{{\mathbf{x}}}^{*})+\rho\|\overline{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|^{2}$ $\leq f({\mathbf{x}}_{\mathrm{feas}})+\beta g_{+}({\mathbf{x}}_{\mathrm{feas}})+\rho\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2}$
-(21) $=f({\mathbf{x}}_{\mathrm{feas}})+\rho\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2}.$
-Now because $\beta\geq\frac{B_{f}+\rho_{f}D^{2}}{B_{g}}$, $\rho=\rho_{f}+\beta\rho_{g}=\rho_{f}$, and $\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2}\leq D^{2}$ from 1(a), we arrive at
-$g_{+}(\widehat{{\mathbf{x}}}^{*})=g_{+}(\overline{{\mathbf{x}}}^{*})\leq\frac{1}{\beta}(f({\mathbf{x}}_{\mathrm{feas}})-f(\overline{{\mathbf{x}}}^{*})+\rho_{f}\|{\mathbf{x}}_{\mathrm{feas}}-{\mathbf{x}}^{*}\|^{2})\leq\frac{B_{f}+\rho_{f}D^{2}}{\beta}\leq B_{g}.$
-For case (ii), we still have Eq.
-21.
-Thus, $f(\overline{{\mathbf{x}}}^{*})+\beta g_{+}(\overline{{\mathbf{x}}}^{*})+\rho\|\overline{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|^{2}\leq f({\mathbf{x}}_{\mathrm{feas}})+\rho D^{2}$ by 1(a).
-Now by $\rho_{g}D^{2}<B_{g}$, $\beta\geq\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-\rho_{g}D^{2}}$, and $\rho=\rho_{f}+\beta\rho_{g}$, it follows
-$g_{+}(\widehat{{\mathbf{x}}}^{*})=g_{+}(\overline{{\mathbf{x}}}^{*})\leq\frac{1}{\beta}\Bigl{(}f({\mathbf{x}}_{\mathrm{feas}})-f(\overline{{\mathbf{x}}}^{*})+\rho D^{2}\Bigr{)}\leq\frac{B_{f}+\rho D^{2}}{\beta}\leq B_{g}.$
-For case (iii), we have $f(\overline{{\mathbf{x}}}^{*})+\beta g_{+}(\overline{{\mathbf{x}}}^{*})+\rho\|\overline{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|^{2}\leq f({{\mathbf{x}}}^{*})+\beta g_{+}({{\mathbf{x}}}^{*})$ by the definition of $\overline{\mathbf{x}}^{*}$.
-Using $|f(\overline{{\mathbf{x}}}^{*})-f({{\mathbf{x}}}^{*})|\leq B_{f}$, $\beta\geq\frac{2B_{f}}{B_{g}}$, and $g_{+}({{\mathbf{x}}}^{*})\leq\frac{B_{g}}{2}$, we derive that $g_{+}(\widehat{{\mathbf{x}}}^{*})=g_{+}(\overline{{\mathbf{x}}}^{*})\leq\frac{B_{f}}{\beta}+g_{+}({{\mathbf{x}}}^{*})\leq B_{g}$.
-Hence, for all three cases under event $\mathcal{A}_{2}$, we have $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$. Thus, by 1(c) and $\lambda_{\widehat{\mathbf{x}}^{*}}=0$, we have
-(22) $\textnormal{dist}\left(\mathbf{0},\partial f(\widehat{\mathbf{x}}^{*})+\lambda_{\widehat{\mathbf{x}}^{*}}\partial g(\widehat{\mathbf{x}}^{*})\right)\leq l_{f},\,\,g(\widehat{\mathbf{x}}^{*})\leq B_{g},\text{ and }\lambda_{\widehat{\mathbf{x}}^{*}}g(\widehat{\mathbf{x}}^{*})=0.$
-Furthermore, $g(\widehat{\mathbf{x}}^{*})>\nu$ implies $\widehat{\mathbf{x}}^{*}\notin\Omega$.
-Hence by Corollary 2.5 and $g(\widehat{\mathbf{x}}^{*})\leq B_{g}$, it holds $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\partial g_{+}(\widehat{\mathbf{x}}^{*})\right)\geq\beta\theta-l_{f}$.
-This yields
-[…]
-Now, we combine events $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. From Eq. 19 and Eq. 23, and by the definition of $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$, we have
-[…]
-Hence from $\overline{\epsilon}\leq\frac{\beta\theta-l_{f}}{\beta\theta}\epsilon$, it follows that $\mathrm{Prob}_{2}\leq\frac{\overline{\epsilon}}{\beta\theta-l_{f}}\leq\frac{\epsilon}{\beta\theta}.$ We finally compute expectations based on these probabilities.
-It holds
-[…]
-$\leq\overline{\epsilon}+\frac{\overline{\epsilon}l_{f}}{\beta\theta-l_{f}}\leq\epsilon.$
-In addition, we have $\mathbb{E}[g(\widehat{\mathbf{x}}^{*})]{\leq}\nu\mathrm{Prob}_{1}+B_{g}\mathrm{Prob}_{2}\leq\nu+\frac{B_{g}\overline{\epsilon}}{\beta\theta-l_{f}}\leq\epsilon,\text{ and }|\lambda_{\widehat{\mathbf{x}}^{*}}g(\widehat{\mathbf{x}}^{*})|\leq\epsilon,$ where we have used Eq.
-19Eq.
-22, $\nu\leq\frac{\epsilon}{2}$ and $\overline{\epsilon}\leq\frac{\beta\theta-l_{f}}{2B_{g}}$.
-Therefore, $\widehat{\mathbf{x}}^{*}$ is an $\epsilon$-KKT point of problem Eq.
-P in expectation.
-This completes the proof.
-3 A single-loop SPIDER-type stochastic subgradient algorithm
-In this section, we present our algorithm for solving problem Eq.
+    The values of parameters in Eq.
+17 depend on several constants, which may be unknown in general.
+Nevertheless, we show in Section 4.1 that these constants can be computed for certain applications, mitigating uncertainty and facilitating precise complexity assessments.
+3 A SPIDER-type stochastic subgradient algorithm
+In this section, we present our algorithm, termed 3S-Econ, for solving problem Eq.
 P.
@@
 The algorithm framework is outlined in Section 3.1.
-In Section 3.2, we prove that in expectation, the algorithm can find a near-$\epsilon$ stationary point of problem Eq.
+In Section 3.2, we prove that in expectation, 3S-Econ can find a near-$\epsilon$ stationary point of problem Eq.
 PP.
@@
 3.1 Algorithm framework
-We employ the SPIDER technique [16] to obtain an estimator $u^{(k)}$ of $g$ at the $k$-th iteration.
-Specifically, every $q$ iterations, we draw a big batch of $S_{1}$ samples from $\mathbb{P}_{g}$, while for other iterations, a smaller batch of $S_{2}$ samples is drawn.
+We employ the SPIDER technique [24] to obtain an estimator ${\mathbf{u}}^{(k)}\in\mathbb{R}^{m}$ of $\mathbf{g}$} at the $k$-th iteration.
+Specifically, every $q$ iterations, we draw a big batch of $S_{1}$ samples from $\mathbb{P}_{{\mathbf{g}}}$, while for other iterations, a smaller batch of $S_{2}$ samples is drawn.
 The estimator is set as
-(25) $u^{(k)}=\left\{\begin{array}[]{lll}g({\mathbf{x}}^{(k)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{1},&\text{ if }\mathrm{mod}(k,q)=0,\\[2.84526pt] u^{(k-1)}+g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k-1)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{2},&\text{ otherwise},\end{array}\right.$
-where $\mathcal{B}_{k}$ contains identically independent samples from $\mathbb{P}_{g}$, and $g(\mathbf{x},\mathcal{B})$ is defined in Eq.
-6.
-With $u^{(k)}$, we then take a sample subgradient of $f$ and $g$ at ${\mathbf{x}}^{(k)}$ and perform the update in Eq.
-1.
+(26) ${\mathbf{u}}^{(k)}=\left\{\begin{array}[]{lll}{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{1},&\text{ if }\mathrm{mod}(k,q)=0,\\[2.84526pt] {\mathbf{u}}^{(k-1)}+{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{2},&\text{ otherwise},\end{array}\right$.}
+where ${\mathbf{u}}^{(k)}=[u_{1}^{(k)},\ldots,u_{m}^{(k)}]^{\top}$, $\mathcal{B}_{k}$ contains identically independent samples from $\mathbb{P}_{{\mathbf{g}}}$, and $g_{i}(\mathbf{x},\mathcal{B})$ is defined in Eq.
+7.
+With ${\mathbf{u}}^{(k)}$, we then take a sample subgradient of $f$ and $\{g_{i}\}_{i\in[m]}$ at ${\mathbf{x}}^{(k)}$ and perform the update in Eq.
+2.
 The pseudocode is given in Algorithm 1.
-Algorithm 1 A SPIDER-type Stochastic Subgradient algorithm
-1:  Input: initial point ${\mathbf{x}}^{(0)}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, $\beta>0$, $q>0$, $S_{1},S_{2}>0$, and $K>0$.
+As detailed in Theorem 3.6, for a given error tolerance $\bar{\epsilon}>0$, we set the penalty parameter $\beta$ to a sufficiently large predefined constant, the smoothing parameter $\nu=O(\bar{\epsilon})$, the step size $\alpha_{k}=O(\bar{\epsilon}^{2})$, and $q=S_{2}=\lceil\sqrt{S_{1}}\rceil$, with $S_{1}=O(\bar{\epsilon}^{-4})$ in the generic stochastic case or $S_{1}=N$ for the finite-sum case.
+Algorithm 1 3S-Econ: A SPIDER-type Stochastic Subgradient algorithm for solving Expectation constrained problem (P)
+1:Input: ${\mathbf{x}}^{(0)}\in\mathcal{X}$, $\beta>0$, $\nu>0$, and positive intergers $q$, $S_{1},S_{2}$, and $K$.
 2:for $k=0,1,\cdots,Kq-1$ do
-3:     Set $u^{(k)}$ by Eq. 25.
-4:     Choose $\alpha_{k}>0$, generate a (stochastic) subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and a (stochastic) subgradient $\boldsymbol{\zeta}_{g}^{(k)}$ of $g$.
+3:  Sample ${\mathcal{B}}_{k}$ and set ${\mathbf{u}}^{(k)}$ by Eq. 26.
+4:  Choose $\alpha_{k}>0$, generate a (stochastic) subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and (stochastic) subgradient $\boldsymbol{\zeta}_{g_{i}}^{(k)}$ of $g_{i}$ for each $i\in[m]$ that are independent of ${\mathcal{B}}_{k}$.
 5:  Update $\mathbf{x}$} by Eq. 2.
@@
 For the convenience of analysis, we denote
-(26) $\boldsymbol{\zeta}_{F^{\nu}}^{(k)}=\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\text{, and }$
-${\mathbf{w}}^{(k)}=\left(\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g}^{(k)}.$
+(27) $\boldsymbol{\zeta}_{F^{\nu}}^{(k)}=\boldsymbol{\zeta}_{f}^{(k)}+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\text{, and }$
+${\mathbf{w}}^{(k)}=\sum_{i=1}^{m}\left(\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}.$
 Then the update in Eq. 2 can be rewritten as
-(27) ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right).$
-Throughout this section, we make the following assumption, where for brevity of notations, we let $\mathbb{E}_{k}[\cdot]:=\mathbb{E}\left[\cdot\mid{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\right]$, representing the conditional expectation given $\{{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\}$ for any $k\geq 0$.
+(28) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathbf{x}}^{(k+1)}=\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right)\right).$
+Throughout this section, we make the following assumption, where for brevity of notation, we let $\mathbb{E}_{k}[\cdot]:=\mathbb{E}\left[\cdot\mid{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\right]$, representing the conditional expectation given $\{{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\}$ for any $k\geq 0$.
 In addition, $\mathrm{Var}(\boldsymbol{\zeta}):=\mathbb{E}[\|\boldsymbol{\zeta}-\mathbb{E}[\boldsymbol{\zeta}]\|^{2}]$ for a random vector $\boldsymbol{\zeta}$.
@@
   • $(a)$
-    Unbiased function estimate and Lipschitz continuity of $g$: For all $k\geq 0$, the samples in $\mathcal{B}_{k}$ are mutually independent.
-It holds $\mathbb{E}_{k}[g({\mathbf{x}}^{(k)},\xi_{g})]=g({\mathbf{x}}^{(k)})$ and $\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k)})\right|^{2}\right]\leq\sigma^{2}$ for any $\xi_{g}\in\mathcal{B}_{k}$ and some $\sigma>0$.
-Moreover, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathrm{dom}}(g)$, it holds $\mathbb{E}\left[\left|g({\mathbf{x}},\xi_{g})-g({\mathbf{y}},\xi_{g})\right|^{2}\right]\leq L_{g}^{2}\|{\mathbf{x}}-{\mathbf{y}}\|^{2}.$
+    Unbiased function estimate and Lipschitz continuity of $\mathbf{g}$}: For all $k\geq 0$, the samples in $\mathcal{B}_{k}$ are mutually independent.
+It holds $\mathbb{E}_{k}[{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})]={\mathbf{g}}({\mathbf{x}}^{(k)})$ and $\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq\sigma^{2}$ for any $\xi_{{\mathbf{g}}}\in\mathcal{B}_{k}$ and some $\sigma>0$.
+Moreover, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{X}}$, it holds $\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{y}},\xi_{{\mathbf{g}}})\right\|^{2}\right]\leq L_{g}^{2}\|{\mathbf{x}}-{\mathbf{y}}\|^{2}.$
   • $(b)$
-    Unbiased Stochastic Subgradients: For each $k$, $\mathbb{E}_{k}[\boldsymbol{\zeta}_{f}^{(k)}]\in\partial f({\mathbf{x}}^{(k)})$ and $\mathbb{E}_{k}[\boldsymbol{\zeta}_{g}^{(k)}]\in\partial g({\mathbf{x}}^{(k)})$.
-In addition, there are $\sigma_{1}\geq 0$ and $\sigma_{2}\geq 0$ such that $\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})\leq\sigma_{1}^{2}$ and $\mathrm{Var}(\boldsymbol{\zeta}_{g}^{(k)})\leq\sigma_{2}^{2}$.
-By 4(b), it follows that $\mathbb{E}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$.
-3.2 Convergence analysis and complexity results
-In this subsection, we present the convergence analysis of Algorithm 1. First, we derive a bound on the difference between $u^{(k)}$ and $g({\mathbf{x}}^{(k)})$.
+    Unbiased Stochastic Subgradients: For each $k$, $\mathbb{E}_{k}[\boldsymbol{\zeta}_{f}^{(k)}]\in\partial f({\mathbf{x}}^{(k)})$ and $\mathbb{E}_{k}[\boldsymbol{\zeta}_{g_{i}}^{(k)}]\in\partial g_{i}({\mathbf{x}}^{(k)})$ for each $i\in[m]$.
+In addition, there are $\sigma_{1}\geq 0$ and $\sigma_{2}\geq 0$ such that $\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})\leq\sigma_{1}^{2}$ and $\mathrm{Var}(\boldsymbol{\zeta}_{g_{i}}^{(k)})\leq\sigma_{2}^{2}$ for each $i\in[m]$.
+By 3(b), it follows that $\mathbb{E}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$.
+3.2 Convergence analysis
+In this subsection, we present the convergence analysis of Algorithm 1. First, we derive a bound on the difference between ${\mathbf{u}}^{(k)}$ and ${\mathbf{g}}({\mathbf{x}}^{(k)})$.
 Lemma 3.1.
-Under 1 and 4, let $\{{\mathbf{x}}^{(k)},u^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1.
-Define $r_{g}^{(k)}:=\mathbb{E}\left[|u^{(k)}-g({\mathbf{x}}^{(k)})|^{2}\right]$.
-Then, for all $k\geq 0$, it holds $r_{g}^{(k)}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
+Under 1 and 3, let $\{{\mathbf{x}}^{(k)},{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1.
+Define $r_{g}^{(k)}:=\mathbb{E}\left[\|{\mathbf{u}}^{(k)}-{\mathbf{g}}({\mathbf{x}}^{(k)})\|^{2}\right]$.
+Then, for all $k\geq 0$, it holds $r_{g}^{(k)}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}m^{2}\sigma_{2}^{2}+\beta^{2}m^{2}l_{g}^{2})$.
 Proof 3.2.
-Case (i): When $\mathrm{mod}(k,q)=0$,  Eq. 25 and 4(a) give
-(28) $r_{g}^{(k)}=\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k)})\right|^{2}\right]\leq\frac{\sigma^{2}}{S_{1}}.$
+Case (i): When $\mathrm{mod}(k,q)=0$,  Eq. 26 and 3(a) give
+(29) $r_{g}^{(k)}=\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq\frac{\sigma^{2}}{S_{1}}.$
 Case (ii): When $\mathrm{mod}(k,q)>0$, it follows that
 […]
-$=r_{g}^{(k-1)}+\mathbb{E}\left[\left|g({\mathbf{x}}^{(k-1)})+g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k-1)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k)})\right|^{2}\right]$
-$=r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{g}\in\mathcal{B}_{k}}\mathbb{E}\left[\left|g({\mathbf{x}}^{(k-1)})+g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k-1)},\xi_{g})-g({\mathbf{x}}^{(k)})\right|^{2}\right]$
-$\leq r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{g}\in\mathcal{B}_{k}}\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k-1)},\xi_{g})\right|^{2}\right]$
-$\leq r_{g}^{(k-1)}+\frac{L_{g}^{2}}{S_{2}}\mathbb{E}\left[\|{\mathbf{x}}^{(k)}-{\mathbf{x}}^{(k-1)}\|^{2}\right]\leq r_{g}^{(k-1)}+\frac{\alpha_{k-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}},$
-where the second equality holds due to the unbiasedness of $g({\mathbf{x}},\xi_{g})$, the third equality follows from the mutual independence of samples in $\mathcal{B}_{k}$, the second inequality follows by 4(a), and the last one uses
-(30) $\mathbb{E}\left[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|^{2}\right]=\textstyle\alpha_{k}^{2}\mathbb{E}\left[\left\|\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\right)\right\|^{2}\right]$
-$\leq 2\alpha_{k}^{2}\left(\mathbb{E}\big{[}\|\boldsymbol{\zeta}_{f}^{(k)}\|^{2}+\beta^{2}\|\boldsymbol{\zeta}_{g}^{(k)}\|^{2}\big{]}\right)$
-$\leq 2\alpha_{k}^{2}\left(\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})+\|\mathbb{E}[\boldsymbol{\zeta}_{f}^{(k)}]\|^{2}+\beta^{2}\mathrm{Var}(\boldsymbol{\zeta}_{g}^{(k)})+\beta^{2}\|\mathbb{E}[\boldsymbol{\zeta}_{g}^{(k)}]\|^{2}\right)$
-$\leq 2\alpha_{k}^{2}(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})=\alpha_{k}^{2}L_{F}^{2}.$
-Since $\mathrm{mod}(k,q)>0$, there exist nonnegative integers $t$ and $s<q$ such that $k=tq+s$. Summing Eq. 29 for all $k=tq+1,\ldots,tq+s-1,tq+s$, and applying Eq. 28, we have
+$=r_{g}^{(k-1)}+\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k-1)})+{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]$
+$=r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{{\mathbf{g}}}\in\mathcal{B}_{k}}\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k-1)})+{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]$
+$\leq r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{{\mathbf{g}}}\in\mathcal{B}_{k}}\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\xi_{{\mathbf{g}}})\right\|^{2}\right]$
+(30) $\leq r_{g}^{(k-1)}+\frac{L_{g}^{2}}{S_{2}}\mathbb{E}\left[\|{\mathbf{x}}^{(k)}-{\mathbf{x}}^{(k-1)}\|^{2}\right]\leq r_{g}^{(k-1)}+\frac{\alpha_{k-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}},$
+where the second equality holds due to the unbiasedness of ${\mathbf{g}}({\mathbf{x}},\xi_{{\mathbf{g}}})$, the third equality follows from the mutual independence of samples in $\mathcal{B}_{k}$, the second inequality follows by 3(a), and the last one uses
+(31) $\mathbb{E}\left[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|^{2}\right]=\mathbb{E}\left[\left\|\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right)\right)-\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}\right)\right\|^{2}\right]$
+$\leq\textstyle\alpha_{k}^{2}\mathbb{E}\left[\left\|\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\right)\right\|^{2}\right]$
+$\leq 2\alpha_{k}^{2}\left(\mathbb{E}\left[\|\boldsymbol{\zeta}_{f}^{(k)}\|^{2}+\beta^{2}m\sum_{i=1}^{m}\left\|\boldsymbol{\zeta}_{g_{i}}^{(k)}\right\|^{2}\right]\right)$
+$\leq 2\alpha_{k}^{2}\left(\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})+\|\mathbb{E}[\boldsymbol{\zeta}_{f}^{(k)}]\|^{2}+\beta^{2}m\sum_{i=1}^{m}\mathrm{Var}(\boldsymbol{\zeta}_{g_{i}}^{(k)})+\beta^{2}m\sum_{i=1}^{m}\|\mathbb{E}[\boldsymbol{\zeta}_{g_{i}}^{(k)}]\|^{2}\right)$
+$\leq 2\alpha_{k}^{2}(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}m^{2}\sigma_{2}^{2}+\beta^{2}m^{2}l_{g}^{2})=\alpha_{k}^{2}L_{F}^{2}.$
+Since $\mathrm{mod}(k,q)>0$, there exist nonnegative integers $t$ and $s<q$ such that $k=tq+s$. Summing Eq. 30 over $k=tq+1,\ldots,tq+s-1,tq+s$, and applying Eq. 29, we have
 (32) $r_{g}^{(k)}\leq r_{g}^{(tq)}+\sum_{i=1}^{s}\frac{\alpha_{tq+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}.$
-The proof is then completed by combining the above two cases.
+The proof is then completed by combining Case (i) and Case (ii).
 The following corollary is obtained immediately from Eq.
-31 by noticing $r_{g}^{(tq)}=0$ if $\mathbb{P}_{g}$ is a uniform distribution over $N$ data points and ${\mathcal{B}}_{tq}$ contains all the data points.
+32 by noticing $r_{g}^{(tq)}=0$ if $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over $N$ data points and ${\mathcal{B}}_{tq}$ contains all the data points.
 Corollary 3.3.
-Under 1 and 4, suppose that $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
+Under 1 and 3, suppose that $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
 Let ${\mathcal{B}}_{k}=\Xi$ if $\mathrm{mod}(k,q)=0$.
-Then $r_{g}^{(k)}=0$ if $\mathrm{mod}(k,q)=0$ and $r_{g}^{(k)}\leq\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$ otherwise, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
+Then $r_{g}^{(k)}=0$ if $\mathrm{mod}(k,q)=0$ and $r_{g}^{(k)}\leq\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$ otherwise, where $L_{F}$ is given in Lemma 3.1.
 Next, we bound the difference between the Moreau envelope at two consecutive iterates.
+Our analysis follows that of [20, Theorem 3.1], with the key modification in how we handle and bound the term $\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{\mathbf{x}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]$.
 Lemma 3.4.
-Under 1 and 4, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1.
-With $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$, it holds that
+Under 1 and 3, suppose $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1. With $L_{F}$ given in Lemma 3.1, it holds
 (33) $\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k+1)})\right]$
@@
 […]
+$\leq$ $\mathbb{E}_{k}\left[F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})+\rho\left\|\widehat{{\mathbf{x}}}^{(k)}-\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right)\right)\right\|^{2}\right]$
+$=$ $F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})+\rho\left\|{{\mathbf{x}}}^{(k)}-\widehat{{\mathbf{x}}}^{(k)}\right\|^{2}+2\rho\alpha_{k}\mathbb{E}_{k}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right\rangle\right]$
 $\quad\quad\quad\quad+\rho\mathbb{E}_{k}\left[\left\|{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k+1)}\right\|^{2}\right]$
@@
 (34) $\quad\quad\quad\quad+2\rho\alpha_{k}\mathbb{E}_{k}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]+\rho\mathbb{E}_{k}\left[\left\|{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k+1)}\right\|^{2}\right],$
-where the first inequality follows directly from the definition of $F^{\nu}_{\frac{1}{2\rho}}$, and the last one uses $\mathbb{E}_{k}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$ by 4(b) and the $\rho$-weak convexity of $F^{\nu}$ by Lemma 2.6.
+where the first inequality follows directly from the definition of $F^{\nu}_{\frac{1}{2\rho}}$, and the last one uses $\mathbb{E}_{k}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$ by 3(b) and the $\rho$-weak convexity of $F^{\nu}$ by Lemma 2.7.
 Since the function ${\mathbf{x}}\mapsto F^{\nu}({\mathbf{x}})+\rho\left\|{\mathbf{x}}-{{\mathbf{x}}}^{(k)}\right\|^{2}$ is $\rho$-strongly convex, we have
@@
 For the third term in the right hand side (RHS) of Eq. 36, it holds that
-$\textstyle 2\rho\alpha_{k}\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]\leq\mathbb{E}\left[{\alpha_{k}\rho^{2}}\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|^{2}\right]+\mathbb{E}\left[{\alpha_{k}\beta^{2}}{}\|{\mathbf{w}}^{(k)}\|^{2}\right]$
-$=$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+{\alpha_{k}\beta^{2}}{}\mathbb{E}\left[\left\|\left(\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}\right]$
-$\leq$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\left|\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right|^{2}\right]$
-$\leq$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\left|\frac{u^{(k)}}{\nu}-\frac{g({\mathbf{x}}^{(k)})}{\nu}\right|^{2}\right]$
-$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)},$
-where the first equality follows by Eq.
-14 and Eq.
-26, the second inequality comes from 1(c) and 4(b), and the third inequality uses 1-Lipschitz of Proj operator.
-Substituting the above inequality back into Eq.
-35 yields Eq.
-32.
+$2\rho\alpha_{k}\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]\leq\mathbb{E}\left[{\alpha_{k}\rho^{2}}\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|^{2}\right]+\mathbb{E}\left[{\alpha_{k}\beta^{2}}{}\|{\mathbf{w}}^{(k)}\|^{2}\right]$
+$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]$
+$+{\alpha_{k}\beta^{2}}{}\mathbb{E}\left[\left\|\sum_{i=1}^{m}\left(\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\right\|^{2}\right]$
+$\leq$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]$
+$+\alpha_{k}\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\sum_{i=1}^{m}\left|\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\right|^{2}\right]$
+$\leq$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\sum_{i=1}^{m}\left|\frac{u_{i}^{(k)}}{\nu}-\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right|^{2}\right]$
+$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\frac{\alpha_{k}\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)},$
+where the first equality follows from Eq. 14 and Eq. 27, and the second inequality comes from 1(c) and 3(b). Substituting the above inequality back into Eq. 36 yields Eq. 33.
 We now give the convergence result in the following theorem.
 Theorem 3.6.
-Under 1 and 4, given $\overline{\epsilon}>0$, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1 with
-(36) $S_{1}=\lceil 16\overline{\epsilon}^{-2}\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\beta^{2}\sigma^{2}\rceil,\,q=S_{2}=\lceil\sqrt{S_{1}}\rceil,\,$
-(37) $\alpha_{k}=\alpha=\min\left\{\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta\sqrt{l_{g}^{2}+\sigma_{2}^{2}}L_{g}L_{F}}\right\},\text{ and }K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil.$
-Here, $\Delta:=\mathbb{E}[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})]-\min_{{\mathbf{x}}}\mathbb{E}[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}})]$ and $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
-It holds
+Under 1 and 3, given $\overline{\epsilon}>0$, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1 with
+(37) $S_{1}=\lceil 16\overline{\epsilon}^{-2}\nu^{-2}m(l_{g}^{2}+\sigma_{2}^{2})\beta^{2}\sigma^{2}\rceil,\,q=S_{2}=\lceil\sqrt{S_{1}}\rceil,\,$
+(38) $\alpha_{k}=\alpha=\min\Big{\{}\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta\sqrt{m(l_{g}^{2}+\sigma_{2}^{2})}L_{g}L_{F}}\Big{\}},\text{ and }K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil.$
+Here, $\Delta:=F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})-\min_{{\mathbf{x}}}F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}})$, and $L_{F}$ is given in Lemma 3.1. It holds
 (39) $\frac{1}{Kq}\sum_{k=0}^{Kq-1}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq\overline{\epsilon}^{2}.$
 Proof 3.7.
-Since $\alpha_{k}=\alpha$Lemma 3.1 implies
-$\sum_{k=0}^{Kq-1}\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)}\leq\frac{Kq\alpha\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right).$
-Summing Eq. 32 from $k=0$ to $Kq-1$ and applying the above bound, we derive
+[…]
+$\sum_{k=0}^{Kq-1}\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)}\leq\frac{Kq\alpha\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right).$
+Summing Eq. 33 from $k=0$ to Kq1Kq-1 and applying the above bound, we derive
 $\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(Kq)})\right]\leq$ $\,\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})\right]-\frac{\alpha}{{4}}\sum_{k=0}^{Kq-1}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]$
-(39) $+\frac{Kq\alpha\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+Kq{\alpha^{2}{\rho}L_{F}^{2}}{}.$
+(40) $+\frac{Kq\alpha\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+Kq{\alpha^{2}{\rho}L_{F}^{2}}{}.$
 Rearranging the terms and multiplying both sides of Eq. 40 by $\frac{4}{Kq\alpha}$, we obtain
 $\frac{1}{Kq}\sum_{k=0}^{Kq-1}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq$ $\frac{4}{Kq\alpha}\left(\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})\right]-\mathbb{E}\left[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(Kq)})\right]\right)$
-$\quad\quad\quad+\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}$
+$\quad\quad\quad+\frac{4\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}$
 (41) $\leq$ $\frac{4\Delta}{Kq\alpha}+\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}.$
 Now we bound the RHS of Eq.
-40.
+41.
 From $\alpha\leq\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}}$, it follows that $4{\alpha{\rho}L_{F}^{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$.
-The choice $S_{1}=\lceil 16\overline{\epsilon}^{-2}\beta^{2}\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\sigma^{2}\rceil$ yields $\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\sigma^{2}}{S_{1}}\leq\frac{\overline{\epsilon}^{2}}{4}$.
-The conditions $q=S_{2}=\lceil\sqrt{S_{1}}\rceil$ and $\alpha\leq\frac{\overline{\epsilon}\nu}{4\beta\sqrt{l_{g}^{2}+\sigma_{2}^{2}}L_{g}L_{F}}$ imply $\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$.
+The choice $S_{1}=\lceil 16\overline{\epsilon}^{-2}\beta^{2}m\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\sigma^{2}\rceil$ yields $\frac{4\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\sigma^{2}}{S_{1}}\leq\frac{\overline{\epsilon}^{2}}{4}$.
+The conditions $q=S_{2}$ and $\alpha\leq\frac{\overline{\epsilon}\nu}{4\beta\sqrt{m(l_{g}^{2}+\sigma_{2}^{2})}L_{g}L_{F}}$ imply $\frac{4\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$.
 Also, we obtain $\frac{4\Delta}{Kq\alpha}\leq\frac{\overline{\epsilon}^{2}}{4}$ from $K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil$.
@@
 Theorem 3.8.
-Under the conditions of Corollary 3.3, given $\overline{\epsilon}>0$, let $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1 with $S_{1}=N$, (i.e., taking the whole data), and other parameters set to those in Eq.
-36Eq.
-37.
+Under the conditions of Corollary 3.3, given $\overline{\epsilon}>0$, let $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1 with $S_{1}=N$ (i.e., taking the whole data), and other parameters set to those in Eq.
+37Eq.
+38.
 Then Eq.
 39 holds.
-From Theorem 3.6 and Theorem 3.8, we can directly derive the complexity of Algorithm 1 to produce a near-$\overline{\epsilon}$ stationary solution of Eq.
+3.3 Complexity result
+In this subsection, we demonstrate that Algorithm 1 can find an $(\epsilon,\epsilon)$-KKT point for problem Eq.
+P, using the results in Theorem 3.6 and Theorem 2.9.
+From these two theorems, we can directly derive the complexity of Algorithm 1 to produce a near-$\overline{\epsilon}$ stationary solution of Eq.
 PP in expectation by choosing an iterate from $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ uniformly at random.
 Corollary 3.9.
-Under 1 and 4Algorithm 1 can produce a near-$\overline{\epsilon}$ stationary solution of Eq.
-PP in expectation with $Kq=O\left(\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic subgradients and $K(S_{1}+(q-1)S_{2})=O\left(\overline{\epsilon}^{-5}\nu^{-1}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic function evaluations on $g$.
-If $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, the number of stochastic function evaluations becomes $O\left(N+\sqrt{N}\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$.
-3.3 Identifying an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation
-In this subsection, we demonstrate that Algorithm 1 can find an $(\epsilon,\epsilon)$-KKT point for problem Eq.
-P under certain specific scenarios.
-The following proposition addresses the case with a convex constraint or a special weakly-convex case.
+Under 1 and 3Algorithm 1 can produce a near-$\overline{\epsilon}$ stationary solution of Eq.
+PP in expectation with $Kq=O\left(\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic subgradients and $K(S_{1}+(q-1)S_{2})=O\left(\overline{\epsilon}^{-5}\nu^{-1}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic function evaluations on $\mathbf{g}$}.
+If $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, the number of stochastic function evaluations becomes $O\left(N+\sqrt{N}\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$.
+We then present the complexity results in the following proposition.
 Proposition 3.10.
-[…]
+Suppose that 1, 2 with $B_{g}>m\rho_{g}D^{2}$, and 3 hold. Then, given $\epsilon>0$, Algorithm 1 can produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation with
   • $\mathrm{(i)}$
-    $O(\epsilon^{-4})$ OGC/CGC and $O(\epsilon^{-6})$ CFC if $\mathbb{P}_{g}$ is a general distribution;
+    $O(\epsilon^{-4})$ OGC/CGC and $O(\epsilon^{-6})$ CFC if $\mathbb{P}_{{\mathbf{g}}}$ is a general distribution;
   • $\mathrm{(ii)}$
-    $O(\epsilon^{-4})$ OGC/CGC and $O(N+\sqrt{N}\epsilon^{-4})$ CFC if $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
+    $O(\epsilon^{-4})$ OGC/CGC and $O(N+\sqrt{N}\epsilon^{-4})$ CFC if $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
 Proof 3.11.
 Choose $\overline{\epsilon}$, $\beta$, and $\nu$ so that the conditions in Eq.
-17 hold and $\beta>~{}\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-\rho_{g}D^{2}}$.
+17 hold.
 In addition, set $S_{1}$ as in Eq.
-36 if $\mathbb{P}_{g}$ is a general distribution and $S_{1}=N$ , (i.e., taking the whole data), if $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
+37 if $\mathbb{P}_{{\mathbf{g}}}$ is a general distribution and $S_{1}=N$ (i.e., taking the whole data), if $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
 Then set $S_{2},$ $q$, $\alpha_{k}$ and $K$ to those in Eq.
 37 and Eq.
-37.
-The claims now directly follow from Theorem 2.8(i)-(ii) and Corollary 3.9.
+38.
+The claims now directly follow from Theorem 2.9 and Corollary 3.9.
 Remark 3.12.
-If the strong feasibility condition assumed by [9] holds, i.e.,
-(41) $\exists\,{\mathbf{y}}\in{{\mathrm{dom}}(g)}\text{ such that }g({\mathbf{y}})\leq-2\rho_{g}D^{2},$
-we then have 2 with any (big) $B_{g}>0$ when $g$ is convex and 3 with any (big) $B_{g}>0$ when $g$ is weakly convex. Hence, the conditions in the two cases (i)-(ii) of Proposition 3.10 are satisfied.
-Next, we consider a finite-sum weakly convex constraint $g$.
-For this case, we can obtain an ($\epsilon,\epsilon$)-KKT point of problem Eq.
-P with a weaker assumption than Eq.
-41.
-Specifically, we require the (uniform) Slater-type CQ assumption that is assumed in [18, 30].
-The purpose of making this assumption is to ensure $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all $k\geq 0$ so that Theorem 2.8(iii) can be applied.
-Assumption 5 ((uniform) Slater-type CQ).
-There exist $B_{g}>~{}0$, $B>0$, and $\overline{\rho}>\rho_{g}\geq 0$, such that for any ${\mathbf{x}}$ satisfying $g({\mathbf{x}})\leq B_{g}$, it holds $g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for some $\mathbf{y}\in{\mathrm{dom}}(g)$.
-Remark 3.13.
-When $g$ is convex, [18] assumes 1(a) and the existence of ${\mathbf{y}}_{\mathrm{feas}}$ such that $g({\mathbf{y}}_{\mathrm{feas}})<0$.
-These assumptions naturally lead to 5, with the parameters given by ${\mathbf{y}}={\mathbf{y}}_{\mathrm{feas}}$, $\overline{\rho}=\frac{-g({\mathbf{y}}_{\mathrm{feas}})}{D^{2}}$, $B=\frac{-g({\mathbf{y}}_{\mathrm{feas}})}{2}$, and any (big) $B_{g}>0$.
-When $g$ is convex, 5 implies 2 with the same value of $B_{g}$, and when $g$ is $\rho_{g}$-weakly convex with $\rho_{g}\geq 0$, it also implies 3 with the same value of $B_{g}$.
-We will verify these two assertions in Appendix A.
-Moreover, 5 is clearly implied by the strong feasibility condition Eq.
-41.
-Before giving our complexity result in this case, we establish a lower bound relating $g_{+}({\mathbf{x}})$ to the distance from ${\mathbf{x}}$ to the feasible set $\Omega$ in the following lemma.
-Its proof is given in Lemma A.5.
-Lemma 3.14.
-Suppose that $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, and 1 and 5 hold.
-If $\textnormal{dist}({\mathbf{x}},\Omega)\leq\min\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\}$, then $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\gamma=\sqrt{B(\overline{\rho}-\rho_{g})/2}$.
-In the next lemma, we establish that Algorithm 1 guarantees $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all iterations.
-Lemma 3.15.
-Suppose that $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, and 14, and 5 hold.
-In Algorithm 1, let $u^{(k)}=g({\mathbf{x}}^{(k)})$, i.e., ${\mathcal{B}}_{k}=\Xi$ for all $k\geq 0$ in Eq.
-25, and choose $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$.
-Suppose $\|\boldsymbol{\zeta}_{f}^{(k)}\|\leq l_{f}$ for all $k\geq 0$.
-Set
-[…]
-(43) $0<\alpha_{k}\leq\min\left\{\frac{C_{0}}{\beta l_{g}^{2}},{\frac{1}{\beta\rho_{g}},}\frac{\gamma}{2l_{f}\rho_{g}},\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}},\,\frac{B_{g}}{4l_{f}l_{g}}\right\},$
-where $\gamma=\sqrt{B(\overline{\rho}-\rho_{g})/2}$.
-If the initial point ${\mathbf{x}}^{(0)}$ satisfies $\textnormal{dist}({\mathbf{x}}^{(0)},\Omega)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$, then it holds $\textnormal{dist}({\mathbf{x}}^{(k)},\Omega)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ and $g({\mathbf{x}}^{(k)})\leq\frac{B_{g}}{2}$ for all $k\geq 0$.
-Proof 3.16.
-We prove the desired results by induction on $k$.
-The claim holds for $k=0$, by the assumption on ${\mathbf{x}}^{(0)}$.
-Suppose it holds up to iteration $k$.
-We show that the same bounds hold for $k+1$.
-We consider the two cases: (i) $g({\mathbf{x}}^{(k)})\geq~{}C_{0}$; (ii) $g({\mathbf{x}}^{(k)})<~{}C_{0}$.
-(i) In this case, since $\nu\leq C_{0}\leq g({\mathbf{x}}^{(k)})$, the updating rule from Eq. 1 becomes
-${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\boldsymbol{\zeta}_{g}^{(k)}\right)=\left({\mathbf{x}}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}\right)-\alpha_{k}\boldsymbol{\zeta}_{f}^{(k)}.$
-Since $0<g({\mathbf{x}}^{(k)})\leq l_{g}\textnormal{dist}\left(\mathbf{x}^{(k)},\Omega\right)<B_{g}$, it follows from Lemma A.1 that $\boldsymbol{\zeta}_{g}^{(k)}\neq\mathbf{0}$.
-For simplicity of notation, let ${\mathbf{y}}^{(k)}={\mathbf{x}}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}$ and $\mathbf{x}^{\dagger(k)}=\operatorname{Proj}_{\Omega}(\mathbf{x}^{(k)})$.
-It holds
-$\small{\textnormal{dist}^{2}({{\mathbf{y}}}^{(k)},\Omega)\leq\left\|{{\mathbf{y}}}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}=\left\|\mathbf{x}^{(k)}-\alpha_{k}\beta\boldsymbol{\zeta}_{g}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}}$
-$=\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)-2\alpha_{k}\beta\left\langle\boldsymbol{\zeta}_{g}^{(k)},\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\rangle+(\alpha_{k}\beta)^{2}\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}}$
-$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+2\alpha_{k}\beta\left(g(\mathbf{x}^{\dagger(k)})-g(\mathbf{x}^{(k)})+\frac{\rho_{g}}{2}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)+(\alpha_{k}\beta)^{2}\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}}$
-$=\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-g_{+}(\mathbf{x}^{(k)})+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)+\alpha_{k}\beta\left(\alpha_{k}\beta\left\|\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}-g(\mathbf{x}^{(k)})\right)}$
-$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-g_{+}(\mathbf{x}^{(k)})+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2}\right)}$
-$\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}\beta\left(-{\gamma}{}+\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|\right)\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|}$
-(44) $\leq\small{\textnormal{dist}^{2}(\mathbf{x}^{(k)},\Omega)-\alpha_{k}\beta\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|^{2},}$
-where the third equality comes from $g(\mathbf{x}^{\dagger(k)})=0$ and $g(\mathbf{x}^{(k)})=g_{+}(\mathbf{x}^{(k)})$, the second inequality holds by the $\rho_{g}$-weak convexity of $g$, the third one uses $g(\mathbf{x}^{(k)})\geq C_{0}$, $\|\boldsymbol{\zeta}_{g}^{(k)}\|\leq l_{g}$, and $0<\alpha_{k}\beta\leq\frac{C_{0}}{l_{g}^{2}}$, the fourth one follows by Lemma 3.14, and the last one holds because $-{\gamma}{}\leq-2\rho_{g}\textnormal{dist}({\mathbf{x}}^{(k)},\Omega)=-2\rho_{g}\left\|\mathbf{x}^{(k)}-\mathbf{x}^{\dagger(k)}\right\|$ by induction.
-Then it follows that
-$\vspace{-1mm}\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)$ $\leq\textnormal{dist}({{\mathbf{y}}}^{(k)},\Omega)+\alpha_{k}l_{f}\leq\sqrt{1-\alpha_{k}\beta\rho_{g}}\textnormal{dist}(\mathbf{x}^{(k)},\Omega)+\alpha_{k}l_{f}$
-$\leq\textnormal{dist}(\mathbf{x}^{(k)},\Omega),$
-where the second inequality uses Eq.
-44, and the last one follows from $\beta\geq\frac{2l_{f}l_{g}}{\rho_{g}C_{0}}$ and $\textnormal{dist}\left(\mathbf{x}^{(k)},\Omega\right)\geq\frac{g({\mathbf{x}}^{(k)})}{l_{g}}\geq~{}\frac{C_{0}}{l_{g}}$.
-Thus $\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ by induction, and $g({{\mathbf{x}}}^{(k+1)})\leq l_{g}\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\frac{B_{g}}{2}$.
-(ii) In the second case, we have $g({\mathbf{x}}^{(k)})\leq C_{0}.$ From Eq. 1, we have
-$\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)\leq$ $\textnormal{dist}({{\mathbf{x}}}^{(k)},\Omega)+\alpha_{k}(l_{f}+\beta l_{g})$
-(45) $\leq$ $\min\left\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\right\}+\alpha_{k}l_{f}+\frac{C_{0}}{l_{g}}\leq\min\left\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\right\},$
-where the second inequality follows from $\alpha_{k}\leq\frac{C_{0}}{\beta l_{g}^{2}}$ and $\textnormal{dist}\left({{\mathbf{x}}}^{(k)},\Omega\right)\leq\min\{\frac{\gamma}{2\rho_{g}},\frac{B_{g}}{2l_{g}}\}$ by induction hypothesis, and the last one is derived using $\alpha_{k}\leq\min\{\frac{\gamma}{2l_{f}\rho_{g}},\frac{B_{g}}{4l_{f}l_{g}}\}$, and $C_{0}\leq\min\{\frac{\gamma l_{g}}{\rho_{g}},\frac{B_{g}}{4}\}$.
-We then obtain
-$\textnormal{dist}({{\mathbf{x}}}^{(k+1)},\Omega)\leq\frac{g({\mathbf{x}}^{(k+1)})}{\gamma}\leq\frac{1}{\gamma}\left(g({\mathbf{x}}^{(k)})+l_{g}\|{{\mathbf{x}}}^{(k+1)}-{{\mathbf{x}}}^{(k)}\|\right)$
-(46) $\leq$ $\frac{1}{\gamma}\left(g({\mathbf{x}}^{(k)})+\alpha_{k}l_{g}(l_{f}+\beta l_{g})\right)\leq\frac{C_{0}}{\gamma}+\frac{1}{\gamma}\left(\alpha_{k}l_{g}l_{f}+{C_{0}}{}\right)\leq\min\left\{\frac{\gamma}{2\rho_{g}},\,\frac{B_{g}}{2l_{g}}\right\},$
-where the first inequality uses Lemma 3.14 by Eq.
-45, the second/third one comes from the updating rule Eq.
-1 and 1(c), the fourth one holds because $g({\mathbf{x}}^{(k)})\leq C_{0}$ and $\alpha_{k}\leq\frac{C_{0}}{\beta l_{g}^{2}}$, and the last one follows by $C_{0}\,\leq\,\min\{\frac{\gamma^{2}}{8\rho_{g}},\,\frac{\gamma B_{g}}{8l_{g}}\},$ and $\alpha_{k}\,\leq\,\min\{\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}}\}.$ Hence $g({{\mathbf{x}}}^{(k+1)})\leq l_{g}\textnormal{dist}\left({{\mathbf{x}}}^{(k+1)},\Omega\right)\leq\frac{B_{g}}{2}$.
-Finally, combining the two above cases and applying the induction hypothesis, we complete the proof.
-Now we are ready to show the complexity result for the finite-sum weakly convex case under 5.
-Proposition 3.17.
-Suppose $g$ is $\rho_{g}$-weakly convex with $\rho_{g}>0$, $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, and 14, and 5 hold.
-In Algorithm 1, let $u^{(k)}=g({\mathbf{x}}^{(k)})$, i.e., ${\mathcal{B}}_{k}=\Xi$ for all $k\geq 0$ in Eq.
-25, and choose $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$.
-Suppose $\|\boldsymbol{\zeta}_{f}^{(k)}\|\leq l_{f}$ for all $k\geq 0$.
-Then given $\epsilon>0$Algorithm 1 can produce an $(\epsilon,\epsilon)$-KKT point of problem Eq.
-P in expectation with $O(\epsilon^{-4})$ stochastic subgradients of $f$, and $O(\epsilon^{-4})$ deterministic subgradients and function evaluations of $g$.
-Proof 3.18.
-Choose $\overline{\epsilon}$ and $\beta$ to satisfy the conditions in Eq.
-17 and Eq.
-42 and $\beta>\frac{2B_{f}}{B_{g}}$, and let $\nu$ satisfy $0<\nu\leq\min\{C_{0},{B}_{g},\frac{\epsilon}{\beta},\frac{\epsilon}{2}\},$ where $C_{0}$ is given in Lemma 3.15.
-Also, choose the stepsize as
-$\alpha_{k}=\alpha=\min\left\{\frac{C_{0}}{\beta l_{g}^{2}},{\frac{1}{\beta\rho_{g}},}\frac{\gamma}{2l_{f}\rho_{g}},\frac{\gamma^{2}}{4l_{f}l_{g}\rho_{g}},\,\frac{\gamma B_{g}}{4l_{f}l^{2}_{g}},\,\frac{B_{g}}{4l_{f}l_{g}},\,\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta l_{g}L_{g}L_{F}}\right\}.\vspace{-2mm}$
-Since $\boldsymbol{\zeta}_{g}^{(k)}\in\partial g({\mathbf{x}}^{(k)})$ and $u^{(k)}=g({\mathbf{x}}^{(k)})$, it holds $\sigma=\sigma_{2}=0$.
-Hence, we can still obtain Eq.
-38 from Eq.
-40 by setting $q=1$, $S_{2}=N$, (i.e., taking the whole data), and $K=\lceil 16\overline{\epsilon}^{-2}\alpha^{-1}\Delta\rceil$, thus a near-$\overline{\epsilon}$ stationary point ${\mathbf{x}}^{*}$ of Eq.
-PP in expectation can be produced within $K$ iterations.
-Now by Theorem 2.8(iii) and Lemma 3.15, ${\mathbf{x}}^{*}$ is also an $(\epsilon,\epsilon)$-KKT point of problem Eq.
-P in expectation.
-It takes $K=O(\epsilon^{-4})$ stochastic subgradient of $f$ and $K=O(\epsilon^{-4})$ deterministic subgradients and function evaluations of $g$.
-The proof is then completed.
+When $\mathbf{g}$} is deterministic, our algorithm still converges to an $(\epsilon,\epsilon)$-KKT point of problem Eq.
+P in expectation without requiring the condition $B_{g}>m\rho_{g}D^{2}$.
+This condition appears only once in our analysis, specifically in the proof of inequality (22), to ensure well-definedness of parameters $\beta$ and $\nu$ in (17).
+Under this assumption, inequality (22) and the subsequent arguments guarantee that $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$, which then enables us to use the Slater-type CQ assumption.
+In contrast, for the deterministic case, the constraints $g_{i}({\mathbf{x}})\leq 0,\forall\,i\in[m]$ can be equivalently expressed as $\max_{i\in[m]}g_{i}({\mathbf{x}})\leq 0$.
+Leveraging this reformulation, we directly utilize the established convergence result for the special case of $m=1$, detailed in our arXiv first version [44].
+This prior result ensures that, provided the initial point is sufficiently close to the feasible set $\mathcal{X}$}, all iterates produced by our method satisfy $\max_{i\in[m]}g_{i}(\widehat{{\mathbf{x}}}^{(k)})\leq B_{g}$.
+Consequently, the assumption $B_{g}>m\rho_{g}D^{2}$ becomes unnecessary for this case, provided that we can obtain a (near) feasible initial point.
 4 Numerical experiments
-In this section, we evaluate the empirical performance of the proposed Algorithm 1 on two fairness-constrained problems, which are instances of Eq.
-P.
-We compare its performance with three state-of-the-art approaches, including two different double-loop IPP [9, 30] and the SSG in [18].
+In this section, we evaluate the empirical performance of the proposed Algorithm 1 and compare to three state-of-the-art approaches, including two different IPPs [10, 46] and the SSG in [27].
 All the experiments are performed on iMac with Apple M1, a 3.2 GHz 8-core processor, and 16GB of RAM running MATLAB R2024a.
-4.1 Datasets and tested problems
-The experiments are conducted on two datasets: a9a [20], and COMPAS [5].
-The a9a dataset contains 48,842 data points with 123 features, and is used to predict income levels; gender is the protected attribute, and fairness is assessed between males and females.
-The COMPAS dataset includes 6,172 data points and 16 features, focused on predicting recidivism risk, with fairness evaluated between Caucasian and non-Caucasian groups.
-We split each dataset into two subsets with a ratio of 2:1.
-The larger set is used as $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, and the smaller set is further split into a protected set $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and an unprotected set $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$.
+4.1 Tested problems and Datasets
+We conduct experiments on solving two fairness-constrained problems and one Neyman-Pearson classification problem, which are instances of Eq. P.
 Problem 1.
@@
 It is formulated in Eq.
-4.
-For this problem, we have $\rho_{g}=0$ and $\rho_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}$.
-That is, $\Phi(\cdot)-\Phi^{*}-\kappa_{1}$ is convex.
-We obtain the minimum $\Phi^{*}$ and a solution ${\mathbf{x}}^{*}$ by applying a stochastic subgradient method to a large number of iterations on $\min_{{\mathbf{x}}}\Phi({\mathbf{x}})$.
-Afterward, we set $\kappa_{1}=0.001\cdot\Phi^{*}$, $D=5\|{\mathbf{x}}^{*}\|$, and let $\Theta$ consist of 400 points equally spaced between $\uline{{\mathbf{a}}}-0.5\left(\overline{{\mathbf{a}}}-\uline{{\mathbf{a}}}\right)$ and $\overline{{\mathbf{a}}}+0.5\left(\overline{{\mathbf{a}}}-\uline{{\mathbf{a}}}\right)$, where $\overline{{\mathbf{a}}}=\max_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$ and $\uline{{\mathbf{a}}}=\min_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$.
+5.
+We obtain the minimum $\Phi^{*}$ and a solution $\mathbf{x}$$^{*}$ by applying a stochastic subgradient method to a large number of iterations on $\min_{{\mathbf{x}}\in{\mathcal{X}}}\Phi({\mathbf{x}})$.
+Afterward, we set $\kappa_{1}=0.001\cdot\Phi^{*}$, $D=5\|{\mathbf{x}}^{*}\|$, and let $\Theta$ consist of 400 points equally spaced between $\uline{a}-0.5\left(\overline{a}-\uline{a}\right)$ and $\overline{a}+0.5\left(\overline{a}-\uline{a}\right)$, where $\overline{a}=\max_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$ and $\uline{a}=\min_{i}({\mathbf{x}}^{*})^{\top}\mathbf{a}_{i}$.
 For all the compared methods, we initialize ${\mathbf{x}}^{(0)}={\mathbf{x}}^{*}$.
+For this problem, we have $l_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|$, $l_{g}=\frac{1}{n}\sum_{i=1}^{n}\|{\mathbf{a}}_{i}\|$, $\rho_{f}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}$, and $\rho_{g}=0$ from [27].
+That is, $\Phi(\cdot)-\Phi^{*}-\kappa_{1}$ is convex, and 1 is satisfied.
+Moreover, 2 holds with parameters $B_{g}=1$, $\overline{\rho}=\frac{-\max_{i\in[m]}g_{i}({\mathbf{x}}^{*})}{D^{2}}$, and $B=\frac{-\max_{i\in[m]}g_{i}({\mathbf{x}}^{*})}{2}$.
 Problem 2. The second problem is a classification problem with demographic parity [1] and the smoothly clipped absolute deviation (SCAD) regularization term [23]. It is formulated as
-(47) $\min_{{\mathbf{x}}}\Phi({\mathbf{x}})+\lambda\|\varphi({\mathbf{x}})\|_{1},\,\,\text{{s.t.}}\,\,\Psi_{0}({\mathbf{x}})\leq\kappa_{2},$
+(42) $\min_{{\mathbf{x}}:\|{\mathbf{x}}\|_{\infty}\leq D}\Phi({\mathbf{x}})+\lambda\|\varphi({\mathbf{x}})\|_{1},\,\,\text{{s.t.}}\,\,\Psi_{0}({\mathbf{x}})\leq\kappa_{2},$
 where $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, $\Psi_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}$, $\Phi$ is the same as that in Eq.
-4, $\lambda>0$, $\kappa_{2}>0$ is a slackness parameter.
+5, $D>0$, $\lambda>0$, $\kappa_{2}>0$ is a slackness parameter.
 The functions $\Psi_{0}$ and $\varphi$ are defined as follows:
 $\small{\Psi_{0}({\mathbf{x}}):=\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\sigma\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\sigma\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\right)\right|,\text{ and }}$
-$\small{[\varphi({\mathbf{x}})]_{i}=\begin{cases}2\left|x_{i}\right|,&0\leq\left|x_{i}\right|\leq 1,\\ -x_{i}^{2}+4\left|x_{i}\right|+1,&1<\left|x_{i}\right|\leq 2,\\ 3,&\left|x_{i}\right|>2,\end{cases}\quad\text{ for }i=1,2,\ldots,d}$
+$\small{[\varphi({\mathbf{x}})]_{i}=\begin{cases}2\left|x_{i}\right|,&0\leq\left|x_{i}\right|\leq 1,\\ -x_{i}^{2}+4\left|x_{i}\right|+1,&1<\left|x_{i}\right|\leq 2,\\ 3,&\left|x_{i}\right|>2,\end{cases}\quad\text{ for }i=1,2,\ldots,d},$
 where $\varphi$ promotes sparsity in $\mathbf{x}$}.
 Here, $\Psi_{0}$ is a continuous relaxation of the demographic parity $\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\geq 0\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\geq 0\right)\right|$.
-For this problem, we have $\rho_{f}=\rho_{g}=\max\left\{2\lambda,\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}\right\}.$ In the experiments, we set $\lambda=0.02$, $\kappa_{2}=0.02$, and ${\mathbf{x}}^{(0)}=\mathbf{0}$ for all the compared methods.
+For this problem, we discuss in Appendix B that its constraint satisfies 2, and from [27], we obtain $\rho_{f}=\rho_{g}=\max\big{\{}2\lambda,\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}\big{\}}$, and the Lipschitz modulus $l_{g}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|$, $l_{f}=5\lambda+\frac{1}{n}\sum_{i=1}^{n}\|{\mathbf{a}}_{i}\|$.
+In the experiments, we set $D=5$, $\lambda=0.02$, $\kappa_{2}=0.02$, and ${\mathbf{x}}^{(0)}=\mathbf{0}$ for all the compared methods.
+Problem 3.
+The third tested problem is a multi-class Neyman-Pearson classification problem formulated in Eq.
+3.
+We set $D=0.3$, $\kappa=4.5$, and let $\phi$ be the sigmoid function $\sigma$.
+Similar to those constants of Problem 2, we have $l_{f}=l_{g}=\max_{i\in[M]}\frac{1}{|{\mathcal{D}}_{i}|}\sum_{{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}}\|{\mathbf{a}}_{i}\|$, and $\rho_{f}=\rho_{g}=\max_{i\in[M]}\frac{1}{|{\mathcal{D}}_{i}|}\sum_{{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}}\|{\mathbf{a}}_{i}\|^{2}$.
+As discussed in Appendix B, the constraint of the tested problem also satisfies 2.
+Datasets.
+The experiments of Problems 1–2 are conducted on two datasets: a9a [30] and COMPAS [5].
+The a9a dataset contains 48,842 data points with 123 features and is used to predict income levels; gender is the protected attribute, and fairness is assessed between males and females.
+The COMPAS dataset includes 6,172 data points and 16 features, focused on predicting recidivism risk, with fairness evaluated between Caucasian and non-Caucasian groups.
+We split each dataset into two subsets with a ratio of 2:1.
+The larger set is used as $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, and the smaller subgroup is further partitioned into two subsets based on a binary protected attribute: $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$.
+The experiments for Problem 3 are conducted using the MNIST dataset [34], which contains 60,000 grayscale images of handwritten digits (0–9), each image having $\times28$ pixels and an associated ground-truth label.
+We set $M=10$ and define each subset ${\mathcal{D}}_{i}$ as the collection of images corresponding to the image labeled by digit $\mathrm{mod}(i,10)$, for $i\in[M]$.
 4.2 Implementation details
@@
 For 3S-Econ-D, we set $q=1$ and $S_{1}=S_{2}=n$ (i.e., taking the whole data) and select $\alpha_{k}=10^{-2}$ as our default setting.
-For 3S-Econ-S, we set $S_{1}=n$ and $S_{2}=q=\lceil\sqrt{n}\rceil$ and select $\alpha_{k}=\frac{1}{100\sqrt{k/q}}$ as our default setting [^2] To determine the default settings, we test 3S-Econ-D and 3S-Econ-S on Problem 1 using the a9a dataset, with varying choices of $\beta$ and $\alpha_{k}$.
-Specifically, for all $k\geq 0$, we choose $\beta\in\{1,10,100,1000\}$ and $\alpha_{k}\in\{0.001,0.01,\frac{1}{100\sqrt{k/q}},\frac{1}{1000\sqrt{k/q}}\}$.
-Through our experiments, we observe that 3S-Econ-D achieves the best performance with $\beta=10$ and $\alpha_{k}=0.01$.
-Similarly, for 3S-Econ-S, $\beta=10$ and $\alpha_{k}=\frac{1}{100\sqrt{k/q}}$ yield the best performance.
+For 3S-Econ-S, we set $S_{1}=n$ and $S_{2}=q=\lceil\sqrt{n}\rceil$ and select $\alpha_{k}=\frac{1}{100\lceil\sqrt{k/q}\rceil}$ as our default setting [^3] To determine the default settings of the stepsize, we test 3S-Econ-D and 3S-Econ-S on Problem 1 using the a9a dataset, with varying choices of $\alpha_{k}$.
+Specifically, for all $k\geq 0$, we choose $\alpha_{k}$ from $\{0.001,0.01,\frac{1}{100\lceil\sqrt{k/q}\rceil},\frac{1}{1000\lceil\sqrt{k/q}\rceil}\}$.
+Through our experiments, we observe that 3S-Econ-D achieves the best performance with $\alpha_{k}=0.01,\forall\,k$.
+Similarly, for 3S-Econ-S, $\alpha_{k}=\frac{1}{100\lceil\sqrt{k/q}\rceil},\forall\,k,$ yields the best performance.
 Based on these observations, we fix these parameters as the default settings and apply them across all datasets and tested problems.
 .
-We adopt the tuning techniques in [9, 30] for the two distinct double-loop IPP methods and the techniques in [18] for the SSG.
+We adopt the tuning techniques in [10, 46] for the two distinct IPP methods and the techniques in [27] for the SSG.
 A strongly convex constrained subproblem is solved approximately in each outer iteration of the IPPs.
@@
 The optimal parameter set for these methods is selected based on the smallest objective value after 5000 iterations.
-We report the objective function value (FV), i.e., $f({\mathbf{x}})$, the constraint violation (CVio), i.e., $[g({\mathbf{x}})]_{+}$, and the stationarity violation (SVio). At each iteration, we solve
-$\min_{{\mathbf{x}}}\big{\{}f({\mathbf{x}})+\rho_{f}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2},\,\text{{s.t.}}\,g({\mathbf{x}})+\rho_{g}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2}\leq 0\big{\}}$
-to a desired accuracy to obtain $\widehat{{\mathbf{x}}}^{(k)}$ and then use $\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|$ to measure SVio [^3] Specically, we solve the strongly convex problem by the SSG [18] to a tolerance of $10^{-2}$..
+We report the objective function value (FV), i.e., $f({\mathbf{x}})$, the constraint violation (CVio), i.e., $\|[{\mathbf{g}}({\mathbf{x}})]_{+}\|_{1}$, and the stationarity violation (SVio).
+At each iteration, we solve
+$\min_{{\mathbf{x}}}\big{\{}f({\mathbf{x}})+\rho_{f}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2},\,\text{{s.t.}}\,g_{i}({\mathbf{x}})+\rho_{g}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2}\leq 0,\,i\in[m]\big{\}}$
+to a desired accuracy to obtain $\widehat{{\mathbf{x}}}^{(k)}$ and then use $\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|$ to measure SVio [^4] Specifically, we solve the strongly convex problem by the SSG [27] to a tolerance of $10^{-2}$..
+4.3 Performance of 3S-Econ-D with different $\beta$ and $\nu$
+In this subsection, we investigate the numerical performance of 3S-Econ-D with different values of $\beta$ and $\nu$ on solving Problem 1 by using the a9a dataset, respectively.
+We vary the penalty parameter $\beta\in\{1,10,100,1000\}$ with fixed $\nu=10^{-5}$, and subsequently vary the smoothing parameter $\nu\in\{10^{-5},10^{-4},10^{-3},10^{-2}\}$ with fixed $\beta=10$.
+The results are depicted in Fig.
+1.
+From the figure, we observe that when $\beta=1$, the CVio remains relatively large, indicating inadequate penalization, while excessively large $\beta$ slows down convergence.
+Among the varying values, $\beta=10$ yields the most balanced trade-off between convergence speed and constraint satisfaction.
+In addition, we observe that the CVio closely aligns with $\nu$, consistent with our theoretical results.
+On all subsequent experiments, we select $\beta=10$ and $\nu=10^{-5}$ as our default settings.
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 1: Performance of 3S-Econ-D with different $\beta$ and $\nu$ on solving Problem 1 by using the a9a dataset.
+Top row: different values of $\beta$ and fixed $\nu=10^{-5}$; Bottom row: different values of $\nu$ and fixed $\beta=10$.
 4.4 Comparisons among different methods
 In Fig.
-1, we compare the 3S-Econ-D method with three state-of-the-art deterministic approaches on solving Problems 1–2.
+2, we compare our 3S-Econ-D method with three state-of-the-art deterministic approaches on solving Problems 1–2 over iterations [^5] Here, for the two IPP methods, the iteration count refers specifically to the inner-loop iterations..
 It is clear that 3S-Econ-D outperforms the others in FVs and SVio values, not only converging faster but also achieving lower errors.
@@
 Since for the deterministic version, we can explicitly check the feasibility.
-Thus by keeping feasible iterates, 3S-Econ-D will significantly outperform the other three deterministic methods.
+By keeping feasible iterates, 3S-Econ-D will significantly outperform the other three deterministic methods.
 (a) Problem 1, a9a
@@
 The FV and CVio values on Problems 1-2 with two datasets are shown in Fig.
-2, “data passes” for the $x$-axis label is a shorthand for “data passes ($g$)” (DP ($g$)), referring to the number of times to access all the data involved in the constraint function.
-Despite more oscillations due to the stochasticity, 3S-Econ-S can produce comparable results as 3S-Econ-D with significantly fewer data passes.
+3, “data passes” for the $x$-axis label is a shorthand for “data passes ($\mathbf{g}$})” (DP ($\mathbf{g}$})), referring to the number of times to access all the data involved in the constraint function.
+Despite more oscillations due to the stochasticity, 3S-Econ-S can produce comparable results to 3S-Econ-D with significantly fewer data passes.
 (a) Prob. 1, a9a
@@
 We terminate all the deterministic methods if SVio is less than $10^{-3}$, while for the stochastic algorithms, the threshold is set to $5\times 10^{-3}$.
-In addition, we impose a maximum “DP ($g$)” to 200,000 for Problem 1 and 720,000 for Problem 2.
+In addition, we impose a maximum “DP ($\mathbf{g}$})” of 200,000 for Problem 1 and 720,000 for Problem 2.
 Since computing $\widehat{{\mathbf{x}}}^{(k)}$ with high precision at each iteration $k$ is time-consuming, we calculate SVio every few steps.
-Upon termination, in terms of “DP ($g$)” that dominate “DP ($f$)” for all methods, we observe that for Problem 1, 3S-Econ-D is 5 times faster than other deterministic methods on the a9a dataset and 2 times faster on the COMPAS dataset; for Problem 2, 3S-Econ-D is 2-15 times faster than other deterministic methods on the a9a dataset, and 1.5-2 times faster on the COMPAS dataset.
-Furthermore, 3S-Econ-S delivers even more impressive speedups: for Problem 1, it is 219 times faster than SSG-S on the a9a dataset and 108 times faster on the COMPAS dataset; for Problem 2, it is 466 times faster than SSG-S on the a9a dataset and 26 times faster on the COMPAS dataset.
-Additionally, in most cases, our methods demonstrate superior performance in terms of both FV and SVio when compared to other algorithms.
+Upon termination, our proposed methods demonstrate superior performance in terms of both FV and SVio compared to other tested algorithms in most cases.
+In terms of “DP ($\mathbf{g}$})”, which consistently dominates “DP (ff)” for all methods, we observe that for Problem 1, 3S-Econ-D achieves speedups ranging from 1.5 to 5 times compared to other deterministic algorithms, while 3S-Econ-S delivers even more impressive speedups, approximately 108 to 466 times faster.
+When considering the metric “(CPU) Time (s)”, both 3S-Econ-D and 3S-Econ-S remain significantly faster than their counterparts.
+Although the acceleration factor of 3S-Econ-S in terms of runtime is less pronounced compared to that measured by “DP ($\mathbf{g}$})”, this discrepancy primarily arises from implementation-related overhead inherent in MATLAB, such as matrix-vector multiplications.
 measures SSG IPP-SSG IPP-ConEx 3S-Econ-D SSG-S 3S-Econ-S
@@
 DP ($f$) 1.00e+05 1.00e+05 1.00e+05 1.50e+04 1.10e+03 455
-DP ($g$) 2.00e+05 2.00e+05 2.00e+05 3.00e+04 2.00e+05 910
+DP ($\mathbf{g}$}) 2.00e+05 2.00e+05 2.00e+05 3.00e+04 2.00e+05 910
+Time (s) 1.38e+03 1.37e+03 3.70e+03 582 1.03e+03 537
 FV 9.26e-02 1.03e-01 9.70e-02 8.47e-02 8.08e-02 8.48e-02
@@
 DP ($f$) 1.00e+05 1.00e+05 1.00e+05 3.70e+04 3.06e+03 938
-DP ($g$) 2.00e+05 2.00e+05 2.00e+05 7.40e+04 2.00e+05 1.85e+03
+DP ($\mathbf{g}$}) 2.00e+05 2.00e+05 2.00e+05 7.40e+04 2.00e+05 1.85e+03
+Time (s) 285 178 236 127 273 32.9
 FV 6.91e-03 6.79e-03 6.85e-03 6.31e-03 6.49e-03 6.41e-03
@@
 DP ($f$) 4.02e+04 3.60e+05 3.60e+05 2.10e+04 282 55
-DP ($g$) 8.04e+04 7.20e+05 7.20e+05 4.20e+04 5.13e+04 110
+DP ($\mathbf{g}$}) 8.04e+04 7.20e+05 7.20e+05 4.20e+04 5.13e+04 110
+Time (s) 305 2.35e+03 7.76e+03 217 134 35.6
 FV 5.11e-01 5.07e-01 5.14e-01 5.05e-01 5.07e-01 5.10e-01
@@
 DP ($f$) 2.89e+05 3.60e+05 3.10e+05 1.54e+05 1.80e+03 2.17e+03
-DP ($g$) 5.78e+05 7.20e+05 6.20e+05 3.07e+05 1.17e+05 4.35e+03
+DP ($\mathbf{g}$}) 5.78e+05 7.20e+05 6.20e+05 3.07e+05 1.17e+05 4.35e+03
+Time (s) 60.0 57.5 156 40.9 27.3 23.8
 FV 2.04e-01 2.04e-01 2.04e-01 2.04e-01 2.06e-01 2.06e-01
@@
 SVio 8.54e-04 8.30e-04 7.99e-04 9.43e-04 4.06e-03 4.70e-03
-Table 2: Comparisons among all the methods.
-5 Conclusions
-We have presented a novel single-loop first-order method for solving nonconvex nonsmooth stochastic optimization with expectation constraints.
+Table 2: Comparisons among different methods on solving Problems 1 and 2.
+In Fig.
+4, we solve Problem 3 on the MNIST dataset and present the values of FV, CVio, and SVio versus CPU time for SSG, IPP-ConEx, IPP-SSG, and our proposed methods.
+Overall, our methods reach lower FV and SVio values more quickly compared to other methods, while achieving feasibility at most iterates.
+Specifically, 3S-Econ-S attains an SVio of $1.29\times 10^{-3}$ at 21.9 seconds, and 3S-Econ-D reaches an SVio of $1.53\times 10^{-3}$ at 172 seconds, significantly outperforming SSG, which achieves its lowest SVio of $1.07\times 10^{-2}$ at about 225 seconds.
+Additionally, IPP-ConEx remains stagnant at the zero solution, likely due to the zero point being a stationary solution of its proximal subproblem.
+Figure 4: Comparisons of different methods on Problem 3 by using MNIST dataset.
+5 Conclusions and Discussions
+We have presented a novel FOM for solving nonconvex nonsmooth stochastic optimization with expectation constraints.
 Built on an exact penalty formulation that applies a smoothed penalty function on the inequality expectation constraint, our method incorporates the SPIDER-type estimation of the constraint function into the stochastic subgradient method.
-Through establishing the equivalence between the near-stationary solution of the penalty problem and the near-KKT solution of the original problem, we achieve an optimal $O(\epsilon^{-4})$ iteration complexity result to produce an $(\epsilon,\epsilon)$-KKT point.
+Through establishing the equivalence between the near-stationary solution of the penalty problem and the near KKT solution of the original problem, we achieve an optimal $O(\epsilon^{-4})$ iteration complexity result to produce an $(\epsilon,\epsilon)$-KKT point.
 In terms of sample complexity, our result is lower than a few state-of-the-art results by a factor of $\epsilon^{-2}$ in either constraint sample subgradient or constraint function values, or by a factor of $\sqrt{N}$ when the constraint is the average of $N$ sample functions.
-Moreover, our method delivers significantly superior numerical performance over state-of-the-art methods on solving fairness-constrained problems.
+Moreover, our method delivers significantly superior numerical performance over state-of-the-art methods on solving fairness-constrained problems and Neyman-Pearson classification problem.
+[…]
 Acknowledgement
-We would like to express our sincere gratitude to Dr. Yankun Huang and Prof. Qihang Lin for providing their code in [18].
-Appendix A Comparisons among our CQs and others
-As mentioned in Remark 3.13, the strong feasibility condition is stronger than the (uniform) Slater-type CQ stated in 5.
-In the following lemmas, we demonstrate that 2 and 3 can be implied by the uniform Slate-type CQ under mild conditions.
-Lemma A.1.
-Under 1 and 5, if $0\leq g({\mathbf{x}})\leq B_{g}$, then $\min_{\boldsymbol{{\boldsymbol{\zeta}}_{g}}\in\partial g(\mathbf{x})}\left\|\boldsymbol{{\boldsymbol{\zeta}}_{g}}\right\|\geq\theta,$ where $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$.
-Proof A.2.
-Becasue $g({\mathbf{x}})\leq B_{g}$ and 5, There exists a point $\mathbf{y}$ such that $g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$.
-Since $g(\mathbf{z})+\frac{\overline{\rho}}{2}\|\mathbf{z}-\mathbf{x}\|^{2}$ is $(\overline{\rho}-\rho_{g})$-strongly convex, we obtain that for any $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$, it holds
-(48) $g(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}+\left\langle\boldsymbol{\zeta}_{g},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B.$
-Applying Young’s inequality to Eq. 48 and noticing $g({\mathbf{x}})\geq 0$, we derive
-$-\frac{\left\|\boldsymbol{\zeta}_{g}\right\|^{2}}{2(\overline{\rho}-\rho_{g})}\leq\left\langle\boldsymbol{\zeta}_{g},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B-g({\mathbf{x}})\leq-B.$
-This indicates $\|\boldsymbol{\zeta}_{g}\|\geq\sqrt{2B(\overline{\rho}-\rho_{g})}$ and completes the proof.
-Lemma A.3.
-Suppose that 1 and 5 hold and $g$ is convex.
-Then for any ${\mathbf{x}}$ satisfying $g({\mathbf{x}})\leq B_{g}$, it holds that $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\theta$ is given in Lemma A.1.
-Proof A.4.
-Define $\mathbf{x}^{+}=\operatorname{Proj}_{\Omega}(\mathbf{x})\in\underset{{\mathbf{y}}:g(\mathbf{y})\leq 0}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}.$ If $\mathbf{x}\in\Omega$, the result holds trivially.
-Therefore, we assume $\mathbf{x}\not\in\Omega$, which implies $g\left(\mathbf{x}^{+}\right)=0$.
-Under the Slater-type CQ, there exist $\lambda^{+}\geq 0$ and $\boldsymbol{\zeta}_{g}\in\partial g\left(\mathbf{x}^{+}\right)$ such that $\mathbf{x}^{+}-\mathbf{x}+\lambda^{+}\boldsymbol{\zeta}_{g}=\mathbf{0}.$ We assert that $\lambda^{+}>0$, because otherwise, the above inclusion would imply $\mathbf{x}=\mathbf{x}^{+}$, which contradicts the assumption that $\mathbf{x}\not\in\Omega$.
-By the convexity of $g$, we have
-(49) $\lambda^{+}\left(g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\right)\geq\lambda^{+}\left\langle\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle=\left\langle\lambda^{+}\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle=\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}.$
-Dividing both sides of the inequalities by $\lambda^{+}$ in Eq. 49 yields
-$g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\geq\frac{\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}}{\lambda^{+}}=\|\boldsymbol{\zeta}_{g}\|\|\mathbf{x}-\mathbf{x}^{+}\|\geq\sqrt{2B(\overline{\rho}-\rho_{g})}\left\|\mathbf{x}-\mathbf{x}^{+}\right\|,$
-where the last inequality is by Lemma A.1.
-Noting that $g(\mathbf{x})=g_{+}(\mathbf{x})$ for $\mathbf{x}\in\Omega$ and $g\left(\mathbf{x}^{+}\right)=0$, we obtain $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, and complete the proof.
-Finally, we remove the convexity assumption from the last lemma and present the following result.
-Lemma A.5.
-Suppose that 1 and 5 hold, and $\rho_{g}>0$.
-Then for any ${\mathbf{x}}$ satisfying $\textnormal{dist}({\mathbf{x}},\Omega)\leq\min\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\}$, it holds that $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\gamma=\sqrt{\frac{B(\overline{\rho}-\rho_{g})}{2}}$.
-Proof A.6.
-Define $\mathbf{x}^{+}=\operatorname{Proj}_{\Omega}(\mathbf{x})\in\underset{{\mathbf{y}}:g(\mathbf{y})\leq 0}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}.$ If $\mathbf{x}\in\Omega$, the result holds trivially.
-Therefore, we assume $\mathbf{x}\not\in\Omega$, which implies $g\left(\mathbf{x}^{+}\right)=0$.
-It is Straightforward that $\mathbf{x}^{+}$ is also the optimal solution of the convex problem $\underset{{\mathbf{y}}:g(\mathbf{y})+\frac{\overline{\rho}}{2}\|{\mathbf{y}}-\mathbf{x}^{+}\|^{2}\leq 0}{\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$.
-By 5, the Slater’s condition holds for the above problem.
-Hence, there exist $\lambda^{+}\geq 0$ and $\boldsymbol{\zeta}_{g}\in\partial g\left(\mathbf{x}^{+}\right)$ such that $\mathbf{x}^{+}-\mathbf{x}+\lambda^{+}\boldsymbol{\zeta}_{g}=\mathbf{0}.$ We assert that $\lambda^{+}>0$, as otherwise, the above inclusion would imply $\mathbf{x}=\mathbf{x}^{+}$, which contradicts to $\mathbf{x}\not\in\Omega$.
-By the $\rho_{g}$-weak convexity of $g$, we have
-(50) $\lambda^{+}\left(g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\right)\geq\lambda^{+}\left(\left\langle\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}\right)$
-$=$ $\left\langle\lambda^{+}\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle-\frac{\lambda^{+}\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}=\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}-\frac{\lambda^{+}\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}.$
-Since $0<g({\mathbf{x}})\leq l_{g}\textnormal{dist}\left(\mathbf{x},\Omega\right)\leq B_{g}$, dividing both sides of Eq. 50 by $\lambda^{+}$ yields
-$g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)$ $\geq\frac{\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}}{\lambda^{+}}-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}=\|\boldsymbol{\zeta}_{g}\|\|\mathbf{x}-\mathbf{x}^{+}\|-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}$
-$\geq\sqrt{2B(\overline{\rho}-\rho_{g})}\left\|\mathbf{x}-\mathbf{x}^{+}\right\|-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2},$
-where the last inequality follows from Lemma A.1.
-Noting $g(\mathbf{x})=g_{+}(\mathbf{x})$ for $\mathbf{x}\not\in\Omega$, $g\left(\mathbf{x}^{+}\right)=0$ and $\|\mathbf{x}-\mathbf{x}^{+}\|=\textnormal{dist}({\mathbf{x}},\Omega)\leq\frac{\sqrt{2B(\overline{\rho}-\rho_{g})}}{\rho_{g}}$, we obtain $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$ and complete the proof.
+We would like to express our sincere gratitude to Dr.
+Yankun Huang and Prof.
+Qihang Lin for providing their code in [27].
+We also thank the three anonymous referees and the associate editor for their constructive comments.
+Appendix A Proof of Lemma 2.4
+Proof A.1.
+Let $\mathcal{J}_{3}=\{i:g_{i}({\mathbf{x}})\geq 0\}$.
+Then by the condition on ${\mathcal{J}}_{1}$, it holds $\mathcal{J}_{1}\cap\mathcal{J}_{3}\neq\emptyset$, and the set difference $\mathcal{J}_{1}\cap\mathcal{J}_{3}^{c}\subset\{i:g_{i}({\mathbf{x}})<0\}$.
+Since $0<g_{+}({\mathbf{x}})\leq B_{g}$, then by 2, there exists $\mathbf{y}\in\mathrm{relint}(\mathcal{X})$ such that $g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for all $i\in[m]$.
+Now, for any $\boldsymbol{\zeta}_{g_{i}}\in\partial[g_{i}({\mathbf{x}})]_{+},{i\in\mathcal{J}_{1}}$, any $\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}({\mathbf{x}}),{i\in\mathcal{J}_{2}}$, and any ${\mathbf{u}}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$, we have
+$\sum_{i\in\mathcal{J}_{1}}[g_{i}(\mathbf{x})]_{+}+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}g_{i}(\mathbf{x})+\left\langle\sum_{i\in J_{1}}{\boldsymbol{\zeta}}_{g_{i}}+\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$
+$=$ $\sum_{i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}}\left(\textstyle g_{i}(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}\right)+\textstyle\left\langle\sum_{i\in{\mathcal{J}}_{1}\cap{\mathcal{J}}_{3}}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$
+$\quad+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\left(\textstyle g_{i}(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}\right)+\left\langle\textstyle\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}{\boldsymbol{\zeta}}_{g_{i}},\mathbf{y}-\mathbf{x}\right\rangle$
+(43) $\leq$ $\sum_{i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}}\left(g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\right)+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\left(g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\right)\leq-B,$
+where the first equality follows from $g_{i}(\mathbf{x})]_{+}=g_{i}(\mathbf{x}),\forall\,i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3},$ and $g_{i}(\mathbf{x})]_{+}=0$, ${\boldsymbol{\zeta}}_{g_{i}}=\mathbf{0},\forall\,i\in\mathcal{J}_{1}\cap\mathcal{J}_{3}^{c}$, the first inequality comes from the $(\overline{\rho}-\rho_{g})$-strongly convexity of $\sum_{i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}}(g_{i}(\cdot)+\frac{\overline{\rho}}{2}\|\cdot-\mathbf{x}\|^{2})+\iota_{\mathcal{X}}(\cdot)$ and the convexity of $\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\big{(}g_{i}(\cdot)+\frac{\overline{\rho}}{2}\|\cdot-\mathbf{x}\|^{2}\big{)}$, as ${\mathcal{J}}_{1}\cap{\mathcal{J}}_{3}\neq\emptyset$ and $\lambda_{i}\geq 0,\forall\,{i\in J_{2}}$, the last inequality holds by ${\mathcal{J}}_{1}\cap{\mathcal{J}}_{3}\neq\emptyset$, $\lambda_{i}\geq 0,\forall\,{i\in J_{2}}$, and $g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for all $i\in[m]$.
+The inequality A.1, together with Young’s inequality, gives
+$-\frac{\left\|\sum_{i\in J_{1}}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}}+\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}\boldsymbol{{\boldsymbol{\zeta}}}_{g_{i}}\right\|^{2}}{2(\overline{\rho}-\rho_{g})}$
+$\textstyle\leq$ $\left\langle\sum_{i\in J_{1}}{\boldsymbol{\zeta}}_{g_{i}}+\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}\partial g_{i}({\mathbf{x}})+{\mathbf{u}},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$
+$\textstyle\leq$ $-B-\sum_{i\in\mathcal{J}_{1}}[g_{i}(\mathbf{x})]_{+}-\sum_{i\in\mathcal{J}_{2}}\lambda_{i}g_{i}(\mathbf{x})\leq-B,$
+where the last inequality holds by $g_{i}({\mathbf{x}})\geq 0,\forall\,i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}$, and $\lambda_{i}g_{i}({\mathbf{x}})\geq 0,\forall\,i\in J_{2}$.
+Since the choice of subgradients and $\mathbf{u}$} is arbitrary, the proof is completed.
+Appendix B Problems 2–3 in Section 4.1 satisfy 2
+We first verify in this section that the fairness-constrained logistic‐loss problem in Section 4.1 satisfies 2.
+Let $\rho_{g}$ and $l_{g}$ denote the weak convexity modulus and Lipschitz continuity modulus of $\mathbf{g}$} in Problem 2, respectively.
+Under mild conditions, the following lemma establishes the Slater-type CQ.
+Lemma B.1 ([27, Theorem B.1]).
+[…]
+$q({\mathbf{x}}):=\max\left\{\max_{i=1,\dots,n_{\mathrm{p}}}\!\frac{\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{p}})(1-\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{p}}))}{n_{\mathrm{p}}},\max_{i=1,\dots,n_{\mathrm{u}}}\!\frac{\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{u}})(1-\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{u}}))}{n_{\mathrm{u}}}\right\}.$
+Then the following statements hold:
+  • $\mathrm{(i)}$
+    Let $q_{\mathrm{min}}=\inf\{q(\mathbf{x})||\overline{\Psi}(\mathbf{x})\mid\in[(\kappa+\underline{\kappa})/2,(\kappa+\bar{\kappa})/2]\}.$ It holds $q_{\mathrm{min}}>0$.
+  • $\mathrm{(ii)}$
+    Let $\underline{q}\in(0,q_{\mathrm{min}}]$.
+Then Problem 2 in Section 4.1 satisfies the Slater-type CQ with any $\bar{\rho}>\rho_{g}$, and with parameters $0<B\leq\min\left\{\frac{\kappa-\underline{\kappa}}{2},\frac{\underline{q}^{2}}{\rho_{g}+4\bar{\rho}},\frac{\underline{q}^{2}\kappa}{l_{g}^{2}}\right\}$ and $0<B_{g}\leq\min\left\{\frac{\bar{\kappa}-\kappa}{2},\frac{\underline{q}^{2}}{\rho_{g}+4\bar{\rho}},\frac{\underline{q}^{2}\kappa}{l_{g}^{2}}\right\}.$
+We note that though $q_{\mathrm{min}}$ is difficult to compute, for our tested Problem 2, we can directly establish a positive lower bound on $q({\mathbf{x}})$ uniformly about ${\mathbf{x}}\in{\mathcal{X}}$.
+In particular, let $D_{a}=\min\left(\min_{1\leq i\leq n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|,\min_{1\leq j\leq n_{\mathrm{u}}}\|{\mathbf{a}}_{j}^{\mathrm{u}}\|\right)$, i.e., there is ${\mathbf{a}}_{i}^{\mathrm{p}}$ or ${\mathbf{a}}_{j}^{\mathrm{u}}$ such that $\|{\mathbf{a}}_{i}^{\mathrm{p}}\|=D_{a}$ or $\|{\mathbf{a}}_{j}^{\mathrm{u}}\|=D_{a}$.
+Since $\|{\mathbf{x}}\|\leq D,\forall\,{\mathbf{x}}\in{\mathcal{X}}$, it holds $|{\mathbf{x}}^{\top}{\mathbf{a}}^{\mathrm{p}}_{i}|\leq D_{a}D$ or $|{\mathbf{x}}^{\top}{\mathbf{a}}^{\mathrm{u}}_{j}|\leq D_{a}D.$ Consequently, we have $q({\mathbf{x}})\geq\frac{\sigma(D_{a}D)\bigl{(}1-\sigma(D_{a}D)\bigr{)}}{\max\{n_{\mathrm{p}},\,n_{\mathrm{u}}\}}=:\tilde{q},\forall\,{\mathbf{x}}\in{\mathcal{X}}$, and thus $q_{\mathrm{min}}\geq\tilde{q}$.
+Therefore, we can set $\underline{q}=\tilde{q}$ in the above Statement (ii).
+Next, we verify that Problem 3 defined in Section 4.1 also satisfies 2.
+For each $i\in\{2,\ldots,M\}$, let $g_{i-1}$ be the component constraint function that corresponds to ${\mathcal{D}}_{i}$ in Problem 3, and suppose that ${\mathbf{a}}_{i}]_{1}=1,\forall\,{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}$.
+Under this assumption, one can easily show that for any $i\in\{2,3,\ldots,M\}$,
+(44) $\|\nabla g_{i-1}(\mathbf{x})\|$ $\geq\frac{1}{\left|\mathcal{D}_{i}\right|}\sum^{M}_{\begin{subarray}{c}l=2\\ l\neq i\end{subarray}}\sum_{{\mathbf{a}}_{i}\in\mathcal{D}_{i}}\sigma\left(\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}\right)\left(1-\sigma\left(\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}\right)\right)\left[\mathbf{a}_{i}\right]_{1},$
+by the fact $\|\nabla g_{i-1}(\mathbf{x})\|\geq\|\nabla_{[{\mathbf{x}}_{i}]_{1}}g_{i-1}(\mathbf{x})\|$.
+Let $D_{a}=\max_{i=2,3,\ldots,M}\{\min_{{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}}\|{\mathbf{a}}_{i}\|\}$, i.e., for each $i\in\{2,3,\ldots,M\}$, there is ${\mathbf{a}}_{i}\in{\mathcal{D}}_{i}$ such that $\|{\mathbf{a}}_{i}\|\leq D_{a}$.
+Since we have $\|{\mathbf{x}}\|\leq D,\forall\,{\mathbf{x}}\in{\mathcal{X}}$, it holds $|\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}|\leq 2D_{a}D$.
+By an analogous argument, we establish that for all $i\in\{2,3,\ldots,M\}$,
+$\|\nabla g_{i-1}(\mathbf{x})\|\geq\frac{\sigma(2D_{a}D)\bigl{(}1-\sigma(2D_{a}D)\bigr{)}}{\max_{i=2,3,\ldots,M}\{|\mathcal{D}_{i}|\}}=:\tilde{q}$.
+Hence, following steps (83)–(85) in the proof of [27, Theorem B.1], we are able to show that Problem 3 satisfies the Slater-type CQ as defined in 2. We do not repeat the details here.